### v1

# captionUnknown document class (or package)

Abstract
As large language models (LLMs) evolve, their integration with 3D spatial data (3D-LLMs) has seen rapid progress, offering unprecedented capabilities for understanding and interacting with physical spaces. This survey provides a comprehensive overview of the methodologies enabling LLMs to process, understand, and generate 3D data. Highlighting the unique advantages of LLMs, such as in-context learning, step-by-step reasoning, open-vocabulary capabilities, and extensive world knowledge, we underscore their potential to significantly advance spatial comprehension and interaction within embodied Artificial Intelligence (AI) systems. Our investigation spans various 3D data representations, from point clouds to Neural Radiance Fields (NeRFs). It examines their integration with LLMs for tasks such as 3D scene understanding, captioning, question-answering, and dialogue, as well as LLM-based agents for spatial reasoning, planning, and navigation. The paper also includes a brief review of other methods that integrate 3D and language. The meta-analysis presented in this paper reveals significant progress yet underscores the necessity for novel approaches to harness the full potential of 3D-LLMs. Hence, with this paper, we aim to chart a course for future research that explores and expands the capabilities of 3D-LLMs in understanding and interacting with the complex 3D world. To support this survey, we have established a project page where papers related to our topic are organized and listed: https://github.com/ActiveVisionLab/Awesome-LLM-3D.
Index Terms:
3D Scene Understanding, Large Language Models, Vision Language Models, Computer Vision.
1 Introduction
The advent of Large Language Models (LLMs) has marked a transformative era in natural language processing, enabling machines to understand, generate, and interact with human language in previously unimagined ways. However, the physical world around us is inherently three-dimensional, and understanding spatial 3D environments is crucial for many real-world applications that involve perception, navigation, and interaction within these 3D spaces. With recent advancements, the application of LLMs has extended well beyond text. The fusion of LLMs with 3D data presents a unique opportunity to enhance computational models’ understanding of and interaction with the physical world, leading to innovations across various domains, including autonomous systems [1, 2, 3, 4, 5], augmented reality [6, 7, 8, 9], robotic navigation [10, 11, 12], and robotic manipulation [13, 14, 15].
Recent works have demonstrated the potential of integrating LLMs with 3D data to interpret, reason, or plan in complex 3D environments, by leveraging the inherent strengths of LLMs, including zero-shot learning [16, 17], advanced reasoning [13, 18, 19], and extensive knowledge [20, 21]. However, the integration of LLMs with 3D data is not trivial. Issues such as 3D data representation, model scalability, and computational efficiency remain significant hurdles. Moreover, ensuring the models can operate in real-world settings requires overcoming obstacles related to data diversity and environmental complexity. Addressing these challenges is crucial for the full realization of LLMs’ potential in 3D applications to create dynamic and context-aware AI systems.
This survey paper provides a critical examination of the intersection of LLMs with 3D data, offering an exhaustive overview of current methodologies, applications, and challenges in this domain. We start by providing the relevant background on common 3D representations, a brief introduction to LLMs and an overview of Vision-Language Models (VLMs) and Vision Foundation Models (VFMs) in Sec. 2. In Sec. 3, we detail the 3D vision language tasks that current methods aim to solve, outlining the current evaluation metrics and protocols.
Figure 1: A general timeline of methods leading up to Large Language Models (LLMs) in 3D.
Next, in Sec. 4, we analyze the data formats, processing techniques, and model architectures that have shown promise in enhancing 3D understanding through the lens of LLM capabilities. We showcase a variety of areas where the amalgamation of LLMs and 3D data has been successfully demonstrated, such as: using LLMs’ world-knowledge [20, 22] and reasoning capabilities [21, 23] to enhance 3D task performance, using LLMs as multi-modal interfaces [24, 18] and embodied agents [15, 13], or generating complex scenes with LLMs [25, 26]. Beyond LLMs, a few research works have proposed end-to-end architectures that unify 3D perception with language capabilities [27, 28]. Moreover, a wide range of works have explored distilling the knowledge of off-the-shelf 2D Vision Language Models (VLMs) for open vocabulary 3D scene understanding [29, 30] as well as text-driven 3D generation [31, 32]. This survey provides a complete overview of these methods in Sec. 5 to show a complete picture of the field of 3D+Language. We then outline the datasets used to train and evaluate these methods in Sec. 6. Finally, Sec. 7 highlights the challenges and potential future research directions in this field.
2 Background
This section offers essential background knowledge on 3D representations, Large Language Models (LLMs), 2D Vision-Language Models (VLMs), and Vision Foundation Models (VFMs).
2.1 3D Representations
Choosing a 3D representation to describe, model, and understand our world is a crucial topic that contributes to the understanding of the current progress in 3D-LLMs. It is also a fundamental research area in computer vision. The field has recently seen substantial growth due to advancements in deep learning, computational resources, and 3D data availability. We briefly describe the most common 3D representations currently in use.
Point Clouds represent 3D shapes with a set of data points in space, storing each point position in a 3D Cartesian coordinate system. In addition to storing the position, other information (e.g. color, normal) can also be stored with each point. Point cloud-based methods [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] are noted for their low storage footprint but lack surface topology information. Typical sources for obtaining point clouds include LiDAR sensors, structured light scanners, time-of-flight cameras, stereo-views, photogrammetry, etc.
Voxel Grids [44, 45, 46, 47, 48, 49, 50, 51] consist of unit cubes in 3D space, akin to pixel representations in 2D [52]. Each voxel minimally encodes the occupancy information (binary or probabilistically) but can additionally encode distance to a surface, as in the Signed Distance Function (SDF) [53, 54, 55, 56, 57] or Truncated Signed Distance Function (TSDF) [58, 59, 60, 61, 62, 63]. Nevertheless, the memory footprint can become excessive when high-resolution details are required.
Polygonal Mesh representations are composed of vertices and surfaces and describe complex 3D shapes compactly. However, their unstructured and non-differentiable nature [64] poses challenges in integrating them with neural networks for end-to-end differentiable pipelines. Some solutions addressing this problem, such as gradient approximation-based approaches [65, 66, 67, 68], can only use handcrafted gradient calculation. Other solutions like differentiable rasterizers [69, 70, 71] may lead to imprecise render results such as blurry content.
Neural Fields [72] have gained a growing interest in the 3D research community over recent years, departing from traditional representations reliant on geometric primitives [73]. A neural field is a mapping from spatial coordinates to scene properties (such as occupancy, color, radiance, etc), but unlike a voxel grid – in which the mapping is from a discrete cell to the value at that voxel – in a neural field the mapping is a learned function, typically a multi-layer perceptron. In this way neural fields implicitly learn compact, continuous, and differentiable 3D shape and scene representations.
One set of neural fields has focused on implicit surface representations. Occupancy Networks [74, 75, 76] encode shapes in a continuous 3D occupancy function represented by neural networks, using 3D point locations and features from point clouds, low-resolution voxels, or images to estimate occupancy probabilities. Concurrently, Deep SDF networks [77, 78, 79, 80] use neural networks to estimate SDFs from 3D coordinates and latent vectors. Recent methods like NeuS [79] and NeuS2 [80] notably improve surface reconstruction fidelity and efficiency for both static and dynamic objects.
Another set of approaches, referred to as Neural Radiance Fields (NeRF) [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94], display strong photorealistic rendering capabilities of the 3D world. These methods encode scene details with positional encoding techniques [95, 96, 90, 97], and utilize MLPs to predict radiance values (color and opacity) along camera rays. However, the necessity for MLPs to infer color and occupancy details for each sampled point in space (including those in empty spaces) requires significant computational resources. Thus, there is a strong incentive to reduce NeRFs’ computational overhead for real-time applications.
Hybrid Representations attempt to combine NeRF techniques with traditional volumetric-based approaches, facilitating high quality, real-time rendering [98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. For instance, combining voxel grids [95, 96, 98, 97] or multi-resolution hash grids [102] with neural networks greatly reduces the NeRF’s training and inference time.
3D Gaussian Splatting [103, 106, 107] is a variation on point clouds where each point contains additional information representing the radiance being emitted in the spatial region around that point as an anisotropic 3D Gaussian ‘blob’. These 3D Gaussians are typically initialized from SfM point clouds [108], and optimized using differentiable rendering. 3D Gaussian Splatting achieves state-of-the-art novel-view synthesis at a fraction of NeRF’s computation by utilizing efficient rasterization [109] rather than raytracing.
2.2 Large Language Model (LLM)
Traditional natural language processing (NLP) encompasses a wide range of tasks aimed at enabling systems to understand, generate and manipulate text. Early approaches in NLP relied on techniques like rule-based systems, statistical models, and earlier neural architectures like recurrent neural networks [110, 111]. The recent introduction of Large Language Models (LLMs), employing transformer architectures [95] and training on vast text corpora [112], has achieved unprecedented performance and sparked a new wave of excitement in the field. Since this paper focuses on 3D LLMs, we provide relevant background knowledge of LLMs here. For an in-depth exploration of LLMs at large, we refer to the latest surveys in the area [113, 114, 115, 116, 117].
2.2.1 LLM Architectures
In the context of LLMs, the “encoder-decoder” and “decoder-only” architectures are prominently used for NLP tasks.
Encoder-decoder architectures [95, 118, 119] consist of two main components: an encoder $f_{enc}$ and a decoder $f_{dec}$. The encoder and decoder components are typically implemented using transformers [95], which employ attention mechanisms to capture long-range dependencies in the input and output sequences. The encoder takes the input sequence $X=(x_{1},x_{2},\ldots,x_{N})$ and maps it into a sequence of latent representations $H=(h_{1},h_{2},\ldots,h_{N})$ that capture the contextual information, and the decoder generates the output sequence $Y=(y_{1},y_{2},\ldots,y_{T})$ based on $H$. Mathematically, the encoding process can be expressed as $H=f_{enc}(X)$, and the entire latent sequence $H$ is generated at once from $X$. The decoder, however, generates the output sequence $Y$ sequentially: $y_{t}=f_{dec}(y_{<t},H)$ where $y_{<t}=(y_{1},y_{2},\ldots,y_{t-1})$.
Decoder-only architectures [120, 121, 122], on the other hand, are a variant of the transformer architecture that uses only the decoder component. It is particularly suitable for language modeling tasks, where the goal is to predict the next token given the previous tokens. The decoder-only architecture can be mathematically expressed as $y_{t}=f_{dec}(y_{<t})$.
Tokenization is a preprocessing method to break the input text into a sequence of tokens, the basic data unit in language models. The number of tokens is finite, and each token can correspond to a word, sub-word, or a single letter. During inference, the input text is converted to a sequence of tokens and fed to the models, which predict the output tokens that are then converted back to text. The tokenization has a great impact on the performance of the language models, as it affects how the models perceive text. Various tokenization techniques are used, such as word-level tokenization, subword tokenization (e.g. byte-pair encoding [123], WordPiece [124], SentencePiece [125]), and character-level tokenization [126].
2.2.2 LLM Emergent Abilities
One major difference between LLM and traditional non-LLM methods is the emergent abilities that become available in large models but are not present in the small models [115]. The term “emergent abilities” refers to new, complex capabilities that arise as LLMs scale in size and complexity. These abilities enable advanced understanding and generation of natural language, problem-solving across various domains without specific training, and adapting to new tasks through in-context learning. In the following, we introduce several common emergent abilities in the scope of LLMs.
In-Context Learning  [127, 128, 129] refers to the ability of LLMs to understand and respond to new tasks or queries based on the context provided within the prompt, without requiring explicit retraining or fine-tuning. Milestone papers (GPT-2/GPT-3 [130, 127]) demonstrate in-context learning in a few-shot manner, where a model is provided with a few examples of a task within the prompt and then asked to tackle a different example without prior explicit training. State-of-the-art LLMs, such as GPT-4 [131], exhibit remarkable in-context learning abilities, understanding complex instructions, and performing a wide range of tasks from simple translations to generating code and creative writing, all based on the context provided within the prompt.
Reasoning in the context of LLMs, often referred to as “chain-of-thought” prompting [132, 133], involves the model generating intermediate steps or reasoning paths when tackling complex problems or questions. This approach allows LLMs to break tasks into smaller, manageable parts, facilitating a more structured and understandable solution process. To achieve this, training involves datasets that include a variety of problem-solving tasks [134, 135], logical puzzles [136, 137], and datasets designed to mimic reasoning under uncertainty [138]. Current state-of-the-art LLMs [127, 131, 139, 140, 141, 142, 143] demonstrate advanced reasoning capabilities typically when model sizes become larger than 60B to 100B parameters [132].
Instruction-following  [144, 127, 140, 145, 146] refers to the models’ capability to understand and execute commands, or follow instructions as specified by the user. This involves parsing the instruction, understanding its intent, and generating an appropriate response or action. Methods for adapting this ability to novel tasks may require instruction tuning from datasets [114] containing various instructions paired with correct responses or actions. Techniques such as supervised learning, reinforcement learning from human feedback, and interactive learning can further enhance performance.
2.2.3 LLM Fine-tuning
In the context of 3D-LLMs, LLMs are either utilized directly in their pre-trained state or undergo fine-tuning to accommodate novel multi-modal tasks. However, fine-tuning the entirety of an LLM’s parameters presents significant computational and memory challenges due to the large amount of parameters involved. Therefore, parameter-efficient fine-tuning (PEFT) [147] has become increasingly popular for adapting LLMs to specific tasks by updating only a relatively small subset of the model’s parameters instead of retraining the entire model. The following part lists four common PEFT methods used in LLMs.
Low-Rank Adaptation (LoRA) and variants [148, 149, 150] update parameters through low-rank matrices. Mathematically, a forward pass for LoRA [148] during the fine-tuning can be expressed as $h=W_{0}x+BAx$. $W_{0}$ is the frozen weight of LLM, while $BA$ is a low-rank matrix parameterized by newly introduced matrices $A$ and $B$, which are updated at fine-tuning stages. This method offers several distinct benefits. During fine-tuning, only $B$ and $A$ are optimized, significantly reducing the computational overhead associated with gradient calculation and parameter updates. Once the fine-tuning concludes and the weights are merged, there is no additional inference cost compared to the original model, as represented by the equation: $h=(W_{0}+BA)x$. Moreover, there’s no need to save multiple copies of LLMs for different tasks since multiple LoRA instances can be saved instead, reducing the storage footprint.
Layer Freezing [151, 142, 152] freezes selected layers of a pre-trained model while updating others during training. This often applies to layers closer to the model’s input or output, depending on the task’s nature and the model architecture. In 3D-LLM methods, for example, all layers except input and output embeddings [153] might be frozen to mitigate overfitting risks with task-specific datasets, retaining the pre-trained general knowledge and reducing parameters that need to be optimized.
Prompt Tuning [154, 155, 156] guides LLMs to perform specific tasks by framing them within prompts, adjusting model inputs in contrast to conventional fine-tuning which adjusts model parameters. Manual prompt engineering [132, 157, 158, 159] is the most intuitive method, yet it is difficult for experienced prompt-tuning engineers to find the optimal prompt. Another set of approaches is automatic prompt generation and optimization. One prevalent approach is to search for the exact optimal input prompt text, known as hard prompt, such as [160, 161, 162, 163]. Alternatively, one can use optimization approaches to optimize the embedding of the prompt (soft prompt[154, 164, 165, 166, 167, 168, 169, 170].
Adaptive Fine-tuning customizes the model architecture for specific tasks by adding or removing layers or modules [171, 172, 173]. This can include integrating new data modalities, such as visual information, alongside textual data. The core idea of adaptive fine-tuning is leveraging small neural network modules inserted between the layers of a pre-trained model. During adaptive fine-tuning, only the parameters of these adapter modules are updated, while the original model weights remain fixed.
2.3 2D Vision-Language models
Vision-Language Models are a family of models which aim to capture and exploit the relationship between text and images/video, and are capable of performing interactive tasks between the two modalities. The majority of VLMs have transformer-based architectures. By leveraging the attention module, visual and textual contents are conditioned on each other, which enables mutual interaction. In the following paragraphs, we briefly cover the applications of VLMs to discriminative and generative tasks.
Discriminative tasks involve predicting a certain characteristic about the data. VLMs, such as CLIP [174] and ALIGN [175], have shown exceptional performance in zero-shot transferability to unseen data in image classification. Both models comprise two modules: a visual encoder and a text encoder. Given an image and its category, CLIP and ALIGN are trained by maximizing the similarity between the image embedding and the textual embedding of the sentence “a photo of a {image category}”. The zero-shot transferability is achieved by replacing “{image category}” with possible candidates during inference and searching for sentences matching best with the image. These two works have inspired numerous follow-ups such as [176, 177, 178, 179], which further boost image classification accuracy. The models can also distill learned knowledge for other tasks, including object detection [180, 181, 182, 183], image segmentation [184, 185, 186, 187, 188], document understanding [189, 190] and video recognition [191].
Generative tasks utilize VLMs to generate text or images from input data. By leveraging large-scale training data, a single VLM can usually perform several image-to-text generation tasks, such as image captioning and visual question answering (VQA). Notable examples include SimVLM [192], BLIP [193], and OFA [194], etc. More powerful VLMs such as BLIP-2 [195], Flamingo [196], and LLaVA [197], are able to handle multi-turn conversations and reasoning based on the input image. Following the introduction of the diffusion model, text-to-image generation has also become a focus of the research community [198, 199]. By training on a large amount of image-text pairs, diffusion models [200, 201] can generate high-quality images based on textual input. This capability has also been extended to generating videos [202], 3D scenes [31], and dynamic 3D objects  [203]. Along with generative tasks, it is also possible to edit existing images through the textual prompt [204, 205, 206, 207].
2.4 Vision Foundation Models (VFMs)
Vision Foundation Models (VFMs) are large neural networks that aim to extract image representations that are sufficiently diverse and expressive to be directly deployed in various downstream tasks, mirroring the role pre-trained LLMs serve in downstream NLP tasks. One notable example is DINO [208], which uses a self-supervised teacher-student training paradigm. The representation learned achieves good results on both image classification and semantic image matching. The attention weights in DINO can also serve as segmentation masks for semantic components of the observed scene. Followup works like iBOT [209] and DINOv2 [210] further improve the representation by introducing the masked image modeling (MIM) loss. SAM, a transformer-based image segmentation model [211], is trained on a dataset consisting of 1.1 billion images with semantic masks and demonstrates strong zero-shot transfer capability. DINO (Zhang et al.[212] – not to be confused with DINO (Caron et al.[208] – employs a DETR-like [213] architecture and mixed query selection for object detection. The follow-up work Grounding-DINO [214] introduces text supervision to boost accuracy. Stable Diffusion [201], a text-to-image generator, has also been used as a feature extractor for ‘real’ images by running a single diffusion step on clean, or artificially noised images and extracting intermediate features [215, 169] or attention masks [216]. These features have been recently used for segmentation [216] and image matching [215, 169] tasks, due to the size and diversity of the training sets used for diffusion models, and due to observed emergent properties of diffusion features, such as zero-shot correspondence across images [215].
3 Tasks and Metrics
To understand the role of language in 3D understanding, it is important to first understand the tasks that 3D vision-language models attempt to solve. Research has grown to include a broad spectrum of research tasks, each with their own set of commonly used datasets and evaluation metrics. Here, we aim to list current 3D vision-language tasks, and their corresponding evaluation metrics. We broadly categorize the tasks by their input and output modalities.
We then begin our analysis of methods for solving these tasks in Sec. 4 and Sec. 5. Then, in Sec. 6, we detail the datasets that are currently used for training and evaluation for these tasks.
3.1 3D Captioning (3D $\rightarrow$ Text)
Given the 3D data of a scene or object, the task of 3D captioning is to generate a corresponding short, natural language description. Here, we decompose this task into a few common variants of the problem, based on the type of data being captioned, and the type of captions that are generated.
Object-Level Captioning requires the model to generate a short, natural language description of a single 3D object. This caption should focus on the key characteristics of the object, including its shape and semantic characteristics.
Scene-Level Captioning refers to the task of generating a short, natural language caption for an entire 3D scene. These captions typically focus on global scene information (such as room types and styles), key objects in the scene, and their relationships. We consider “grounded captioning”, where the model outputs a description of the relationships between objects in the scene, potentially alongside positional information for those objects, to be a variant of scene captioning.
3D Dense Captioning refers to the joint task of localizing instances of objects in a 3D scene and describing them using natural language captions. In this instance, the output may also contain positional information about the objects being captioned. Often, the referring descriptions from 3D grounding datasets are used to produce the captioning and location data needed for 3D dense captioning. For example, captions in Scan2Cap [217] are generated using the referring expressions from ScanRefer [218].
Evaluation Metrics for 3D captioning require comparing the generated captions against ground truth captions for testing samples.
Exact Match (EM) requires that the generated caption exactly matches the ground truth. Exact Match has different accuracy thresholds, denoted as EM@$K$, which means that the correct answer is within the top “$K$” answers generated by the model. Commonly used thresholds are EM@1 and EM@10. However, natural language captions with the same semantic meaning can be represented in many ways, so the dominant metrics for captioning are automated text generation metrics [219] that aim to measure matching n-grams or semantic similarity rather than complete sentence matches. BLEU [220] matches n-grams between the predicted and true captions, with “BLEU@$x$” referring to matching n-grams of length “$x$” (typical values are in the range 1-4). This still requires matching exact words, but is slightly more robust to rearrangements in phrasing. ROUGE [221] similarly aims to match n-grams, with the commonly used ROUGE-L focusing on the structural similarity of sentences. METEOR [222], is based upon the precision and recall of unigram matches, with “matches” also existing between synonyms and words which are morphological variants of each other. CIDEr [223] weights n-grams by their frequency, with higher frequency n-grams given lower weights. As the above metrics rely on n-gram matches, they cannot account for different by semantically similar words. Thus, various metrics that measures semantic content overlap via similarity in learned embedding spaces (e.g. , SentenceSim [224] and BERT Score [225]) have been introduced.
For dense captioning, where captions are localized to parts of the scene, adjusted benchmarks are needed. Often, BLEU, ROUGE, METEOR and CIDEr scores are still used, however the score is set to zero if the intersection over union (IoU) between the predicted bounding box and the object is less than a threshold “$k$”. Typical “$k$” values are 0.25 and 0.5 [19, 217, 226]. However, these metrics focus on the captioning recall while ignoring false positives. This is addressed by more recent work that additionally measures the precision and F-1 score of generated captions w.r.t. the BLEU, ROUGE, METEOR and CIDEr metrics [227].
3.2 3D Grounding (3D + Text $\rightarrow$ 3D Position)
Given a 3D scene and a “referring expression” that describes an object in the scene relative to other objects, 3D grounding involves generating a position, bounding box or segmentation mask for the target object(s).
Single-Object Grounding involves locating a single query object in a scene given reference information, such as language descriptions [218, 228] or additional gestures [229].
Multi-Object Grounding involves locating multiple objects using a referring expression. There are two main variants for this kind of grounding. The first involves a single-sentence description that may be ambiguous, potentially referring to zero, one, or multiple target objects of the same category in the 3D scene [230]. The second variant uses paragraph-length referring expressions that describe multiple objects belonging to potentially different categories, and the spatial relationships between them [231].
Evaluation Metrics for 3D grounding require comparing predicted locations, most often in the form of bounding boxes, with ground truth locations of objects from testing samples. Acc@$K$IoU [218] is the widely used metric in 3D visual grounding, which measures the percentage of positive predictions which have a intersection over union (IoU) with the ground truth greater than a threshold $K$, which is usually set to 0.25 or 0.5. It is worth noting that some datasets evaluate the performance in different scenarios. For example, ScanRefer [218] divides the datasets into unique/multiple/overall splits. Some methods measure the average IoU [153, 232] whereas other methods measure the average distance between the centers of the bounding boxes [153]. For multi-object grounding, the F1 score is used as the metric [230]. They first get the one-to-one matching between the predicted and ground truth bounding boxes according to IoUs. Then the pairs with IoUs higher than a threshold are regarded as true positives.
3.3 3D Conversation (3D + Text $\rightarrow$ Text)
It is also natural to consider tasks where questions are asked about a 3D scene, either in a single-turn setting or a more natural multi-turn conversational setting.
3D Question Answering (3D-QA) is a task where the model is required to generate answers to questions asked by users given a 3D scene. The topic of questions has a diverse range and the model has to understand both the 3D scene and the question to generate the correct response. The question includes both simple tasks such as determining the existence of an object and more difficult ones like spatial reasoning. As there are several well-established benchmarks, and the majority of questions in the benchmarks are factual with unique answers, 3D-QA is a popular task to evaluate a multi-task model’s capability.
3D Situated Question Answering (3D-SQA) is a special case of 3D-QA. The key difference is that 3D-QA requires the model to answer questions from the perspective of a spectator with access to all information about the scene, while 3D-SQA needs the answer from the perspective of a player in a pre-defined situation. For example, 3D-SQA may ask “how many chairs are in front of me?” given the situation of “standing behind and facing the dining table”.
3D Dialogue requires the model to have a coherent and natural multi-turn conversation with users about the 3D scene, instead of one round QA. For example, a user may want to know about a room, so they continuously ask questions about each part of the room, while the model is expected to respond correctly and coherently.
Evaluation Metrics involve comparing a model’s responses with ground truth responses for testing samples. For 3D-QA and 3D-SQA, the dominant metric is the Exact Match (EM), meaning that the answer generated by the model must exactly match with the correct answer. This is because the majority of questions in existing 3D-QA benchmarks [233, 234, 235, 236] are factual questions where there is only one definitively correct answer. For 3D Dialogue and Task Planning whose answers are non-unique, the semantic metrics such as BLEU [220], ROUGE [221], METEOR [222], CIDEr [223], and SPICE [237] are applied to evaluate the similarity between generated responses and reference answers provided by benchmarks. They are also used in 3D-QA, particularly the ScanQA benchmark, to measure the semantic similarity alongside the accuracy.
3.4 3D Embodied Agents (3D + Text $\rightarrow$ Action)
It is also useful to consider tasks that involve interactions with the 3D scene, conditioned on a specific text prompt describing the desired action or goal.
3D Task Planning is the task where users provide a high-level objective and the model is required to outline low-level steps to fulfill this objective. For example, given a 3D scene of a room, users may ask how to clean the room and the model needs to offer detailed steps to clean it.
3D Navigation refers to the task of enabling 3D agents, such as robots or virtual characters, to move and orient themselves within 3D spaces. This involves understanding and interpreting the 3D environment, identifying obstacles, and planning safe, efficient paths to reach designated goals.
3D Manipulation refers to the ability of 3D agents to physically interact with objects in their environment. This can range from picking up and moving objects to more complex sequences of actions such as assembling parts or opening doors.
Evaluation Metrics for 3D Task Planning also rely on matching the textual/token output of a model with the ground truth actions for test samples. BLEU [220], ROUGE [221], METEOR [222], CIDEr [223], and SPICE [237] are applied to evaluate the similarity between generated responses and ground-truth answers.
For 3D Navigation, there are two primary metrics to evaluate the performance. 1) Success Rate (SR) measures whether the 3D agent reaches the target locations within a predefined distance threshold. 2) Success Rate Weighted by Path Length (SPL) [238], which is calculated as the SR weighted by the ratio of the ground truth length and actual path length, aims to reflect how efficiently the model achieved its goal. Other metrics include Oracle Success Rate (OSR), Trajectory Length (TL) and Goal Process (GP) [239]. In addition to the above metrics that measure whether the agent successfully reached the target and its efficiency, it is also useful to consider how well the agent path matched the path specified by the language (when language is used to specify a detailed paths). One such metric is Success weighted by normalized Dynamic Time Warping (SDTW) [240] which combines the SR with the difference between the agent path and a ground truth path for a given instruction. Please note that our discussion focuses exclusively on metrics used within 3D-LLMs methods. We encourage readers to refer to Gu et al. [241] for summary of navigation metrics.
For 3D Manipulation, the key metric is Success Rate [242], which for manipulation is defined as the number of successful manipulations divided by the total number of task samples. As discussed in Sec. 4.5, different datasets have different conventions for how their actions are represented using text, such as using structured outputs, using normalized numeric scores, or introducing new tokens.
3.5 Text-to-3D Generation (Text $\rightarrow$ 3D)
Beyond using text to describe and interact with existing 3D scenes, it is also possible to generate 3D objects and scene via language specification. Here we give a brief summary of this area, see Lee et al. [243] for a more in-depth survey.
3D Object Generation involves generating 3D models of individual objects from text descriptions. The text input can provide details about the object’s category, attributes, part structure, and other properties that should be reflected in the generated 3D shape.
3D Scene Generation is the task of creating full 3D environments such as rooms or outdoor spaces based on text scene descriptions. This involves generating 3D models for objects specified in the text as well as intelligently arranging and composing multiple 3D object models given the constraints specified in the text, like object categories, counts, spatial relationships, and scene attributes.
Figure 2: Architectures for aligning 3D with text for LLMs. Here we show four high-level architectures: (a) 3D-only model that aligns 3D features to the LLM’s input space, (b) 3D+text model where 3D features and text are both aligned, (c) Q-Former style model where text is used during to condition the alignment of the 3D features, and optionally given to the LLM itself (dashed arrow), and (d) text-only approach which converts 3D representations into text strings, avoiding the need to train an alignment module.
3D Editing refers to modifying existing 3D assets like shapes or scenes based on text instructions. This could involve adding, removing, or transforming objects, changing materials or colors, or altering high-level scene properties according to the given text.
Evaluation Metrics for 3D generation tasks assess the quality of the generated shape/scenes and how well the generated content matches the input text. Common metrics to measure generated geometry include Chamfer Distance (CD) and Mesh-Volume/Surface Distance (MVD). CD is calculated by summing squared point-to-point distances w.r.t. ground-truth 3D data, while MVD calculates the volume/surface between two meshes to measure the geometric error. To assess overall quality, classification accuracy checks if semantic properties are preserved, while the Fréchet Inception Distance (FID) captures realism and diversity. To check if the generated shape matches the input text, it is common to measure the similarity of the text with either aligned embeddings of the 3D shape (e.g. ULIP [244]) or rendered images (e.g. CLIP [174]). It is also common to use human studies for evaluation. However, recent work [245] has shown it is possible to use LVLMs like GPT-v4 as alternative to using human judges. For text-based 3D editing, CD and IoU evaluate how well instructed edits were applied to input geometry without excessive distortion.
4 3D Tasks with LLMs
3D scene understanding tasks have been widely studied. At its core, scene understanding entails recognizing and categorizing all objects present within a designated 3D environment, a process known as semantic [246, 247, 248, 46, 249, 250] or instance-level [251, 252, 253, 254, 255, 256, 257] understanding. This stage is critical, as it forms the basis upon which more nuanced interpretations are built. Subsequently, a higher level of scene understanding focuses on spatial understanding, which refers to the construction of spatial scene graphs  [258, 259] and the semantics of object relationships [260, 261]. Going further, potential interactions can be predicted, such as affordances [262, 13, 14, 15, 263], scene changes [264, 265], and understanding the broader context of the scene, e.g. the functionality and aesthetic styles [266]. 3D data also presents unique challenges that are not present in 2D, such as the relatively high cost of obtaining and labeling 3D data, sparse 3D data structures that are not uniformly dense or aligned to a grid, and the need to reconcile multiple (possibly occluded) viewpoints of the same objects [261, 256]. To this end, researchers have leveraged the power of language, where the semantics and relationships within a 3D world can be embedded. Recent efforts in integrating Large Language Models (LLMs) with 3D data have shown promise in achieving multi-level understanding and interaction, leveraging LLMs’ inherent strengths, namely zero-shot learning, in-context learning, step-by-step reasoning, and extensive world knowledge.
In Sec. 4.1, along with Fig. 2, we provide a brief description of how LLMs process 3D scene information, highlighting how 3D features are aligned with language so that they can be interpreted and reasoned with via LLMs, which is foundational for subsequent sections. The rest of this section is structured to align the taxonomy presented in Fig. 3, which describes the role LLMs have played in solving 3D tasks. We start by showing how the world-knowledge (sometimes referred to as ‘common-sense knowledge’) and reasoning abilities of LLMs can enhance performance on 3D tasks in Sec. 4.2. In Sec. 4.3, we elaborate on how to integrate multiple 3D tasks into one LLM to achieve multi-task learning. We explore how LLMs can be used as a unified interface for combining other modalities in Sec. 4.4. We then describe how LLMs serve as embodied agents to interact with the 3D world in Sec. 4.5. Finally, we present how LLMs serve as assistants for generating semantically diverse 3D objects and scenes in Sec. 4.6.
In addition, we provide Tab. I to contrast 3D-LLMs methods across three axes: 3D components, LLMs components, and the alignment of 3D vision and language, aiming to offer a high-level insight into the various approaches within this evolving field.
Figure 3: Taxonomy of 3D with LLM methods. In Sec. 4, we analyze the role LLMs have played in solving 3D tasks from five perspectives : Enhancing 3D Tasks, Multi-Task Learning, 3D Multi-modal Interfaces, Embodied Agents, and 3D Generation.
4.1 How do LLMs process 3D scene information?
Traditional LLMs are limited to text as both input and output, making the ability to ingest 3D information a primary concern for all 3D-LLM methods. The general idea is to map 3D objects or scene information into the language space, enabling LLMs to understand and process these 3D inputs. Specifically, this typically involves two steps: (i) using a pre-trained 3D encoder to process the corresponding 3D representations, yielding original 3D features; (ii) employing an alignment module to translate these 3D features into 3D tokens that LLMs can process, akin to the tokenization process mentioned in Sec. 2.2.1. Pre-trained LLMs can then use these aligned 3D tokens when generating outputs.
Given the diversity of 3D representations, as described in Sec. 2.1, there are multiple ways to obtain 3D features. As shown in 3D Geometry column in Tab. I, Point Clouds [21, 24, 19, 171, 172, 267, 268, 269, 153, 270, 271, 266, 272] are most common due to their simplicity and compatibility with various pre-trained 3D encoders, which makes them a popular choice for multi-task and multi-modal learning methods. Multi-view Images [11, 14, 17, 20, 270, 273] are also frequently used as research on 2D feature extraction is well-established, which means that 3D feature extraction requires only an additional 2D to 3D lifting scheme. RGB-D data, readily obtained using depth cameras, is commonly used in 3D embodied agent systems to extract viewpoint-relevant information for navigation and understanding. 3D Scene Graphs are a more abstract 3D representation, excelling in modeling the presence of objects and their relationship, and capturing high-level information of scenes. They are often utilized in 3D scene classification [274] and planning tasks [275]. NeRFs are currently less used in 3D-LLM methods [21]. We believe this is due to their implicit nature which makes them more difficult to tokenize and integrate with feed-forward neural networks.
Current methods use different architectures (see Fig. 2) and modules that align 3D features with LLM input spaces (see 3D+LLM column in Tab. I). For models which only accept 3D inputs (Fig. 2a), a linear layer [269, 24, 266] or an MLP [171, 172] is used as the alignment module to transform 3D features into the LLM input space. Models which accept 3D and text as input often use two separate branches to align 3D features and text (Fig. 2b). Some works [171, 172] employ a one-layer vanilla transformer to allow 3D object features to attend to each other during alignment. Others, such as [270, 271], create transformer-based alignment modules, where the standard transformer architecture is tweaked to better suit the different types of 3D data, such as dense point clouds and sparse LiDAR scans. Meanwhile, text is encoded using the pre-existing LLM text embedding table. Other works [153, 268, 276] follow the Q-Former style approach of [195] to align 3D features and text (Fig. 2c), introducing fixed length query tokens as additional input, and following a BERT-based structure to facilitate interaction between 3D and text features during alignment. Mostly, these three types of architectures described above achieve alignment by utilizing 3D captioning datasets [218], where the captioning loss, i.e.the cross-entropy loss between captions generated by LLMs and the brief, ground truth description of the scene, is used to fine-tune the alignment module, while freezing the pre-trained 3D feature extractor and LLM.
Finally, some models [21, 20, 22, 23, 273, 12, 275, 277, 14, 17] employ closed-sourced models like ChatGPT and do not train an alignment module at all (Fig. 2d). Instead of aligning 3D features with the LLM input space, text descriptions are directly generated from the 3D data, such as by describing 3D bounding boxes, positions and relationships, or by using pre-existing captions. These text descriptions are input into ChatGPT. No additional alignment module is proposed in these works, and thus no training is required.
4.2 LLMs for Enhancing 3D Task Performance
LLMs trained on large amounts of data have been shown to acquire common-sense knowledge about the world [278]. The potential of LLMs’ world knowledge and reasoning capabilities has been explored to enhance 3D scene understanding and reformulate the pipelines of several 3D tasks. In this section, we focus on methods which aim to use LLMs to improve the performance of existing methods on 3D vision-language tasks. When applying LLMs to 3D tasks, we can categorize their use into two distinct groups: knowledge-enhanced and reasoning-enhanced approaches. Knowledge-enhanced approaches tap into the vast world knowledge embedded within LLMs to enhance 3D task performance. This may provide contextual insights, filling in knowledge gaps, or might enhance the semantic understanding of 3D environments. Alternatively, rather than relying on their world knowledge, reasoning-enhanced approaches leverage the ability of LLMs to infer step-by-step, thereby offering better generalization to tackle more complex 3D challenges. The following two sections describe these approaches respectively.
3D Component LLM Component 3D+LLM Method Subsection 3D Geometry Vision Model Tuning LLM Abilities Fine-tuning LLM Base # Parameters Hardware Alignment Module Date Chen et al. [274] 3.2 SG CLIP F WK None GPT2 1.5B 1 RTX3080 - 09/22 ConceptFusion [18] 3.2/3.4 RGB-D OpenSeg F IF/R None GPT3 175B 1 3090 - 02/23 ViewRefer [20] 3.2 MVI Multi-View Transformer T WK None GPT3 175B 4 A100 Transformer 03/23 LLM-Grounder [21] 3.2 PC/NeRF OpenScene/LERF F IF/ICL/R None GPT3.5/4 - - - 09/23 Abdelreheem et al. [22] 3.2 Mesh Vision Transformer F IF/WK None GPT3.5 - 1 3090 - 09/23 Transcribe3D [23] 3.2 PC Groupfree Transformer F IF/R None GPT3.5/4 - - - 10/23 zero-shot 3DVG [16] 3.2 RGB-D Mask3d F IF/ICL/R None GPT3.5/4 - - Transformer 11/23 3DAP [273] 3.2 MVI - - IF/ICL PT GPT4V - - - 12/23 3D-LLM [153] 3.3 PC Mask2Former/SAM F IF/ICL/R LF OPT/Flan-T5 9B/2.7B/3B 64 V100 QFormer 07/23 Chat-3D [171] 3.3 PC Point-BERT F IF/ICL/R AF Vicuna 7B - Linear layer 08/23 LEO [270] 3.3 PC/MVI OpenClip/PointNet++ F IF/ICL/R LoRA Vicuna 7B 8 A100 Transformer 11/23 LL3DA [19] 3.3 PC/3D-BB ScanNet scene encoder F IF/ICL/R AF OPT 1.3B 8 3090 QFormer 11/23 Point-LLM [267] 3.3 PC Point-BERT F IF/ICL/R Full LLaMA 7B/13B 8 A100 Linear layer 12/23 GPT4Point [268] 3.3 PC Point-BERT F IF/ICL/R LF OPT/Flan-T5 6.7B/3B 8 A100 QFormer 12/23 Chat-3D v2 [172] 3.3 PC Uni-3D F IF/ICL/R AF Vicuna 7B 4 A40 MLP 12/23 LiDAR-LLM [271] 3.3 PC/VM VoxelNet F IF/ICL/R AF LLaMA 7B 4 A100 Transformer 12/23 3DMIT [269] 3.3 PC EPCL/Uni3D F IF/ICL/R LoRA Vicuna 7B 8 A100 Linear layer 01/24 Scene-LLM [266] 3.3 PC/VM ConceptFusion F IF/ICL/R LF LLaMA-2 7B 32 A100 Linear layer 03/24 Point-Bind [272] 3.4 PC I2P-MAE F IF/ICL/R LF LLaMA 7B 8 A100 Linear layer 09/23 JM3D-LLM [279] 3.4 PC PointNet++/PointMLP/Point-BERT F IF/ICL/R LF Vicuna 7B 3 A100 MLP 10/23 LLM-Planner [12] 3.5 VM Seg&Depth F IF/ICL/R None GPT3 175B - - 03/23 SayPlan [275] 3.5 SG SG generator F IF/ICL/R None GPT4 - - - 07/23 VoxPoser [13] 3.5 RGB-D OWL-ViT/SAM F IF/ICL/R None GPT4 - - - 07/23 UniHSI [277] 3.5 RGB-D - - IF/ICL/R None GPT3.5/4 - 1 A100 - 09/23 LAN-grasp [14] 3.5 MVI OWL-ViT F IF/ICL/R None GPT4 - - - 10/23 Agent3D-Zero [17] 3.3/3.5 MVI - - IF/ICL/R PT GPT4V - - - 03/24 NaviLLM [11] 3.5 MVI EVA-CLIP-Large F IF/ICL/R LF Vicuna 7B 8 A100 Transformer 12/23 MultiPLY [24] 3.3/3.4/3.5 PC ConceptGraph T IF/ICL/R AF Vicuna 13B 128 V100 Linear layer 01/24 ManipLLM [15] 3.5 RGB-D CLIP F IF/ICL/R AF LLaMA 7B 1 A100 Linear layer 03/24 3D-VLA [276] 3.3/3.5 PC/MVI Mask2Former/SAM F IF/ICL/R LoRA Flan-T5 3B 384 V100 QFormer 03/24 PolyGen [280] 3.6 n-gon Mesh Custom Transformer F - Full Custom - 4 V100 - 02/20 LLMR [281] 3.6 Unity, Mesh Dall-E-2, CLIP F IF/ICL/R None GPT4 - 1 3080 - 09/23 3D-GPT [26] 3.6 Blender, Mesh - - IF/R None GPT3.5/4 - - - 10/23 MeshGPT [282] 3.6 Mesh GPT-2-style Transformer F - Full GPT-2-style $\sim$345M 4 A100 - 11/23 ShapeGPT [283] 3.6 SDF 3D VQ-VAE, CLIP F IF/R Full T5 - 4 A100 T5 11/23 Holodeck [284] 3.6 Mesh - - IF/R None GPT4 - 8 RTX 8k CLIP/SBERT 12/23 GALA-3D [25] 3.6 3DGS MVDream, Control-Net F IF/R None Any (GPT3.5) - - - 02/24
TABLE I: Summarization of 3D-LLMs methods. The 3D Geometry column collects the 3D geometric information being used for each method, such as Point Cloud (PC), Multi-view Images (MVI), RGB+Depth (RGB-D), 3D Bounding Box (3D-BB), Scene Graph (SG), Voxel Map (VM) and NeRF. The tuning column outlines whether the vision model is finetuned during training (True/False). IF, ICL, R, and WK denote the LLM Abilities of Instruction-following, In-Context Learning, Reasoning, and World Knowledge. The Fine-tuning column summarizes how LLM components are fine-tuned, such as Prompt Tuning (PT), Low-Rank Adaptation (LoRA), Adaptive Fine-tuning (AF), Layer Freezing (LF), or Full Fine-tuning (Full). The Hardware column shows methods that are trained either with numbers of Nvidia GPU and the specific GPU type or with no training involved.
4.2.1 Knowledge-enhanced approaches
There are several methods that utilize LLM world knowledge. Chen et al. [274] use LLMs for 3D room classification from RGB-D images. Here, the knowledge embedded in LLMs is used to determine the room category based on the object category information contained in the room. Firstly, this approach creates scene graphs from Matterport3D [285] data with nodes for regions and objects, and object nodes linked to room nodes. Next, key objects are selected to form queries per room type. LLMs score descriptions extracted from the selected objects, with the highest score predicting the room label. Spatial information like sizes or locations can also be provided.
ViewRefer [20] uses LLMs to expand the grounding texts with view-related descriptions. For instance, given the original text “Facing the front of the couch, the table that is to the right of the couch”, the LLM is used to create a similar sentence but from another speaker’s point of view, e.g. “With back to the front of the couch, pick the table that is on the left side of the couch”. By rephrasing input texts and their opposite-view synonyms several times, the model improves cross-view grounding. It also adopts a fusion transformer with inter-view attention and includes learnable multi-view prototypes, which capture inter-view knowledge and further boost 3D grounding performance.
Abdelreheem et al. [22] address the problem of semantic correspondence in 3D shapes. They classify 3D shapes by feeding rendered views into a BLIP2 model to generate lists of class proposals. ChatGPT [286] unifies these into a single class per shape. ChatGPT also generates semantic part names and pairwise mappings (e.g. arm $\rightarrow$ wing). A 3D segmenter then segments shapes based on semantic regions, utilizing the part mappings to produce sparse correspondence maps.
The knowledge-enhanced strategies described above enable strong performance especially in zero-shot scenarios where no labeled 3D data is available for a particular object or scene type. This allows open-ended reasoning about object parts, relationships, and semantics beyond fixed ontologies, as demonstrated by (i) Chen et al. [274] generating spatial and semantic object descriptions, (ii) ViewRefer  [20] describing multi-view object relationships, and (iii) Abdelreheem et al. [22] generating and matching object part semantics across shapes.
Method
Object
Captioning
Scene
Captioning
Dense
Captioning
Grounding QA
Situated
QA
Dialogue Planning Navigation Manipulation
3D-LLM [153]
Chat-3D [171]
LEO [270]
LL3DA [19]
PointLLM [267]
GPT4Point [268]
Chat-3D V2 [172]
LiDAR-LLM [271]
3DMIT [269]
MultiPLY [24]
3D-VLA [276]
Agent3D-Zero [17]
Scene-LLM [266]
TABLE II: Comparison of tasks included in various 3D multi-task learning methods. QA stands for Question Answering. The methods are sorted chronologically from top (oldest) to bottom (most recent).
4.2.2 Reasoning-enhanced approaches
Along with world knowledge, the reasoning abilities of LLMs also help tackle other 3D tasks, particularly visual grounding in complicated 3D scenes with detailed geometry and multiple objects. In such cases, text descriptions of the objects should include their appearance and spatial relationship with surrounding items. Ordinary grounding methods [287] often struggle in this setting due to their inability to understand detailed text descriptions. LLM-Grounder [21], Transcribe3D [23], and zero-shot 3DVG [16] approach this problem by leveraging the reasoning capability of LLMs to analyze the text descriptions and generate a sequence of instructions for locating the objects using existing grounding toolboxes. Specifically, an LLM first identifies anchor and target objects from the text descriptions. It then analyzes the spatial relationship (or described attributes) between multiple candidate objects based on their coordinates returned by the grounding tool to select the candidate that best matches the text description. In addition, (i) Transcribe3D [23] and LLM-Grounder [21] employ a multi-round interactive question-and-answer process to help users clarify their intentions, prompting them to provide instructions with more information to arrive at more accurate results, whereas (ii) LLM-Grounder includes multiple grounding tool choices, such as OpenScene [29] or LERF [30] to accommodate different 3D representations like point clouds or NeRF. A common drawback of these approaches is the “blindness” of the LLM since it is provided only with abstract textual descriptions of the 3D scene instead of the original point cloud of the scene. This may lead to the loss of crucial scene details. Consequently, when a 3D scene contains multiple objects of the same class, the absence of necessary scene details means that ambiguities of the text-based references cannot be resolved, which limits overall performance.
Besides visual grounding, the reasoning capability of LLMs facilitates other tasks as well. 3DAP [273] leverages GPT-4V to infer 3D information of an object from its 2D images using a visual prompting technique, where it annotates the input image with a 3D axis to enhance the LLM’s awareness of the 3D scale. ConceptFusion [18] uses GPT3 to generate instructions using pre-defined elementary spatial comparison modules to enable more complex spatial reasoning using their proposed 3D feature map.
4.3 LLMs for 3D Multi-Task Learning
Many works focus on using the instruction-following and in-context learning capabilities of LLMs to unify multiple 3D tasks into a single language space. By employing different text prompts to denote different tasks, these studies aim for LLMs to serve as a unifying dialogue interface. Achieving multi-task learning using an LLM typically involves several key steps, beginning with the construction of 3D-text data pairs [19, 153, 270]. These pairs require crafting task instructions in text form and defining the output for each different task. Next, 3D data (typically in the form of point clouds) are fed into 3D encoders [288, 34] to extract 3D features. Alignment modules [171, 153, 172, 266] are subsequently used to (i) align 3D features with text embeddings from the LLMs across multiple levels (object-level, relationship-level, and scene-level) and (ii) translate the 3D features into tokens interpretable by the LLMs. Finally, an appropriate training strategy [272, 19, 153, 270, 269] needs to be selected, such as single or multi-stage 3D-language alignment training and multi-task instruction fine-tuning.
In the remainder of this section, we explore these aspects in detail. We additionally summarize the scope and capabilities of each method reviewed in this section in Table II.
4.3.1 Data for Multi-Task Learning
As demonstrated in Table II, we divide tasks into four categories: captioning, grounding, question answering (QA), and embodied agent tasks (i.e., planning, navigation, and manipulation). Accordingly, each task’s textual output follows a predefined format. For captioning and QA tasks, the output is plain text and is not constrained by a specific format. The output of the grounding task is a 3D bounding box, typically the coordinates of the referred object’s center along with its 3D size. Normally, the values for point and size are normalized to fall within the range of 0-255 [19] which limits the range of tokens the LLM is required to predict. For planning, the model outputs a series of steps to execute a task in text form, whereas for navigation, the output is a series of spatial coordinates. For manipulation, the output is action sequences in text form. Existing methods follow these guidelines to construct their multi-task instruction fine-tuning datasets.
Once the text format has been decided upon, different methods use different strategies to obtain text annotations for their datasets. Several approaches utilize human labelers to generate ‘ground truth’ annotations for each sample [218, 228, 234, 233], however this can be an expensive and time-consuming process. Another approach is to use ChatGPT [286] to generate text annotations for each sample, which is a strategy utilized by 3DMIT [269], LiDAR-LLM [271], Chat-3D [171] and Chat-3D v2 [172]. Here, 3D scene data is converted into text (usually by describing object bounding boxes and spatial relations textually), and a task description is created to describe the desired output. To guide ChatGPT towards the expected output format for the task, demonstration examples are provided, which allows ChatGPT to perform in-context learning to generate plausible text annotations for other 3D scenes. Alternatively, other multi-task datasets [19, 266] are constructed simply be merging existing 3D Vision-Language (VL) datasets [218, 228, 234, 153]. Some multi-task datasets are constructed using a combination of these three approaches, as seen in LEO [270], LiDAR-LLM [271] and 3D-LLM [153], seeking to combine the accuracy of human annotations with the scalability provided by using LLM-generated annotations.
4.3.2 Training an LLM for multiple 3D tasks
The first step in training an LLM for multiple 3D tasks involves obtaining meaningful 3D features, where the extraction method varies depending on the type of 3D scene. For individual object point clouds, Point-LLM [267], Chat-3D [171], and GPT4Point [268] utilize Point-BERT [288] to extract 3D object features. For indoor scenes, LEO [270] uses PointNet++ [34] for feature extraction, whereas Chat-3D v2 [172] and 3DMIT [269] segment the scene and use Uni-3D [289] to extract features for each segmented part. Meanwhile, MultiPLY [24] integrates the extracted object features into a scene graph [290] to represent the entire scene. 3D-LLM [153] and Scene-LLM [266] lift features from 2D multi-view images into a 3D representation. 3D-LLM [153] lifts 2D semantic features from Mask2Former [291] or SAM [211]. Scene-LLM [266] follows ConceptFusion [18] to fuse both global information and local details, mapping pixel-wise CLIP features into point-wise 3D features. For outdoor 3D scenes, LiDAR-LLM [271] uses VoxelNet [292] to extract 3D voxel features.
For alignment modules, as discussed in Sec. 4.1, various network architectures are used. Notably, MultiPLY [24] employs distinct linear layers to align features from each modality. Chat-3D [171] and Chat-3D v2 [172] use a one-layer vanilla transformer to allow 3D object features to attend to each other during alignment. LEO [270] and LiDAR-LLM [271] utilize modified transformers as their alignment modules to better suit different types of 3D data (dense point clouds vs. sparse LiDAR). LEO [270] modifies the self-attention mechanism to explicitly encode spatial relations between object pairs in point clouds. In contrast, LiDAR-LLM [271] employs both self-attention and cross-attention mechanisms to align Bird’s Eye View (BEV) features with text features. 3D-LLM [153] and GPT4Point [268] adopt Q-Former, while LL3DA[19] adds an additional branch on top of Q-Former that allows query tokens to interact with visual prompts provided by users.
LLMs can be fine-tuned to incorporate multiple 3D tasks using different strategies discussed in Sec. 2.2.3. LEO [270], and 3DMIT [269] use low-rank adaptation (LoRA) for fine-tuning. As a result, the total trainable parameters, including those of alignment modules and 3D encoders, amount to less than 10% of the original LLMs’ parameters, significantly enhancing training efficiency. Chat-3D [171], LL3DA [19], Chat-3D v2 [172], LiDAR-LLM [271], and MultiPLY [24] adopt adaptive fine-tuning. Specifically, these models include modules that align spatial information in 3D scenes with language, e.g. a transformer layer, to capture object relationships. These modules, along with the pre-trained 3D encoders and LLMs, are fine-tuned for alignment. 3D-LLM [153], Scene-LLM [266], Point-LLM [267], and GPT4Point [268] employ layer freezing. By freezing most LLM layers and fine-tuning certain layers like the embedding layer, this strategy preserves linguistic capabilities while improving 3D understanding. Lastly, Agent3D-Zero [17] uses prompt tuning, a training-free method for guiding LLMs toward 3D task comprehension. This approach utilizes tailored prompts, adding grid lines and tick marks onto a BEV image of 3D scene, which help the 2D VLMs understand the 3D geometry.
[…]
In contrast to these single-stage alignment methods, LiDAR-LLM [271], Chat-3D [171], and Chat-3D v2 [172] each employ a 2-stage 3D-language alignment process. LiDAR-LLM [271] focuses on enhancing both local and global scene perception through 3D captioning tasks in two phases: first, concentrating on single-view captions and then expanding to panoramic scene descriptions. They develop instance-level perception capabilities through a blend of captioning and grounding tasks. Chat-3D [171] first aligns 3D objects with text using datasets for 3D object classification [296, 293, 297] aiming to maximize the cosine similarity between mapped 3D object features and object category word embeddings by updating only the alignment module. In the second stage of scene-level alignment, it utilizes ScanRefer [218] to enable captioning capabilities, updating an additional transformer layer specifically to model spatial relationships of objects. Similarly, Chat-3D v2 [172] incorporates object-level and scene-level alignments, with the second stage additionally training a position embedding layer. For training efficiency, both LL3DA [19] and 3DMIT [269] skip the alignment stages and focus solely on the stage of instruction tuning described below.
Nearly all multitask learning methods ultimately require the ability to complete various 3D tasks based on instructions. Thus, as the final stage of training, each method typically uses their own constructed multitask instruction-following dataset for instruction fine-tuning [^1] Instruction fine-tuning refers to the process of further fine-tuning LLMs on a dataset consisting of (INSTRUCTION, OUTPUT) pairs [298]. Since all task outputs are unified into textual form, the training loss used is the standard autoregressive loss used in LLMs. This stage usually involves jointly training the alignment module and the LLM. An exception is Agent3D-Zero [17], which completes various 3D tasks by feeding GPT4V with 2D images from different viewpoints, thus it does not require any training.
4.4 LLMs as 3D Multi-Modal Interfaces
In addition to exploring 3D multi-task learners, some recent studies incorporate information across different modalities to further improve the capability of the models and enable novel interactions. Apart from text and 3D scenes, multi-modal 3D-LLMs may also include 2D images, audio, or touch information in scenes as inputs.
Most works aim to construct a common representation space across different modalities. Since several existing works [299, 300] have already provided pre-trained encoders that map text, images, or audio to common spaces, some works choose to learn a 3D encoder that aligns 3D embeddings to the embedding space of pre-trained encoders for other modalities. JM3D-LLM [279] learns a 3D point cloud encoder that aligns the embedding space of the point cloud to the text-image embedding space of SLIP [301]. It renders a sequence of images of the point cloud and constructs a hierarchical text tree during training to enable a detailed alignment. Point-Bind [272] also learns a similar 3D encoder and aligns it to ImageBind [302] to unify the embedding space of images, text, audio, and point clouds. This enables tackling different tasks such as retrieval, classification, and generation among the various modalities using different task heads. However, one notable limitation is that this approach only works on small-scale object-level scenes as it is computationally expensive for a 3D encoder to process a large scene with millions of points. Moreover, most pre-trained multi-modal encoders like CLIP are designed for single-object scenes and are not suitable for large scenes with multiple objects and local details.
Large scenes instead require more meticulous design to incorporate multiple modalities. ConceptFusion [18] constructs an enhanced feature map that fuses both global information and local details for each constituent image of a large scene. This is achieved by using pre-trained feature extractors [187, 188] that have already been aligned to different modalities including text and audio. It then maps the feature maps to the point cloud of the scene using traditional SLAM approaches. MultiPLY [24] adopts a similar representation to ConceptGraph [290]. It identifies all salient objects in the scene, obtains the global embedding for each object, and finally constructs a scene graph. The resultant representation is a scene embedding aligned with the embedding space of Llama [140]. The embeddings of other modalities including audio, temperature, and sense of touch can also be mapped to the same space using linear projection. All embeddings are tokenized and sent to the LLM at once. Compared with methods on object-level scenes, approaches that can process large scenes reduce costs by relying on pre-trained encoders to bridge the modality gap, instead of learning new ones from scratch.
4.5 LLMs for Embodied Agents
3D embodied agents can be created using the planning, tool-using, and decision-making abilities of LLMs. Such abilities enable LLMs to generate intelligent decisions that encompass navigation within a 3D environment [270, 275, 11], interaction with objects [14], and selecting the appropriate tools to execute specific tasks [24]. This section describes how 3D embodied agents perform the tasks of planning, navigation, and manipulation.
4.5.1 3D Task Planning
For embodied agents, ‘task planning’ refers to the ability to generate steps to execute a specific task, given the task description and the 3D environment. Task planning usually serves as a prerequisite for navigation and manipulation tasks [12, 275] as the accuracy of planning directly influences the performance of subsequent tasks.
LEO [270] and LLM-Planner [12] utilize LLMs to generate step-by-step plans which they adjust dynamically based on environmental perceptions. LEO [270] emphasizes scene-aware planning grounded in the current scene’s configurations, whereas LLM-Planner [12] adopts GPT3 [127] to divide planning into high-level subgoals and low-level actions, and replan when the agent gets stuck during task execution. 3D-VLA [276] integrates 3D perception, reasoning, and action through a generative world model. It focuses on enhancing planning capabilities by utilizing its generative model to predict future state representations (e.g. goal images and point clouds). Agent3D-Zero [17] introduces Set-of-Line Prompting (SoLP) that enhances the VLM’s comprehension of the scene’s geometric aspects by generating diverse observational viewpoints. Specifically, SoLP superimposes grid lines and tick marks onto BEV images, and prompts the VLM to provide more accurate camera positions and orientations, which enables the task of the VLM to understand 3D spatial concepts. UniHSI [277] addresses the task of Human-Scene Interaction (HSI), which involves generating interactions between humans and objects within 3D environments based on input language commands. It uses an LLM as a planner to translate language commands into task plans represented as Chains of Contacts (CoC), a sequence that represents the chronological relationships between human joint points and object positions. While the aforementioned approaches focus on planning within a single scene, SayPlan [275] can handle multiple rooms and floors by (i) leveraging 3D scene graphs for semantic search and (ii) integrating classical path planning with an iterative replanning pipeline for plan refinement.
4.5.2 3D Navigation
3D Navigation refers to the ability of an embodied agent to move and orient itself within 3D environments, often based on visual inputs and language instructions. Each of the described methods – LEO [270], Agent3D-Zero [17], LLM-Planner [12], and NaviLLM [11] – implements 3D navigation in different ways. LEO [270] processes egocentric 2D images and object-centric 3D point clouds alongside textual instructions. It generates a sequence of action tokens that correspond to executable navigation commands like ‘move forward’ or ‘turn right’. LEO employs ‘shortest path navigation trials’, which provide a less noisy and more straightforward learning environment compared to human demonstrations. Agent3D-Zero [17] navigates by continuously selecting new viewpoints based on assessments of the environment. It incorporates historical data from previous viewpoints to refine its navigation path towards specific goals, such as finding a printer in an office setting. LLM-Planner [12] employs a hierarchical approach, that first generates high-level plans as sequences of subgoals, which are then translated into a series of primitive actions by a low-level planner. This makes the overall process adaptable to the immediate environment. NaviLLM [11] transforms various embodied navigation tasks into generation problems using schema-based instructions. These instructions include $4$ elements: a task defined by a word sequence, observations of all reachable viewpoints, the history of past visual observations, and output hints that guide action generation (e.g. selecting a direction or object).
4.5.3 3D Object Manipulation
Manipulation in the context of 3D embodied agents refers to their ability to physically interact with objects, ranging from moving objects to complex sequences such as assembling parts or opening doors. The core idea used to enable LLMs to perform manipulation tasks lies in tokenizing action sequences. To let LLMs output specific actions, it is first necessary to define action tokens that allow LLMs to generate said actions, based on the task and 3D scene context. Subsequently, platforms like CLIPort [242] or motion planning modules in robot arms translate these tokenized actions into physical movements to be executed by the agent.
[…]
While these systems enable robots to perform complex tasks by mapping instructions to actions, they overlook semantic understanding of manipulable objects and are often unable to distinguish between suitable and unsuitable parts for manipulation. To address this issue, VoxPoser [13], LAN-grasp [14], and ManipLLM [15] focus on “affordance”, and create affordance maps to represent objects and features in their surroundings that can be utilized to perform specific tasks, such as graspable handles [14, 15], pressable buttons [15], or movable objects [13]. Specifically, VoxPoser [13] uses an LLM to decompose free-form language instructions, infer affordances and constraints, and compose 3D voxel maps by interacting with VLMs using code interfaces. These maps enable generating closed-loop robot trajectories that are robust to dynamic changes, with capability to learn from online experiences in contact-rich environments. LAN-grasp [14] employs foundation models to deepen robots’ understanding of objects for semantically appropriate grasping by combining multiple models to identify graspable parts, without the need for retraining. ManipLLM [15] predicts manipulation outcomes by identifying 3D coordinates for contact points and gripper orientations from text prompts, RGB images, and depth maps.
4.6 LLMs for 3D Generation
Traditionally, 3D modeling has been a complex time-intensive process with a high barrier to entry, requiring detailed attention to geometry, texture, and lighting to achieve realistic results. In this section, we scrutinize the integration of LLMs with 3D generative techniques, where we show how language provides a way to generate contextualized objects in scenes and provides innovative solutions to 3D content creation and manipulation.
4.6.1 Object-level Generation
Shape-GPT [283] quantizes 3D shapes into discrete “shape word” tokens using a shape-specific 3D VQ-VAE. This enables the integration of shape data into a multi-modal input for the T5 language model [139] along with text and images. This multi-modal representation enables T5 to learn cross-modal interactions for, e.g. text-to-shape generation and shape editing/completion. GPT4Point [268] uses a two-stream approach - aligning point cloud geometry with text via Point-QFormer, which then feeds into coupled LLM and diffusion paths for text comprehension and high-fidelity 3D object generation conforming to text input.
In contrast, MeshGPT [282] and PolyGen [280] do not condition the generation on text, but they still adopt an autoregressive approach akin to sequence modelling in LLMs. MeshGPT uses graph convolutions to encode mesh geometry/topology into rich embeddings compressed via Residual Vector Quantization and fed into a GPT-style transformer to autoregressively predict tokens/embeddings for generating meshes with desired properties. PolyGen [280] is an autoregressive transformer-based model of 3D meshes, which utilizes pointer networks. It consists of a vertex model that unconditionally models mesh vertices, and a face model for mesh faces conditioned on input vertices using autoregressive networks to output face indices and vertex coordinates for generating diverse high-quality meshes.
4.6.2 Scene-scale Generation
Holodeck [284] and GALA-3D [25] employ multi-stage pipelines to progressively refine an initial coarse 3D scene layout from text into a detailed realistic 3D environment. Holodeck employs specialized modules to craft a basic layout, select materials, and incorporate elements like doors and windows based on GPT-4’s spatial reasoning and placement/style suggestions. It then populates the layout with Objaverse assets matched to GPT-4’s textual descriptions. An optimizer arranges these objects adhering to spatial relational constraints obtained from GPT-4 encouraging realistic object layouts and interactions.
GALA-3D [25] first generates coarse layouts from text using an LLM, then translates them into a 3D Gaussian representation. This representation serves as the foundation for creating detailed 3D content using instance-level text-to-image diffusion priors. It employs compositional optimization to fine-tune the parameters of the layout-guided Gaussians, ensuring the final scene aligns with the text for object placement, scale, and interaction.
Both leverage complementary strengths of LLMs to extract high-level semantic layouts and generative models/optimization to convert these into geometrically and physically plausible 3D scenes.
4.6.3 Procedural Generation and Manipulation
LLMR [281], 3D-GPT [26] and SceneCraft [303] adopt modular architectures with specialized components/agents for interactive 3D world creation and code generation from natural language. LLMR consists of distinct components for generating code to construct scenes in Unity, understanding existing scene objects and properties to enable modifications, identifying required functions to execute instructions, and evaluating final code quality. Similarly, 3D-GPT has components for interpreting instructions and determining required generation functions, enriching descriptions with detailed modeling attributes, and translating enriched descriptions to Python code for the Blender API. Overall, these methods showcase task decomposition and specialization of LLM components to handle instruction interpretation, function mapping, and robust code generation.
5 3D Tasks with VLMs
While Section 4 discussed methods integrating LLMs in 3D tasks, a vast body of research has explored various aspects of 3D understanding through the lens of 2D Vision-Language Models (VLMs). VLMs include much richer visual information that can be linked directly to 3D. This section reviews contributions from a series of recent papers covering language-driven open-world understanding, instance-level understanding, unified end-to-end architectures, spatial reasoning, generation, and beyond.
5.1 Open-Vocabulary 3D Scene Understanding
Open-vocabulary 3D scene understanding aims to recognize and describe scene elements using natural language descriptions instead of predefined category labels. OpenScene [29] takes a zero-shot approach by predicting dense features for 3D scene points that are co-embedded with CLIP’s text and image pixel embeddings in a shared feature space, enabling task-agnostic training and open-vocabulary querying to identify objects, materials, affordances, activities, and room types. CLIP-FO3D [304] follows a similar approach, modifying CLIP to extract dense pixel features from 3D scenes which are projected to point clouds, then training a 3D model via distillation to transfer CLIP’s knowledge. Semantic Abstraction [305] extracts relevancy maps from CLIP as abstracted object representations to generalize to novel semantics, vocabularies, and domains. Open-Fusion [306] combines the SEEM [307] vision-language model with TSDF 3D mapping for real-time open-vocabulary scene creation and querying, utilizing region-based embeddings and confidence maps.
Approaches such as PLA [308] and RegionPLC [309] leverage contrastive learning to combine captions with 2D and 3D data modalities to associate visual and semantic information. PLA [308] uses 3D-caption pairs and contrastive learning to associate multi-view images with captions for learning visual-semantic representations, while RegionPLC [309] proposes region-aware contrastive learning by combining region-level captions from 2D models mapped to 3D points. OVIR-3D [310] fuses 2D region proposals and text-aligned features from an off-the-shelf 2D detector into 3D instances for efficient open-vocab retrieval. CoDA [311] uses 3D geometry priors from annotated base categories and CLIP’s 2D semantic priors in its 3D Novel Object Discovery (3D-NOD) strategy. Its Discovery-driven Cross-Modal Alignment (DCMA) aligns 3D and image/text features for novel object localization and classification.
Instance-level scene understanding works such as OpenMask3D [312] and Open3DIS [313] utilize predicted class-agnostic 3D instance masks and 2D segment-level CLIP embeddings to enable open-vocabulary 3D instance segmentation. OpenIns3D [314] achieves open-vocabulary understanding without aligned images, using a “Mask-Snap-Lookup” pipeline that predicts 3D mask proposals, generates synthetic scene images, and assigns categories to masks via a language module. Rozenberszki et al.[315] propose to leverage CLIP features to ground 3D feature learning for 3D semantic and instance segmentation.
Language grounding with NeRFs has shown promising results in open-vocabulary scene understanding. Several methods, such as DFF [316], LERF [30], VL-Fields [317], and 3D-OVS [318], distill the knowledge of 2D feature extractors like DINO or CLIP into 3D feature fields by minimizing the error of volume-rendered features with respect to the 2D features, enabling query-based local editing and grounding language into neural implicit representations. LERF [30] optimizes a dense, scale-conditioned 3D language field by volume rendering CLIP embeddings. LangSplat [319] and N2F2 [320] demonstrate efficient open-vocabulary querying and interaction within a 3D Gaussian Splatting representation by leveraging hierarchical supervision and multiscale feature fields.
5.2 Text-Driven 3D Generation
Section 4.6 covered methods that used LLMs for 3D generation. Here, we survey text-to-3D generation methods that leverage guidance from 2D VLMs [174] and text-to-image diffusion models [321, 322] using differentiable rendering. Early works such as DreamFields [323], CLIP-Mesh [32], CLIP-Forge [324] and Text2Mesh [325] explore zero-shot 3D generation guided by CLIP.
DreamFusion [31] introduced Score Distillation Sampling (SDS) where the parameters of a 3D representation are optimized by making its renderings from arbitrary perspectives appear highly realistic, as evaluated by a pre-trained 2D diffusion model. It uses the text-to-image Imagen model [322] to optimize a NeRF representation via SDS. Magic3D [326] proposes a two-stage framework: generating a coarse model with a low-resolution diffusion prior and sparse 3D hash grid, then optimizing a textured 3D mesh model with an efficient differentiable renderer and high-resolution latent diffusion model [321]. Fantasia3D [327] disentangles geometry and appearance, using a hybrid DMTet [328] representation and spatially varying BRDFs. ProlificDreamer [329] introduced Variational Score Distillation (VSD), a particle-based framework treating 3D parameters as random variables to improve fidelity and diversity. Dream3D [330] utilizes explicit 3D shape priors and text-to-image diffusion models for enhanced text-guided 3D synthesis. MVDream [331] employs a multi-view consistent diffusion model trainable on few-shot data for personalized generation. Text2NeRF [332] combines NeRF representations with pre-trained text-to-image diffusion models to generate diverse indoor/outdoor 3D scenes from language. Apart from simultaneously generating geometries and appearance, several studies explore the possibility of solely synthesizing the textures based on given geometries [333, 334, 335].
For human avatars, AvatarCraft [336] uses diffusion models to guide neural implicit field geometry/texture learning from text prompts. Additionally, it enables animating these human avatars by deforming the neural implicit field with an explicit warping field that maps the target human mesh to a template human mesh. AvatarCLIP [337] proposes a zero-shot CLIP-supervised framework for 3D avatar generation, geometry sculpting, texture mapping and motion synthesis from text. CG-HOI [338] uses diffusion models to characterize dynamic human-object interactions from text. GenZI [339] distills information about human interactions via a pre-trained vision-language model, producing zero-shot 3D human-scene interaction synthesis from text prompts.
Exploring compositional generation, CG3D [340] generates scalable 3D scenes using explicit 3D Gaussian radiance fields by compositing individual objects without bounding boxes. Po et al. [341] introduce locally conditioned diffusion for granular scene control via text prompts and bounding boxes. GraphDreamer [342] generates compositional scenes from scene graphs by decomposing them into global-local descriptions to optimize object SDFs.
Overall, these approaches combine diffusion models, vision-language models, neural representations, and 3D priors for text-to-3D generation of objects, avatars, and scenes.
5.3 End-to-End Architectures for 3D Vision & Language
Transformer models pre-trained on large 3D-text datasets learn powerful joint representations that bridge the visual and language modalities. 3D-VisTA [343] is a transformer model that employs self-attention to jointly model 3D visual and text data, enabling effective pre-training on objectives like masked language/object modeling and scene-text matching. UniT3D [227] takes a unified transformer approach, combining a PointGroup 3D detection backbone, BERT text encoder, and multi-modal fusion module, with joint pre-training on synthetically generated 3D-language data. SpatialVLM [344] takes a different tack, and co-trains VLMs on a large synthetic 3D spatial reasoning dataset, boosting performance on 3D spatial visual question answering tasks and enabling applications like chain-of-thought reasoning for robotics. Multi-CLIP [345] pre-trains a 3D scene encoder to align scene features with CLIP’s text and image embeddings, aiming to transfer CLIP’s knowledge for improved 3D understanding on tasks like visual question answering.
In addition to pre-training approaches, researchers have explored architectures that unify 3D perception with language capabilities in an end-to-end framework. D3Net [27] combines dense captioning and visual grounding with a 3D object detector, a speaker for generating captions from detections, and a listener for discriminating objects using the captions. Uni3DL [28] operates on point clouds, with modules for text encoding, point encoding, semantic/mask prediction, and diverse task outputs like segmentation, detection, grounding, and captioning. InstanceRefer [346] uses panoptic segmentation and linguistic cues to filter instance candidates based on language descriptions for visual grounding tasks in 3D point clouds, while LanguageRefer [347] combines language embeddings with spatial embeddings from 3D bounding boxes. 3DVG-Transformer [348] also tackles 3D grounding in point clouds, with a coordinate-guided contextual aggregation module and multiplex attention for effective feature fusion.
Dataset     Object Captioning     Scene Captioning     Dense Captioning     Single Object Grounding     Multiple Object Grounding     Question Answering     Situated Question Answering     Dialogue     Task Planning     Real     Synthetic     Object     Indoor     Outdoor     Human-Annotated     Model-Annotated     Template-Based
Cap3D [349]
Text2Shape [350]
SceneVerse [351]
nu-Caption [271]
nu-Grounding [271]
ScanRefer [218]
ReferIt3D [228]
Multi3DRefer [230]
Chat-3D v2 [172]
EmbodiedScan [352]
ScanEnts3D [353]
WildRefer [354]
RIORefer [355]
ARKitSceneRefer [356]
ScanERU [229]
DenseGrounding [231]
ScanQA (Azuma et al.[234]
ScanQA (Ye et al.[357]
3DMV-VQA [236]
NuScenes-QA [235]
CLEVR3D [358]
SQA-3D [233]
3D-LLM [153]
ScanScribe [294]
M3DBench [359]
GPT4Point [268]
LAMM [360]
TABLE III: An overview of datasets used for 3D-Related tasks using Large Language Models. For each dataset, we show which tasks that dataset has been used to demonstrate, whether the data is captured from the real world or synthetically generated, whether the 3D data is of an object, an indoor scene or an outdoor scene, and how the annotations are obtained. We focus primarily on newer datasets which are used to evaluate methods in recent research papers.
6 Datasets
We now provide a high-level overview of the datasets that are used to train and evaluate 3D vision-language models. In Table III, we provide a list of datasets alongside the tasks they are used for, as well as information about the 3D scans and annotations. In Fig. 4, we present these datasets on a timeline, showing where each dataset sourced 3D information from. Current 3D vision-language datasets are almost exclusively generated by taking existing 3D vision datasets, and applying human, model or templated annotations to samples. As seen in Table III, a majority of existing datasets focus on real, indoor scenes, and this is partially explained by observing that the majority of existing datasets use 3D scans from ScanNet [361] and 3RScan [295]. Many of the datasets presented here share the same 3D data, and instead differ primarily through their choice of annotation strategy, and the 3D vision-language task they are designed to be used for.
Datasets for 3D navigation and manipulation using language are often designed around specific requirements and have a large overlap with existing bodies of research. We refer readers to existing survey papers [362, 363] for an overview of these datasets. Similarly, for text-to-3D generation datasets, we direct readers to the recent survey by Lee et al. [243]. We omit further discussion here due to prior extensive coverage, and because many methods use 2D vision-language data rather than 3D-specific datasets.
Figure 4: Timeline of datasets. A timeline showing how existing datasets are combined and annotated to form new datasets for 3D vision language tasks. Datasets in orange are foundational 3D datasets without language annotations and datasets in blue are the annotated datasets used in 3D vision language tasks. Many existing datasets use 3D data from the same sources, such as ScanNet and 3RScan, and instead differ primarily through their choice of annotation strategy and targeted 3D vision task. Note that WildRefer also introduces new 3D data and annotations for vision-language tasks.
Cap3D [349] is a 3D object captioning dataset developed upon 660k objects from the Objaverse [293] dataset. It is constructed by generating 2D image captions from multiple views of the 3D object, and consolidating those captions using image-text alignment and LLMs.
Text2Shape [350] is a human-captioned form of the 8,447 tables and 6,591 chairs found in ShapeNet [296], combined with a procedurally generated dataset of primitive shapes labeled with template-based captions. It was originally used for generative text-to-3D shape tasks.
SceneVerse [351] is a large-scale, multi-purpose dataset of annotated scenes made by compiling 68k scenes from existing 3D datasets. SceneVerse contains a total of 2.5 million vision-language pairs for object captioning, scene captioning and generating relative descriptions, primarily generated through the use of 3D scene graphs and LLMs.
nu-Caption [271] is a captioned version of 420k LiDAR scans from the nuScenes [364] dataset, annotated using GPT-4 and 2D MLLMs. Captions include general scene descriptions, detailed descriptions of objects and their relationships, and the identification of potential risks on the road.
nu-Grounding [271] builds on nu-Caption by focusing on the task of grounding, using the annotations from nuScenes to to create 280k pairs of questions and answers for visual grounding and grounded captioning.
ScanRefer [218] introduces the task of 3D RGB-D grounding using a natural language expression by creating 51,583 human-annotated ‘referring expressions’ of 11,046 objects found in 800 scenes of ScanNet [361]. The input consists of a point cloud of a scanned 3D scene alongside a free-form description of a specified target object, and the output is the corresponding bounding box of the object. ScanRefer provides an evaluation server and online benchmark  [^2] https://kaldir.vc.in.tum.de/scanrefer_benchmark/ to enable comparison between different methods.
ReferIt3D [228] introduces a number of datasets (Nr3D, Sr3D and Sr3D+) consisting of objects from 707 ScanNet scenes. Similarly to ScanRefer, these object have been annotated with referring expressions, with a focus on queries where the scene contains multiple instances of the target class, and the referring expression is required to disambiguate between them. Nr3D contains 41,503 human-annotated, free-form utterances to refer to objects in 3D scenes, Sr3D contains 83,572 template-based utterances, and SR3D+ is a version of Sr3D with augmented utterances. ReferIt3D also provides an evaluation server and online benchmark  [^3] https://referit3d.github.io/benchmarks.html to enable comparison between different methods.
Multi3DRefer [230] is a modified version of the ScanRefer dataset. Instead of referring expressions that always refer to one object in the scene, Multi3DRefer contains 6688 zero-target, 42,060 single target and 13178 multiple target referring descriptions collected for 11,609 objects in 800 ScanNet scenes. ChatGPT [286] is also used to rephrase the referring expressions.
Chat-3D v2 [172] is another modified form of ScanRefer, where the referring expressions from 705 scenes in ScanNet are used to construct scene captions that describe the relationships between objects in the scene. The scene captions are generated using GPT-4 [365] by providing the model with ground-truth information about objects. The generated captions contain references to explicit ‘object identifiers’ that directly represent each object in the scene.
EmbodiedScan [352] is an annotated combination of Matterport3D [285], 3RScan [295] and ScanNet [361], designed as a multi-modal, ego-centric dataset for 3D scene understanding. Segment Anything [366] and other annotation tools are used to provide 3D bounding boxes, semantic occupancy, and 970k template-based language descriptions across a total of 5185 scenes.
ScanEnts3D [353] extends both ScanRefer [218] and ReferIt3D [228] by using professional annotators to link every object mentioned in referential sentences to their respective instances within a 3D scene. In the original paper this dataset is only used for training purposes, where it is found to improve the model’s performance on other visual grounding and captioning datasets.
WildRefer [354] presents the STRefer and LifeRefer datasets, emphasizing human-centric in the wild settings with comprehensive 3D and linguistic human-annotations for 3D grounding. STRefer contains 5,458 referring expressions for objects in 662 scenes from the STCrowd [367] dataset, whereas LifeRefer contains 25,380 referring expressions for objects in 3,172 scenes from a new set of 3D scans acquired for this dataset.
RIORefer [355] is a human-annotated version of 3RScan [295], used for 3D grounding. It consists of 63k descriptions of objects across 1,380 scenes. This dataset is introduced as a way of testing the cross-dataset generalization abilities of models, such as in the proposed “ScanRefer to RIORefer generalization” and “RIORefer to ScanRefer generalization” tasks.
ARKitSceneRefer [356] is an annotated version of ARKitScenes [368] emphasizing the 3D grounding of small, everyday objects found in real-world indoor environments. It contains 15k descriptions of objects found in 1,605 scenes.
ScanERU [229] is a modified and human-annotated version of ScanNet that combines 46k referring expressions from ScanRefer with 706 ScanNet scenes that have been modified to include a 3D human model gesturing to the referred object, using locations designated by human annotators.
DenseGrounding [231], like Multi3DRefer, seeks to extend the task of 3D grounding to include multiple objects, however instead of a single referring expression referring to multiple objects, each input is a combined paragraph of referring queries which each correspond to a single object. These paragraphs are constructed using the nearest neighbors of random objects from ScanRefer and ReferIt3D, and combining their referring expressions together to form a paragraph.
ScanQA (Azuma et al.) [234] is an annotated version of ScanNet with 41k question-answer pairs generated across 800 scenes. Questions are automatically generated using the referring expressions in ScanRefer and then refined by human annotators, whereas answers are derived entirely from human-annotators. This is the dataset that is typically referred to by the name “ScanQA”.
ScanQA (Ye et al.) [357] was released contemporaneously with ScanQA (Azuma et al.) and is also a human-annotated form of ScanNet to be used as a 3D question answering dataset. Ye et al.contains 10k question-answer pairs for 806 ScanNet scenes. While Azuma et al.initially generates questions using referring expressions from ScanRefer, Ye et al.use human annotators for the entirety of question creation.
3DMV-VQA [236] is an annotated version of 5k scenes from the Habitat-Matterport 3D Dataset (HM3D) [369], using the semantic information from HM3DSem [370] to generate 50k questions of four types: “concept”, “counting”, “relation” and “comparison”. These questions are generated as templates, and then transformed into natural language questions.
NuScenes-QA [235] consists of 34k scenes from nuScenes [364] annotated with 460k template-style question-answer pairs generated using a constructed scene graph. Questions are divided into 5 types: “existence”, “counting”, “query-object”, “query-status” and “comparison”, and can include spatial reasoning.
CLEVR3D [358] is an annotated form of 3RScan [295] designed for indoor 3D QA tasks. Template-based questions and answers are generated using the scene graph annotations from 3DSSG [371]. Initially, 44k questions are generated for 1,333 scenes, however they use a “compositional scene manipulation” technique to randomly replace objects in the scene graphs with objects from a generated object pool, artificially leading to 171k questions across 8,771 scenes.
SQA-3D [233] is another human-annotated version of ScanNet, proposing the problem of “situated question answering”, where each query consists of a description of the agent’s current location and situation, and a query relating to the scene. SQA-3D contains 33.4k questions across 650 scenes.
3D-LLM [153] is an annotated version of Objaverse, ScanNet and HM3D/HM3DSem designed to handle a broad variety of tasks involving 3D vision and language. BLIP-2 [195] and ChatGPT [286] are used to generate 300k instances of ‘3D-language data’ that are used across a range of 3D vision-language tasks.
ScanScribe [294] is an annotated version 1,185 scenes from ScanNet and 3RScan designed to act as a large-scale 3D scene-text pair dataset for pre-training 3D vision-language models. It utilizes ScanQA, ScanRefer and ReferIt3D for ScanNet data, and uses GPT-3 prompting to generate annotations for 3RScan to create a total of 278k scene descriptions. To increase the diversity of the dataset, scenes are synthetically generated by random replacing 10% of the objects in scenes with a same-category object from Objaverse.
M3DBench [359] is a multi-modal instruction-following dataset consisting of 327k instruction-response pairs across a broad range of tasks. Data is collected from a large set of existing datasets (see Fig. 4), alongside using GPT-prompting to generate additional annotations.
GPT4Point [268] is an annotated form of over 1 million objects from Objaverse-XL, designed for 3D Captioning, 3D QA and other 3D tasks. Annotations are generated automatically through a hierarchical pipeline that fuses and improves captions from multiple views.
LAMM [360] is a large scale multi-modal instruction-tuning dataset that covers tasks for both 2D and 3D vision. The 3D Data is sourced from 3RScan [295], CLEVR3D [358], 3DSSG [371] and ShapeNet [296], and is annotated using GPT [365] and template-based responses.
7 Challenges and Opportunities
Despite progress in integrating LLMs with 3D data, challenges remain in data representation, computational efficiency, and benchmarks, necessitating innovative solutions.
Representation Choice has a large impact on the performance of 3D vision language models. Currently, point clouds are predominantly used to represent both indoor (e.g. vertices of meshes) and outdoor (e.g. LiDAR point clouds) environments due to their simplicity and neural network compatibility. However, they struggle to capture the fine details crucial for accurate, rich spatial models. Developing novel 3D scene representations that more effectively bridge the gap between spatial information and language could unlock new levels of understanding and interaction. By finding innovative ways to encode language and semantic information in 3D representations, such as using distilled language and semantic embeddings [30, 316], could help bridge the gap between these two modalities.
Computational Demands of both 3D data processing and LLMs pose significant challenges. Scalability remains a concern as the complexity of 3D environments and the size of language models increase. Advances in LLM architectures designed for adaptivity and computational efficiency could significantly broaden their application scope.
Improving Benchmarks is essential for comprehensively evaluating and advancing the capabilities of multi-modal LLMs in 3D tasks. Current benchmarks have limited scope, especially regarding 3D reasoning, hampering evaluation of spatial reasoning skills as well as development of 3D decision-making/interaction systems. Moreover, the metrics in use today fall short of capturing the full spectrum of LLMs’ capabilities in 3D environments. Crafting task-specific metrics that more precisely measure performance across diverse 3D tasks is essential. Finally, the granularity of current scene understanding benchmarks is overly simple, limiting insight into complex 3D environment comprehension. A more diverse array of tasks is required.
Safety and Ethical Implications are imperative to consider when employing LLMs for 3D understanding. LLMs can hallucinate and output inaccurate, unsafe information leading to erroneous decisions in critical 3D applications. Moreover, LLMs often fail in ways that are unpredictable and challenging to interpret. They may also inherit societal biases present in training data, disproportionately disadvantaging certain groups when making predictions in real-world 3D scenes. It is crucial to approach the use of LLMs in 3D contexts with caution, employing strategies to create more inclusive datasets, robust evaluation frameworks for bias detection and correction, and mechanisms to minimize hallucinations, ensuring responsible and equitable outcomes.
8 Conclusion
This survey paper provides a thorough exploration of the integration of LLMs with 3D data. Systematically reviewing methodologies, applications, and emergent abilities of LLMs in processing, understanding, and generating 3D data, the survey underlines the transformative potential of LLMs across a spectrum of 3D tasks. From enhancing spatial comprehension and interaction within 3D environments to driving forward the capabilities of embodied AI systems, LLMs emerge as pivotal in advancing the field.
Key findings include the identification of LLMs’ unique advantages such as zero-shot learning, advanced reasoning, and extensive world knowledge, which are instrumental in bridging the gap between textual information and spatial interpretation. The paper showcases the wide array of tasks where LLMs’ integration with 3D data has been successfully demonstrated. The exploration of other 3D vision-language methods alongside LLMs reveals a rich landscape of research aiming to deepen our understanding of the 3D world.
Furthermore, the survey highlights significant challenges such as data representation, model scalability, and computational efficiency, suggesting that overcoming these hurdles is crucial for the full realization of LLMs’ potential in 3D applications. In conclusion, this survey not only offers a comprehensive overview of the current state of 3D tasks using LLMs but also sets the stage for future research directions. It calls for a collaborative effort to explore and expand the capabilities of LLMs in understanding and interacting with the complex 3D world, paving the way for further advancements in the area of spatial intelligence.

### v2

@@
-captionUnknown document class (or package)
+When LLMs step into the 3D World: A Survey and Meta-Analysis of 3D Tasks via Multi-modal Large Language Models
 Abstract
 As large language models (LLMs) evolve, their integration with 3D spatial data (3D-LLMs) has seen rapid progress, offering unprecedented capabilities for understanding and interacting with physical spaces.
-This survey provides a comprehensive overview of the methodologies enabling LLMs to process, understand, and generate 3D data.
+This survey provides a comprehensive overview of the methodologies that enable LLMs to process, understand, and generate 3D data.
 Highlighting the unique advantages of LLMs, such as in-context learning, step-by-step reasoning, open-vocabulary capabilities, and extensive world knowledge, we underscore their potential to significantly advance spatial comprehension and interaction within embodied Artificial Intelligence (AI) systems.
@@
 1 Introduction
-The advent of Large Language Models (LLMs) has marked a transformative era in natural language processing, enabling machines to understand, generate, and interact with human language in previously unimagined ways.
-However, the physical world around us is inherently three-dimensional, and understanding spatial 3D environments is crucial for many real-world applications that involve perception, navigation, and interaction within these 3D spaces.
-With recent advancements, the application of LLMs has extended well beyond text.
-The fusion of LLMs with 3D data presents a unique opportunity to enhance computational models’ understanding of and interaction with the physical world, leading to innovations across various domains, including autonomous systems [1, 2, 3, 4, 5], augmented reality [6, 7, 8, 9], robotic navigation [10, 11, 12], and robotic manipulation [13, 14, 15].
+The advent of Large Language Models (LLMs) has marked a transformative era in natural language processing, enabling machines to understand, generate, and interact with human language in ways previously unimaginable.
+However, the physical world around us is inherently three-dimensional, and understanding spatial environments is crucial for many real-world applications that involve perception, navigation, and interaction within these 3D spaces.
+With recent advances, the application of LLMs has extended well beyond text.
+The fusion of LLMs with 3D data presents a unique opportunity to enhance computational models’ understanding of, and interaction with the physical world, leading to innovations across various domains, including autonomous systems [1, 2, 3, 4, 5], augmented reality [6, 7, 8, 9], robotic navigation [10, 11, 12], and robotic manipulation [13, 14, 15].
 Recent works have demonstrated the potential of integrating LLMs with 3D data to interpret, reason, or plan in complex 3D environments, by leveraging the inherent strengths of LLMs, including zero-shot learning [16, 17], advanced reasoning [13, 18, 19], and extensive knowledge [20, 21].
-However, the integration of LLMs with 3D data is not trivial.
-Issues such as 3D data representation, model scalability, and computational efficiency remain significant hurdles.
-Moreover, ensuring the models can operate in real-world settings requires overcoming obstacles related to data diversity and environmental complexity.
-Addressing these challenges is crucial for the full realization of LLMs’ potential in 3D applications to create dynamic and context-aware AI systems.
-This survey paper provides a critical examination of the intersection of LLMs with 3D data, offering an exhaustive overview of current methodologies, applications, and challenges in this domain.
-We start by providing the relevant background on common 3D representations, a brief introduction to LLMs and an overview of Vision-Language Models (VLMs) and Vision Foundation Models (VFMs) in Sec.
-2.
-In Sec.
-3, we detail the 3D vision language tasks that current methods aim to solve, outlining the current evaluation metrics and protocols.
-Figure 1: A general timeline of methods leading up to Large Language Models (LLMs) in 3D.
-Next, in Sec.
-4, we analyze the data formats, processing techniques, and model architectures that have shown promise in enhancing 3D understanding through the lens of LLM capabilities.
-We showcase a variety of areas where the amalgamation of LLMs and 3D data has been successfully demonstrated, such as: using LLMs’ world-knowledge [20, 22] and reasoning capabilities [21, 23] to enhance 3D task performance, using LLMs as multi-modal interfaces [24, 18] and embodied agents [15, 13], or generating complex scenes with LLMs [25, 26].
-Beyond LLMs, a few research works have proposed end-to-end architectures that unify 3D perception with language capabilities [27, 28].
-Moreover, a wide range of works have explored distilling the knowledge of off-the-shelf 2D Vision Language Models (VLMs) for open vocabulary 3D scene understanding [29, 30] as well as text-driven 3D generation [31, 32].
-This survey provides a complete overview of these methods in Sec.
-5 to show a complete picture of the field of 3D+Language.
-We then outline the datasets used to train and evaluate these methods in Sec.
-6.
-Finally, Sec.
-7 highlights the challenges and potential future research directions in this field.
+However, despite rapid advances in this area, research remains scattered and relatively nascent, with most efforts emerging only since 2023 (see Fig.
+1).
+This fragmentation creates a need for a comprehensive and systematic survey to consolidate the current state of research and provide a clear overview of the methodologies, challenges, and future directions.
+The aim of this survey is to fill these gaps by providing a detailed analysis of the field, enabling researchers to better understand the progress made thus far and to identify key areas for future exploration.
+To the best of our knowledge, we are the first to survey the field of 3D-LLMs.
+We outline the methodologies that are currently being used to allow LLMs to ingest and process 3D information in Fig.
+3, and present a novel taxonomy categorizing 3D-LLM methods based on the role of the language model in solving 3D tasks in Fig.
+4.
+We describe areas where the amalgamation of LLMs and 3D data has been successfully demonstrated, such as: using LLMs’ world-knowledge [20, 22] and reasoning capabilities [21, 23] to enhance 3D task performance, using LLMs as multi-modal interfaces [24, 18] and embodied agents [15, 13], or generating complex 3D scenes with LLMs [25, 26].
+We then provide a multifaceted comparison between existing methods, with a focus on the 3D vision components, LLM components, and alignment methodologies used to train these models (see Tab.
+I).
+Figure 1: An overall timeline to show the developments of 3D representations (Sec.
+2.1), L(V)LMs (Sec.
+2.2 and Sec.
+2.3), 3D w/VLMs&VFMs(Sec.
+5), 3D w/LLMs (Sec.
+4).
+This figure illustrates how the development of 3D representation and L(V)LMs inspired the 3D vision-language methods (3D w/LLMs, 3D w/VLMs VFMs).
+Besides, we can clearly see the rapid growth of 3D-LLM methods starting from 2023, which calls for the attention of more researchers.
+Compared to other LLM-related surveys, which either focus on single tasks like video understanding [27]—or on specific scenarios like robotics [28, 29, 30] or autonomous driving [31], our survey focuses on the broad potential of integrating LLMs with 3D data.
+Based on this, we highlight how the advantages and disadvantages of different 3D representations (Sec.
+2.1), and how these representations can be combined with LLMs to accomplish various 3D tasks (Sec.
+4).
+Additionally, we cover 3D methods utilizing Vision-Language models (VLMs) and Vision-Foundation Models (VFMs) (Sec.
+5).
+Altogether, these solutions could provide insights for various real-world scenarios such as AR/VR, healthcare, game development, and interior design.
+In addition to the above, we offer a thorough review of 3D vision-language tasks (Sec.
+3), categorizing them into five major types based on differences in the input and output modalities.
+We also compare datasets used for training and testing (Sec.
+6), mapping the trajectory of this field through the evolution of datasets (see Fig.
+5).
+Finally, we identify key challenges and outline future directions (Sec.
+7).
 2 Background
-This section offers essential background knowledge on 3D representations, Large Language Models (LLMs), 2D Vision-Language Models (VLMs), and Vision Foundation Models (VFMs).
+This section offers background knowledge on 3D representations, Large Language Models (LLMs), 2D Vision-Language Models (VLMs), and Vision Foundation Models (VFMs).
 2.1 3D Representations
-Choosing a 3D representation to describe, model, and understand our world is a crucial topic that contributes to the understanding of the current progress in 3D-LLMs.
-It is also a fundamental research area in computer vision.
-The field has recently seen substantial growth due to advancements in deep learning, computational resources, and 3D data availability.
-We briefly describe the most common 3D representations currently in use.
-Point Clouds represent 3D shapes with a set of data points in space, storing each point position in a 3D Cartesian coordinate system.
-In addition to storing the position, other information (e.g.
-color, normal) can also be stored with each point.
-Point cloud-based methods [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] are noted for their low storage footprint but lack surface topology information.
-Typical sources for obtaining point clouds include LiDAR sensors, structured light scanners, time-of-flight cameras, stereo-views, photogrammetry, etc.
-Voxel Grids [44, 45, 46, 47, 48, 49, 50, 51] consist of unit cubes in 3D space, akin to pixel representations in 2D [52].
-Each voxel minimally encodes the occupancy information (binary or probabilistically) but can additionally encode distance to a surface, as in the Signed Distance Function (SDF) [53, 54, 55, 56, 57] or Truncated Signed Distance Function (TSDF) [58, 59, 60, 61, 62, 63].
-Nevertheless, the memory footprint can become excessive when high-resolution details are required.
-Polygonal Mesh representations are composed of vertices and surfaces and describe complex 3D shapes compactly.
-However, their unstructured and non-differentiable nature [64] poses challenges in integrating them with neural networks for end-to-end differentiable pipelines.
-Some solutions addressing this problem, such as gradient approximation-based approaches [65, 66, 67, 68], can only use handcrafted gradient calculation.
-Other solutions like differentiable rasterizers [69, 70, 71] may lead to imprecise render results such as blurry content.
-Neural Fields [72] have gained a growing interest in the 3D research community over recent years, departing from traditional representations reliant on geometric primitives [73].
-A neural field is a mapping from spatial coordinates to scene properties (such as occupancy, color, radiance, etc), but unlike a voxel grid – in which the mapping is from a discrete cell to the value at that voxel – in a neural field the mapping is a learned function, typically a multi-layer perceptron.
-In this way neural fields implicitly learn compact, continuous, and differentiable 3D shape and scene representations.
-One set of neural fields has focused on implicit surface representations.
-Occupancy Networks [74, 75, 76] encode shapes in a continuous 3D occupancy function represented by neural networks, using 3D point locations and features from point clouds, low-resolution voxels, or images to estimate occupancy probabilities.
-Concurrently, Deep SDF networks [77, 78, 79, 80] use neural networks to estimate SDFs from 3D coordinates and latent vectors.
-Recent methods like NeuS [79] and NeuS2 [80] notably improve surface reconstruction fidelity and efficiency for both static and dynamic objects.
-Another set of approaches, referred to as Neural Radiance Fields (NeRF) [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94], display strong photorealistic rendering capabilities of the 3D world.
-These methods encode scene details with positional encoding techniques [95, 96, 90, 97], and utilize MLPs to predict radiance values (color and opacity) along camera rays.
-However, the necessity for MLPs to infer color and occupancy details for each sampled point in space (including those in empty spaces) requires significant computational resources.
-Thus, there is a strong incentive to reduce NeRFs’ computational overhead for real-time applications.
-Hybrid Representations attempt to combine NeRF techniques with traditional volumetric-based approaches, facilitating high quality, real-time rendering [98, 99, 100, 101, 102, 103, 104, 105, 106, 107].
-For instance, combining voxel grids [95, 96, 98, 97] or multi-resolution hash grids [102] with neural networks greatly reduces the NeRF’s training and inference time.
-3D Gaussian Splatting [103, 106, 107] is a variation on point clouds where each point contains additional information representing the radiance being emitted in the spatial region around that point as an anisotropic 3D Gaussian ‘blob’.
-These 3D Gaussians are typically initialized from SfM point clouds [108], and optimized using differentiable rendering.
-3D Gaussian Splatting achieves state-of-the-art novel-view synthesis at a fraction of NeRF’s computation by utilizing efficient rasterization [109] rather than raytracing.
-2.2 Large Language Model (LLM)
-Traditional natural language processing (NLP) encompasses a wide range of tasks aimed at enabling systems to understand, generate and manipulate text.
-Early approaches in NLP relied on techniques like rule-based systems, statistical models, and earlier neural architectures like recurrent neural networks [110, 111].
-The recent introduction of Large Language Models (LLMs), employing transformer architectures [95] and training on vast text corpora [112], has achieved unprecedented performance and sparked a new wave of excitement in the field.
-Since this paper focuses on 3D LLMs, we provide relevant background knowledge of LLMs here.
-For an in-depth exploration of LLMs at large, we refer to the latest surveys in the area [113, 114, 115, 116, 117].
+Choosing an effective 3D representation is a critical design decision in 3D-LLMs research.
+It is also a fundamental research area in computer vision and graphics, driven by advances in deep learning, computational resources, and the availability of 3D data.
+In practice, image-based representations, such as multi-view RGB images and RGB-D data, although not strictly 3D, are often employed as proxies for 3D geometry, owing to their ease of acquisition and strong compatibility with existing 2D vision backbones.
+Nevertheless, these representations mainly capture geometry at the pixel level.
+Conversely, structured scene representations, such as 3D scene graphs, offer semantic abstraction over spatial layouts and are increasingly adopted for reasoning tasks in embodied AI.
+Below, we briefly describe the most common 3D representations.
+Point Clouds represent 3D shapes as discrete data points in space, typically storing the XYZ coordinates and optionally additional attributes such as color or surface normals.
+Point cloud-based methods [32, 33] are lightweight and easy to obtain (e.g., from LiDAR, structured light, or photogrammetry), but lack explicit surface topology.
+Voxel Grids [34, 35] divide the 3D space into regular volumetric units (voxels), analogous to pixels in 2D.
+Each voxel can encode occupancy (binary or probabilistic), surface distance (e.g.
+Signed Distance Function [36, 37]), or Truncated SDF [38, 39].
+While intuitive, voxel grids become memory-intensive at high resolution, which motivates sparse voxel structures [40].
+Meshes use interconnected vertices, edges, and faces to compactly model 3D geometry with surface continuity.
+Despite their expressiveness, they are less favored in learning-based pipelines due to their irregular topology.
+Solutions include gradient approximation [41, 42] and differentiable rasterization [43], though these often suffer from rendering artifacts or limited precision.
+Neural Fields [44] represent 3D scenes as continuous functions (typically MLPs) that map spatial coordinates to physical properties such as occupancy, color, or semantics.
+Unlike discrete voxel grids, neural fields model scenes as continuous volumes.
+For example, Occupancy Networks [45, 46] predict whether a point lies inside an object, while Deep SDF [47, 48] learns continuous signed distance fields.
+These representations are compact and flexible but often slow to evaluate.
+Neural Radiance Fields (NeRF) [49, 50] extend neural fields by modeling view-dependent color and density along camera rays.
+This enables photorealistic novel-view synthesis but requires expensive sampling and evaluation, especially in empty space, prompting research on acceleration.
+Hybrid Representations combine neural fields with explicit structures to improve efficiency.
+For instance, voxel grids [51] or multi-resolution hash grids [52] are used to condition lightweight MLPs, significantly reducing NeRF training and inference cost while maintaining high rendering quality.
+3D Gaussian Splatting [53] offers a differentiable point-based representation, where each point emits anisotropic Gaussian blobs instead of sharp vertices.
+This enables real-time, high-fidelity rendering.
+Initializations typically come from SfM-based point clouds [54], though recent works [55, 56] explore SfM-free pipelines.
+Figure 2: Common task categories for 3D Vision Language models.
+Each category contains many types of queries, but we provide a single example from each category.
+The scan is from the ScanNet dataset [57], and the ‘question answering’ and ‘situated question answering’ examples are adapted from ScanQA [58] and SQA3D [59] respectively.
+2.2 Large Language Models (LLMs)
+Natural Language Processing (NLP) aims to enable systems to understand, generate, and manipulate text.
+Early approaches used rule-based systems, statistical models, or neural architectures like recurrent neural nets [60].
+The rise of LLMs, powered by transformers [61] and vast text corpora [62], has transformed the field.
+Since this paper focuses on 3D LLMs, we provide a brief background on LLMs here.
+For a more detailed coverage, see recent surveys on this topic [63, 64, 65].
 2.2.1 LLM Architectures
-In the context of LLMs, the “encoder-decoder” and “decoder-only” architectures are prominently used for NLP tasks.
-Encoder-decoder architectures [95, 118, 119] consist of two main components: an encoder $f_{enc}$ and a decoder $f_{dec}$.
-The encoder and decoder components are typically implemented using transformers [95], which employ attention mechanisms to capture long-range dependencies in the input and output sequences.
-The encoder takes the input sequence $X=(x_{1},x_{2},\ldots,x_{N})$ and maps it into a sequence of latent representations $H=(h_{1},h_{2},\ldots,h_{N})$ that capture the contextual information, and the decoder generates the output sequence $Y=(y_{1},y_{2},\ldots,y_{T})$ based on $H$.
-Mathematically, the encoding process can be expressed as $H=f_{enc}(X)$, and the entire latent sequence $H$ is generated at once from $X$.
-The decoder, however, generates the output sequence $Y$ sequentially: $y_{t}=f_{dec}(y_{<t},H)$ where $y_{<t}=(y_{1},y_{2},\ldots,y_{t-1})$.
-Decoder-only architectures [120, 121, 122], on the other hand, are a variant of the transformer architecture that uses only the decoder component.
-It is particularly suitable for language modeling tasks, where the goal is to predict the next token given the previous tokens.
-The decoder-only architecture can be mathematically expressed as $y_{t}=f_{dec}(y_{<t})$.
-Tokenization is a preprocessing method to break the input text into a sequence of tokens, the basic data unit in language models.
-The number of tokens is finite, and each token can correspond to a word, sub-word, or a single letter.
-During inference, the input text is converted to a sequence of tokens and fed to the models, which predict the output tokens that are then converted back to text.
-The tokenization has a great impact on the performance of the language models, as it affects how the models perceive text.
-Various tokenization techniques are used, such as word-level tokenization, subword tokenization (e.g.
-byte-pair encoding [123], WordPiece [124], SentencePiece [125]), and character-level tokenization [126].
+Large language models are typically built upon either encoder-decoder or decoder-only transformer architectures [61].
+Encoder-decoder models process the input sequence via an encoder to produce contextual embeddings, which are then consumed by the decoder to autoregressively generate the output.
+In contrast, decoder-only architectures [66] generate outputs directly from previous tokens without a separate encoder, making them particularly efficient for generative tasks.
+Another critical component is tokenization, which converts raw text into discrete tokens.
+Choices like byte-pair encoding (BPE) [67], WordPiece [68], and SentencePiece [69] significantly impact model performance and vocabulary expressivity.
+These architectural and preprocessing designs influence how language and multimodal inputs are represented and processed, and they are often adapted or extended in 3D-LLMs to support spatial encoders and cross-modal alignment.
 2.2.2 LLM Emergent Abilities
-One major difference between LLM and traditional non-LLM methods is the emergent abilities that become available in large models but are not present in the small models [115].
-The term “emergent abilities” refers to new, complex capabilities that arise as LLMs scale in size and complexity.
-These abilities enable advanced understanding and generation of natural language, problem-solving across various domains without specific training, and adapting to new tasks through in-context learning.
-In the following, we introduce several common emergent abilities in the scope of LLMs.
-In-Context Learning  [127, 128, 129] refers to the ability of LLMs to understand and respond to new tasks or queries based on the context provided within the prompt, without requiring explicit retraining or fine-tuning.
-Milestone papers (GPT-2/GPT-3 [130, 127]) demonstrate in-context learning in a few-shot manner, where a model is provided with a few examples of a task within the prompt and then asked to tackle a different example without prior explicit training.
-State-of-the-art LLMs, such as GPT-4 [131], exhibit remarkable in-context learning abilities, understanding complex instructions, and performing a wide range of tasks from simple translations to generating code and creative writing, all based on the context provided within the prompt.
-Reasoning in the context of LLMs, often referred to as “chain-of-thought” prompting [132, 133], involves the model generating intermediate steps or reasoning paths when tackling complex problems or questions.
-This approach allows LLMs to break tasks into smaller, manageable parts, facilitating a more structured and understandable solution process.
-To achieve this, training involves datasets that include a variety of problem-solving tasks [134, 135], logical puzzles [136, 137], and datasets designed to mimic reasoning under uncertainty [138].
-Current state-of-the-art LLMs [127, 131, 139, 140, 141, 142, 143] demonstrate advanced reasoning capabilities typically when model sizes become larger than 60B to 100B parameters [132].
-Instruction-following  [144, 127, 140, 145, 146] refers to the models’ capability to understand and execute commands, or follow instructions as specified by the user.
-This involves parsing the instruction, understanding its intent, and generating an appropriate response or action.
-Methods for adapting this ability to novel tasks may require instruction tuning from datasets [114] containing various instructions paired with correct responses or actions.
-Techniques such as supervised learning, reinforcement learning from human feedback, and interactive learning can further enhance performance.
+A defining feature of LLMs is their emergent abilities, which arise as model scale increases [65], enabling zero-shot generalization across diverse tasks without explicit training.
+These include in-context learning [70], where the model adapts to new tasks based on prompts alone; reasoning, often enabled by chain-of-thought prompting [71], where intermediate steps are generated to solve complex problems; and instruction-following [72], where the model executes natural language commands via instruction tuning.
+These abilities underpin the remarkable generality of LLMs and have directly inspired recent 3D-LLMs to adopt similar paradigms—using in-context demonstrations, compositional reasoning, and open-ended instruction-following over multimodal 3D inputs to enhance task generalization without task-specific retraining.
 2.2.3 LLM Fine-tuning
-In the context of 3D-LLMs, LLMs are either utilized directly in their pre-trained state or undergo fine-tuning to accommodate novel multi-modal tasks.
-However, fine-tuning the entirety of an LLM’s parameters presents significant computational and memory challenges due to the large amount of parameters involved.
-Therefore, parameter-efficient fine-tuning (PEFT) [147] has become increasingly popular for adapting LLMs to specific tasks by updating only a relatively small subset of the model’s parameters instead of retraining the entire model.
-The following part lists four common PEFT methods used in LLMs.
-Low-Rank Adaptation (LoRA) and variants [148, 149, 150] update parameters through low-rank matrices.
-Mathematically, a forward pass for LoRA [148] during the fine-tuning can be expressed as $h=W_{0}x+BAx$.
-$W_{0}$ is the frozen weight of LLM, while $BA$ is a low-rank matrix parameterized by newly introduced matrices $A$ and $B$, which are updated at fine-tuning stages.
-This method offers several distinct benefits.
-During fine-tuning, only $B$ and $A$ are optimized, significantly reducing the computational overhead associated with gradient calculation and parameter updates.
-Once the fine-tuning concludes and the weights are merged, there is no additional inference cost compared to the original model, as represented by the equation: $h=(W_{0}+BA)x$.
-Moreover, there’s no need to save multiple copies of LLMs for different tasks since multiple LoRA instances can be saved instead, reducing the storage footprint.
-Layer Freezing [151, 142, 152] freezes selected layers of a pre-trained model while updating others during training.
-This often applies to layers closer to the model’s input or output, depending on the task’s nature and the model architecture.
-In 3D-LLM methods, for example, all layers except input and output embeddings [153] might be frozen to mitigate overfitting risks with task-specific datasets, retaining the pre-trained general knowledge and reducing parameters that need to be optimized.
-Prompt Tuning [154, 155, 156] guides LLMs to perform specific tasks by framing them within prompts, adjusting model inputs in contrast to conventional fine-tuning which adjusts model parameters.
-Manual prompt engineering [132, 157, 158, 159] is the most intuitive method, yet it is difficult for experienced prompt-tuning engineers to find the optimal prompt.
-Another set of approaches is automatic prompt generation and optimization.
-One prevalent approach is to search for the exact optimal input prompt text, known as hard prompt, such as [160, 161, 162, 163].
-Alternatively, one can use optimization approaches to optimize the embedding of the prompt (soft prompt[154, 164, 165, 166, 167, 168, 169, 170].
-Adaptive Fine-tuning customizes the model architecture for specific tasks by adding or removing layers or modules [171, 172, 173].
-This can include integrating new data modalities, such as visual information, alongside textual data.
-The core idea of adaptive fine-tuning is leveraging small neural network modules inserted between the layers of a pre-trained model.
-During adaptive fine-tuning, only the parameters of these adapter modules are updated, while the original model weights remain fixed.
-2.3 2D Vision-Language models
-Vision-Language Models are a family of models which aim to capture and exploit the relationship between text and images/video, and are capable of performing interactive tasks between the two modalities.
-The majority of VLMs have transformer-based architectures.
-By leveraging the attention module, visual and textual contents are conditioned on each other, which enables mutual interaction.
-In the following paragraphs, we briefly cover the applications of VLMs to discriminative and generative tasks.
-Discriminative tasks involve predicting a certain characteristic about the data.
-VLMs, such as CLIP [174] and ALIGN [175], have shown exceptional performance in zero-shot transferability to unseen data in image classification.
-Both models comprise two modules: a visual encoder and a text encoder.
-Given an image and its category, CLIP and ALIGN are trained by maximizing the similarity between the image embedding and the textual embedding of the sentence “a photo of a {image category}”.
-The zero-shot transferability is achieved by replacing “{image category}” with possible candidates during inference and searching for sentences matching best with the image.
-These two works have inspired numerous follow-ups such as [176, 177, 178, 179], which further boost image classification accuracy.
-The models can also distill learned knowledge for other tasks, including object detection [180, 181, 182, 183], image segmentation [184, 185, 186, 187, 188], document understanding [189, 190] and video recognition [191].
-Generative tasks utilize VLMs to generate text or images from input data.
-By leveraging large-scale training data, a single VLM can usually perform several image-to-text generation tasks, such as image captioning and visual question answering (VQA).
-Notable examples include SimVLM [192], BLIP [193], and OFA [194], etc.
-More powerful VLMs such as BLIP-2 [195], Flamingo [196], and LLaVA [197], are able to handle multi-turn conversations and reasoning based on the input image.
-Following the introduction of the diffusion model, text-to-image generation has also become a focus of the research community [198, 199].
-By training on a large amount of image-text pairs, diffusion models [200, 201] can generate high-quality images based on textual input.
-This capability has also been extended to generating videos [202], 3D scenes [31], and dynamic 3D objects  [203].
-Along with generative tasks, it is also possible to edit existing images through the textual prompt [204, 205, 206, 207].
-2.4 Vision Foundation Models (VFMs)
-Vision Foundation Models (VFMs) are large neural networks that aim to extract image representations that are sufficiently diverse and expressive to be directly deployed in various downstream tasks, mirroring the role pre-trained LLMs serve in downstream NLP tasks.
-One notable example is DINO [208], which uses a self-supervised teacher-student training paradigm.
-The representation learned achieves good results on both image classification and semantic image matching.
-The attention weights in DINO can also serve as segmentation masks for semantic components of the observed scene.
-Followup works like iBOT [209] and DINOv2 [210] further improve the representation by introducing the masked image modeling (MIM) loss.
-SAM, a transformer-based image segmentation model [211], is trained on a dataset consisting of 1.1 billion images with semantic masks and demonstrates strong zero-shot transfer capability.
-DINO (Zhang et al.[212] – not to be confused with DINO (Caron et al.[208] – employs a DETR-like [213] architecture and mixed query selection for object detection.
-The follow-up work Grounding-DINO [214] introduces text supervision to boost accuracy.
-Stable Diffusion [201], a text-to-image generator, has also been used as a feature extractor for ‘real’ images by running a single diffusion step on clean, or artificially noised images and extracting intermediate features [215, 169] or attention masks [216].
-These features have been recently used for segmentation [216] and image matching [215, 169] tasks, due to the size and diversity of the training sets used for diffusion models, and due to observed emergent properties of diffusion features, such as zero-shot correspondence across images [215].
+Given the large number of parameters in LLMs, full fine-tuning is often infeasible for 3D-LLMs.
+Instead, parameter-efficient fine-tuning (PEFT) methods have become the standard to adapt LLMs to 3D tasks with reduced computational cost.
+LoRA [73] and its variants [74, 75] insert trainable low-rank matrices into transformer layers, allowing efficient adaptation while keeping most of the original weights frozen.
+Layer freezing [76, 77] updates only a small subset of layers, typically at the input/output, preserving general language knowledge and reducing overfitting.
+Prompt tuning steers model behavior by modifying input prompts instead of model parameters, including handcrafted prompts [71], discrete token search (hard prompts) [78], or learned embeddings (soft prompts) [79].
+Finally, adaptive tuning [80, 81] integrates lightweight neural modules between frozen layers to enable efficient cross-modal adaptation, such as incorporating 3D data alongside text.
+These PEFT strategies are widely adopted in 3D-LLMs to enable scalable and modular alignment of LLMs with 3D inputs.
+2.3 Vision-Language and Vision Foundation Models
+2D Vision and Vision-Language Foundation Models aim to learn universal visual representations and multimodal alignments, enabling transfer to a wide range of downstream tasks.
+Vision-Language Models (VLMs) such as CLIP [82] and ALIGN [83] align image and text embeddings through contrastive training, achieving strong zero-shot transfer in classification and extending to tasks like detection [84], segmentation [85], and video understanding [86].
+Later VLMs like BLIP-2 [87], Flamingo [88], and LLaVA [89] support image-conditioned generation and multi-turn interaction.
+Text-to-image diffusion models [90] further expand the generative capabilities, with extensions to video [91] and 3D generation [92, 93].
+Vision Foundation Models (VFMs) focus on learning rich visual representations via self-supervised or large-scale supervised pretraining.
+DINO [94] and its successors [95, 96] use contrastive and masked image modeling strategies, demonstrating generalization in classification, segmentation, and matching.
+Segment Anything Model (SAM)[97] enables zero-shot segmentation across diverse domains.
+Beyond this, diffusion models[90] also serve as feature extractors [98, 99], revealing semantic priors through intermediate representations.
+These 2D VLMs and VFMs lay the foundation for recent efforts extending multimodal learning into the 3D domain, which we discuss in the following sections.
 3 Tasks and Metrics
@@
 Research has grown to include a broad spectrum of research tasks, each with their own set of commonly used datasets and evaluation metrics.
-Here, we aim to list current 3D vision-language tasks, and their corresponding evaluation metrics.
-We broadly categorize the tasks by their input and output modalities.
-We then begin our analysis of methods for solving these tasks in Sec. 4 and Sec. 5. Then, in Sec. 6, we detail the datasets that are currently used for training and evaluation for these tasks.
+Here, we aim to summarize current 3D vision-language tasks, and their corresponding evaluation metrics.
+We broadly categorize the tasks by their input and output modalities, and provide an example for each category in Fig.
+2.
+We then describe methods for solving these tasks in Sec.
+4 and Sec.
+5.
+Then, in Sec.
+6, we detail the datasets that are currently used for training and evaluation for these tasks.
 3.1 3D Captioning (3D $\rightarrow$ Text)
@@
 Exact Match (EM) requires that the generated caption exactly matches the ground truth.
-Exact Match has different accuracy thresholds, denoted as EM@$K$, which means that the correct answer is within the top “$K$” answers generated by the model.
+Exact Match has different accuracy thresholds, denoted as EM@$K$, which means that the correct answer is within the top “KK” answers generated by the model.
 Commonly used thresholds are EM@1 and EM@10.
 However, natural language captions with the same semantic meaning can be represented in many ways, so the dominant metrics for captioning are automated text generation metrics [102] that aim to measure matching n-grams or semantic similarity rather than complete sentence matches.
-BLEU [220] matches n-grams between the predicted and true captions, with “BLEU@$x$” referring to matching n-grams of length “$x$” (typical values are in the range 1-4).
+BLEU [103] matches n-grams between the predicted and true captions, with “BLEU@xx” referring to matching n-grams of length “xx” (typical values are in the range 1-4).
 This still requires matching exact words, but is slightly more robust to rearrangements in phrasing.
@@
 CIDEr [106] weights n-grams by their frequency, with higher frequency n-grams given lower weights.
-As the above metrics rely on n-gram matches, they cannot account for different by semantically similar words.
+As the above metrics rely on n-gram matches, they cannot account for different but semantically similar words.
 Thus, various metrics that measures semantic content overlap via similarity in learned embedding spaces (e.g.
@@
 For dense captioning, where captions are localized to parts of the scene, adjusted benchmarks are needed.
-Often, BLEU, ROUGE, METEOR and CIDEr scores are still used, however the score is set to zero if the intersection over union (IoU) between the predicted bounding box and the object is less than a threshold “$k$”.
-Typical “$k$” values are 0.25 and 0.5 [19, 217, 226].
+Often, BLEU, ROUGE, METEOR and CIDEr scores are still used, however the score is set to zero if the intersection over union (IoU) between the predicted bounding box and the object is less than a threshold “kk”.
+Typical “kk” values are 0.25 and 0.5 [19, 100, 109].
 However, these metrics focus on the captioning recall while ignoring false positives.
@@
 Evaluation Metrics for 3D grounding require comparing predicted locations, most often in the form of bounding boxes, with ground truth locations of objects from testing samples.
-Acc@$K$IoU [218] is the widely used metric in 3D visual grounding, which measures the percentage of positive predictions which have a intersection over union (IoU) with the ground truth greater than a threshold $K$, which is usually set to 0.25 or 0.5.
+Acc@KKIoU [101] is the widely used metric in 3D visual grounding, which measures the percentage of positive predictions which have a intersection over union (IoU) with the ground truth greater than a threshold $K$, which is usually set to 0.25 or 0.5.
 It is worth noting that some datasets evaluate the performance in different scenarios.
@@
 The key difference is that 3D-QA requires the model to answer questions from the perspective of a spectator with access to all information about the scene, while 3D-SQA needs the answer from the perspective of a player in a pre-defined situation.
-For example, 3D-SQA may ask “how many chairs are in front of me?” given the situation of “standing behind and facing the dining table”.
+For example, 3D-SQA may ask “how many chairs are in front of me?” given the situation of “facing the dining table”.
 3D Dialogue requires the model to have a coherent and natural multi-turn conversation with users about the 3D scene, instead of one round QA.
@@
 They are also used in 3D-QA, particularly the ScanQA benchmark, to measure the semantic similarity alongside the accuracy.
+Figure 3: Architectures for aligning 3D with text for LLMs.
+Here we show four high-level architectures: (a) 3D-only model that aligns 3D features to the LLM’s input space, (b) 3D+text model where 3D features and text are both aligned, (c) Q-Former style model where text is used during to condition the alignment of the 3D features, and optionally given to the LLM itself (dashed arrow), and (d) text-only approach which converts 3D representations into text strings, avoiding the need to train an alignment module.
 3.4 3D Embodied Agents (3D + Text $\rightarrow$ Action)
@@
 Other metrics include Oracle Success Rate (OSR), Trajectory Length (TL) and Goal Process (GP) [121].
-In addition to the above metrics that measure whether the agent successfully reached the target and its efficiency, it is also useful to consider how well the agent path matched the path specified by the language (when language is used to specify a detailed paths).
-One such metric is Success weighted by normalized Dynamic Time Warping (SDTW) [240] which combines the SR with the difference between the agent path and a ground truth path for a given instruction.
 Please note that our discussion focuses exclusively on metrics used within 3D-LLMs methods.
 We encourage readers to refer to Gu et al.
-[241] for summary of navigation metrics.
+[122] for a summary of navigation metrics.
 For 3D Manipulation, the key metric is Success Rate [123], which for manipulation is defined as the number of successful manipulations divided by the total number of task samples.
@@
 This involves generating 3D models for objects specified in the text as well as intelligently arranging and composing multiple 3D object models given the constraints specified in the text, like object categories, counts, spatial relationships, and scene attributes.
-Figure 2: Architectures for aligning 3D with text for LLMs.
-Here we show four high-level architectures: (a) 3D-only model that aligns 3D features to the LLM’s input space, (b) 3D+text model where 3D features and text are both aligned, (c) Q-Former style model where text is used during to condition the alignment of the 3D features, and optionally given to the LLM itself (dashed arrow), and (d) text-only approach which converts 3D representations into text strings, avoiding the need to train an alignment module.
 3D Editing refers to modifying existing 3D assets like shapes or scenes based on text instructions.
@@
 the functionality and aesthetic styles [147].
-3D data also presents unique challenges that are not present in 2D, such as the relatively high cost of obtaining and labeling 3D data, sparse 3D data structures that are not uniformly dense or aligned to a grid, and the need to reconcile multiple (possibly occluded) viewpoints of the same objects [261, 256].
+3D data also presents unique challenges that are not present in 2D, such as the relatively high cost of obtaining and labeling 3D data, working with sparse 3D data structures that are not uniformly dense or aligned to a grid, and the need to reconcile multiple (possibly occluded) viewpoints of the same objects [142, 137].
 To this end, researchers have leveraged the power of language, where the semantics and relationships within a 3D world can be embedded.
-Recent efforts in integrating Large Language Models (LLMs) with 3D data have shown promise in achieving multi-level understanding and interaction, leveraging LLMs’ inherent strengths, namely zero-shot learning, in-context learning, step-by-step reasoning, and extensive world knowledge.
+Recent efforts in integrating LLMs with 3D data have shown promise in achieving multi-level understanding and interaction, leveraging LLMs’ inherent strengths, namely zero-shot learning, in-context learning, step-by-step reasoning, and extensive world knowledge.
 In Sec.
 4.1, along with Fig.
-2, we provide a brief description of how LLMs process 3D scene information, highlighting how 3D features are aligned with language so that they can be interpreted and reasoned with via LLMs, which is foundational for subsequent sections.
+3, we provide a brief description of how LLMs process 3D scene information, highlighting how 3D features are aligned with language so that they can be interpreted and reasoned with via LLMs, which is foundational for subsequent sections.
 The rest of this section is structured to align the taxonomy presented in Fig.
-3, which describes the role LLMs have played in solving 3D tasks.
+4, which describes the role LLMs have played in solving 3D tasks.
 We start by showing how the world-knowledge (sometimes referred to as ‘common-sense knowledge’) and reasoning abilities of LLMs can enhance performance on 3D tasks in Sec.
@@
 I to contrast 3D-LLMs methods across three axes: 3D components, LLMs components, and the alignment of 3D vision and language, aiming to offer a high-level insight into the various approaches within this evolving field.
-Figure 3: Taxonomy of 3D with LLM methods.
-In Sec.
-4, we analyze the role LLMs have played in solving 3D tasks from five perspectives : Enhancing 3D Tasks, Multi-Task Learning, 3D Multi-modal Interfaces, Embodied Agents, and 3D Generation.
+3D Component LLM Component 3D+LLM Method Subsection 3D Geometry Vision Model Tuning LLM Abilities Fine-tuning LLM Base # Parameters Hardware Alignment Module Date Chen et al.
+[148] 4.2 SG CLIP F WK None GPT2 1.5B 1 3080 - 09/22 ConceptFusion [18] 4.2/4.4 RGB-D OpenSeg F IF/R None GPT3 175B 1 3090 - 02/23 ViewRefer [20] 4.2 MVI Multi-View Transformer T WK None GPT3 175B 4 A100 Transformer 03/23 LLM-Grounder [21] 4.2 PC/NeRF OpenScene/LERF F IF/ICL/R None GPT3.5/4 - - - 09/23 Abdelreheem et al.
+[22] 4.2 Mesh Vision Transformer F IF/WK None GPT3.5 - 1 3090 - 09/23 Transcribe3D [23] 4.2 PC Groupfree Transformer F IF/R None GPT3.5/4 - - - 10/23 zero-shot 3DVG [16] 4.2 RGB-D Mask3d F IF/ICL/R None GPT3.5/4 - - Transformer 11/23 3DAP [149] 4.2 MVI - - IF/ICL/R PT GPT4V - - - 12/23 SpatialPIN [150] 4.2 RGB-D LanSAM/Depth Anything F IF/R PT GPT4V - - - 03/24 SIG3D [151] 4.2 PC OpenScene F IF/ICL/R LF Flan-T5 2.7B 1 A100 MLP 10/24 3D-LLM [115] 4.3 PC/VG Mask2Former/SAM F IF/ICL/R LF OPT/Flan-T5 9B/2.7B/3B 64 V100 QFormer 07/23 Chat-3D [80] 4.3 PC Point-BERT F IF/ICL/R AF Vicuna 7B - Linear layer 08/23 LEO [152] 4.3 PC/MVI OpenClip/PointNet++ F IF/ICL/R LoRA Vicuna 7B 8 A100 Transformer 11/23 LL3DA [19] 4.3 PC/3D-BB Vote2Cap-DETR F IF/ICL/R AF OPT 1.3B 8 3090 QFormer 11/23 Point-LLM [153] 4.3 PC Point-BERT F IF/ICL/R Full LLaMA 7B/13B 8 A100 Linear layer 12/23 GPT4Point [154] 4.3 PC Point-BERT F IF/ICL/R LF OPT/Flan-T5 6.7B/3B 8 A100 QFormer 12/23 Chat-3D v2 [81] 4.3 PC Uni3D F IF/ICL/R AF Vicuna 7B 4 A40 MLP 12/23 LiDAR-LLM [155] 4.3 PC/VG VoxelNet F IF/ICL/R AF LLaMA 7B 4 A100 Transformer 12/23 Chat-Scene [156] 4.3 PC/MVI Uni-3D F IF/ICL/R LoRA Vicuna-v1.5 7B 4 A100 MLP 12/23 3DMIT [157] 4.3 PC EPCL/Uni3D F IF/ICL/R LoRA Vicuna 7B 8 A100 Linear layer 01/24 ShapeLLM [158] 4.3 PC ReCon++ F IF/ICL/R Full Vicuna 7B 8 A800 Linear layer 02/24 Scene-LLM [147] 4.3 PC/VG ConceptFusion F IF/ICL/R LF LLaMA-2 7B 32 A100 Linear layer 03/24 MiniGPT-3D [159] 4.3 PC Point-BERT F IF/ICL/R LoRA Phi-2 2.7B 1 3090 QFormer 05/24 LLaNA [160] 4.3 NeRF nf2vec F IF/ICL/R AF LLaMA-2 7B 4 A100 MLP 06/24 GreenPLM [161] 4.3 PC Uni3D F IF/ICL/R LoRA Phi-2 2.7B 1 3090 MLP 09/24 LLaVA-3D [162] 4.3 MVI/RGB-D SigLIP F IF/ICL/R Full Qwen2 7B 16 A100 MLP 09/24 Robin-3D [163] 4.3 PC/MVI Uni-3D F IF/ICL/R LoRA Vicuna-v1.5 7B 8 A6000 MLP 09/24 PerLA [164] 4.3 PC/SG Vote2Cap-DETR F IF/ICL/R AF OPT 1.3B 2 H100 QFormer 11/24 Video-3D LLM [165] 4.3 MVI/RGB-D SigLIP F IF/ICL/R Full Qwen2 7B 8 A100 Transformer 12/24 GPT4Scene [166] 4.3 MVI SigLIP F IF/ICL/R Full Qwen2 7B 8 A100 Transformer 01/25 LSceneLLM [167] 4.3 PC OpenScene F IF/ICL/R Full LLaMA-2 7B 64 V100 MLP 02/25 Inst3D-LMM [168] 4.3 PC/RGB-D Uni-3D F IF/ICL/R LoRA Vicuna-v1.5 7B 8 A100 MLP 03/25 SplatTalk [169] 4.3 MVI/3DGS SigLIP F IF/ICL/R LoRA Qwen2 7B 1 H100 MLP 03/25 Ross3D [170] 4.3 MVI SigLIP F IF/ICL/R Full Qwen2 7B 8 A100 MLP 04/25 3D-LLaVA [171] 4.3 PC 3D U-Net F IF/ICL/R LoRA Vicuna-v1.5 7B 8 3090 MLP 04/25 Spatial-MLLM [172] 4.3 MVI VGGT F IF/ICL/R Full Qwen2.5 3B 4 A800 MLP 05/25 VG LLM [173] 4.3 MVI VGGT F IF/ICL/R Full Qwen2.5 3B 8 H800 Linear layer 05/25 VLM-3R [174] 4.3 MVI CUT3R F IF/ICL/R LoRA Qwen2 7B 16 H200 MLP 05/25 LEO-VL [175] 4.3 MVI/RGB-D Qwen2.5-VL ViT F IF/ICL/R LoRA Qwen2.5 7B 8 A100 MLP 06/25 3DRS [176] 4.3 MVI/RGB-D VGGT F IF/ICL/R Full Qwen2 7B 8 H100 Transformer 06/25 Point-Bind [177] 4.4 PC I2P-MAE F IF/ICL/R LF LLaMA 7B 8 A100 Linear layer 09/23 JM3D-LLM [178] 4.4 PC PointNet++/Point-BERT F IF/ICL/R LF Vicuna 7B 3 A100 MLP 10/23 MultiPLY [24] 4.3/4.4/4.5 PC ConceptGraph T IF/ICL/R AF Vicuna 13B 128 V100 Linear layer 01/24 3DLLM-Mem [179] 4.4 MVI/RGB-D SigLIP F IF/ICL/R Full Qwen2 7B 8 TPU v5p MLP 05/25 LLM-Planner [12] 4.5 VG Seg&Depth F IF/ICL/R None GPT3 175B - - 03/23 SayPlan [180] 4.5 SG SG generator F IF/ICL/R None GPT4 - - - 07/23 VoxPoser [13] 4.5 RGB-D OWL-ViT/SAM F IF/ICL/R None GPT4 - - - 07/23 UniHSI [181] 4.5 RGB-D - - IF/ICL/R None GPT3.5/4 - 1 A100 - 09/23 LAN-grasp [14] 4.5 MVI OWL-ViT F IF/ICL/R None GPT4 - - - 10/23 Agent3D-Zero [17] 4.3/4.5 MVI - - IF/ICL/R PT GPT4V - - - 03/24 NaviLLM [11] 4.5 MVI EVA-CLIP-Large F IF/ICL/R LF Vicuna 7B 8 A100 Transformer 12/23 ManipLLM [15] 4.5 RGB-D CLIP F IF/ICL/R AF LLaMA 7B 1 A100 Linear layer 03/24 3D-VLA [182] 4.3/4.5 PC/MVI Mask2Former/SAM F IF/ICL/R LoRA Flan-T5 3B 384 V100 QFormer 03/24 PolyGen [183] 4.6 n-gon Mesh Custom Transformer F - Full Custom - 4 V100 - 02/20 LLMR [184] 4.6 Unity, Mesh Dall-E-2, CLIP F IF/ICL/R None GPT4 - 1 3080 - 09/23 3D-GPT [26] 4.6 Blender, Mesh - - IF/R None GPT3.5/4 - - - 10/23 MeshGPT [185] 4.6 Mesh GPT-2-style Transformer F - Full GPT-2-style $\sim$345M 4 A100 - 11/23 ShapeGPT [186] 4.6 SDF 3D VQ-VAE, CLIP F IF/R Full T5 - 4 A100 T5 11/23 Holodeck [187] 4.6 Mesh - - IF/R None GPT4 - 8 RTX 8k CLIP/SBERT 12/23 GALA-3D [25] 4.6 3DGS MVDream, Control-Net F IF/R None Any (GPT3.5) - - - 02/24 LLaMA-Mesh [188] 4.6 Mesh - F IF/R Full LLaMA-3.1 8B 32 A100 MLP 11/24
+TABLE I: Summarization of 3D-LLMs methods.
+The 3D Geometry column collects the 3D geometric information being used for each method, such as Point Cloud (PC), Multi-view Images (MVI), RGB+Depth (RGB-D), 3D Bounding Box (3D-$B$), Scene Graph (SG), Voxel Grid (VG), Gaussian Splatting (3DGS) and NeRF.
+The tuning column outlines whether the vision model is finetuned during training (True/False).
+IF, ICL, R, and WK denote the LLM Abilities of Instruction-following, In-Context Learning, Reasoning, and World Knowledge.
+The Fine-tuning column summarizes how LLM components are fine-tuned, such as Prompt Tuning (PT), Low-Rank Adaptation (LoRA), Adaptive Fine-tuning (AF), Layer Freezing (LF), or Full Fine-tuning (Full).
+The Hardware column shows methods that are trained either with numbers of Nvidia GPU and the specific GPU type or with no training involved.
 4.1 How do LLMs process 3D scene information?
@@
 Multi-view Images [11, 14, 17, 20, 152, 149] are also frequently used as research on 2D feature extraction is well-established, which means that 3D feature extraction requires only an additional 2D to 3D lifting scheme.
-RGB-D data, readily obtained using depth cameras, is commonly used in 3D embodied agent systems to extract viewpoint-relevant information for navigation and understanding.
-3D Scene Graphs are a more abstract 3D representation, excelling in modeling the presence of objects and their relationship, and capturing high-level information of scenes.
-They are often utilized in 3D scene classification [274] and planning tasks [275].
-NeRFs are currently less used in 3D-LLM methods [21].
+RGB-D data, readily obtained using depth cameras, is commonly used in 3D embodied agent systems to extract viewpoint-relevant information for navigation and understanding [162, 165, 168, 175, 176].
+3D Scene Graphs provide more abstract 3D representation, excelling in modeling the presence of objects and their relationship, and capturing high-level information of scenes.
+They are often utilized in 3D scene classification [148], understanding [164] and planning tasks [180].
+NeRFs are currently less used in 3D-LLM methods [21, 160].
 We believe this is due to their implicit nature which makes them more difficult to tokenize and integrate with feed-forward neural networks.
 Current methods use different architectures (see Fig.
-2) and modules that align 3D features with LLM input spaces (see 3D+LLM column in Tab.
+3) and modules that align 3D features with LLM input spaces (see 3D+LLM column in Tab.
 I).
 For models which only accept 3D inputs (Fig.
-2a), a linear layer [269, 24, 266] or an MLP [171, 172] is used as the alignment module to transform 3D features into the LLM input space.
+3a), a linear layer [157, 24, 147] or an MLP [80, 81] is used as the alignment module to transform 3D features into the LLM input space.
 Models which accept 3D and text as input often use two separate branches to align 3D features and text (Fig.
-2b).
+3b).
 Some works [80, 81] employ a one-layer vanilla transformer to allow 3D object features to attend to each other during alignment.
@@
 Other works [115, 154, 182] follow the Q-Former style approach of [87] to align 3D features and text (Fig.
-2c), introducing fixed length query tokens as additional input, and following a BERT-based structure to facilitate interaction between 3D and text features during alignment.
-Mostly, these three types of architectures described above achieve alignment by utilizing 3D captioning datasets [218], where the captioning loss, i.e.the cross-entropy loss between captions generated by LLMs and the brief, ground truth description of the scene, is used to fine-tune the alignment module, while freezing the pre-trained 3D feature extractor and LLM.
-Finally, some models [21, 20, 22, 23, 273, 12, 275, 277, 14, 17] employ closed-sourced models like ChatGPT and do not train an alignment module at all (Fig.
-2d).
+3c), introducing fixed length learnable query tokens as additional input, and following a BERT-based structure to facilitate interaction between 3D and text features during alignment.
+Mostly, these three types of architectures described above achieve alignment by utilizing 3D captioning datasets [101], where the captioning loss, i.e., the cross-entropy loss between captions generated by LLMs and the brief, ground truth description of the scene, is used to fine-tune the alignment module, while freezing the pre-trained 3D feature extractor and LLM.
+Figure 4: Taxonomy of 3D with LLM methods.
+In Sec.
+4, we analyze the role LLMs have played in solving 3D tasks from five perspectives: Enhancing 3D Tasks, Multi-Task Learning, 3D Multi-modal Interfaces, Embodied Agents, and 3D Generation.
+Finally, some models [21, 20, 22, 23, 149, 12, 180, 181, 14, 17] employ closed-source models like ChatGPT and do not train an alignment module at all (Fig.
+3d).
 Instead of aligning 3D features with the LLM input space, text descriptions are directly generated from the 3D data, such as by describing 3D bounding boxes, positions and relationships, or by using pre-existing captions.
@@
 The following two sections describe these approaches respectively.
-3D Component LLM Component 3D+LLM Method Subsection 3D Geometry Vision Model Tuning LLM Abilities Fine-tuning LLM Base # Parameters Hardware Alignment Module Date Chen et al.
-[274] 3.2 SG CLIP F WK None GPT2 1.5B 1 RTX3080 - 09/22 ConceptFusion [18] 3.2/3.4 RGB-D OpenSeg F IF/R None GPT3 175B 1 3090 - 02/23 ViewRefer [20] 3.2 MVI Multi-View Transformer T WK None GPT3 175B 4 A100 Transformer 03/23 LLM-Grounder [21] 3.2 PC/NeRF OpenScene/LERF F IF/ICL/R None GPT3.5/4 - - - 09/23 Abdelreheem et al.
-[22] 3.2 Mesh Vision Transformer F IF/WK None GPT3.5 - 1 3090 - 09/23 Transcribe3D [23] 3.2 PC Groupfree Transformer F IF/R None GPT3.5/4 - - - 10/23 zero-shot 3DVG [16] 3.2 RGB-D Mask3d F IF/ICL/R None GPT3.5/4 - - Transformer 11/23 3DAP [273] 3.2 MVI - - IF/ICL PT GPT4V - - - 12/23 3D-LLM [153] 3.3 PC Mask2Former/SAM F IF/ICL/R LF OPT/Flan-T5 9B/2.7B/3B 64 V100 QFormer 07/23 Chat-3D [171] 3.3 PC Point-BERT F IF/ICL/R AF Vicuna 7B - Linear layer 08/23 LEO [270] 3.3 PC/MVI OpenClip/PointNet++ F IF/ICL/R LoRA Vicuna 7B 8 A100 Transformer 11/23 LL3DA [19] 3.3 PC/3D-BB ScanNet scene encoder F IF/ICL/R AF OPT 1.3B 8 3090 QFormer 11/23 Point-LLM [267] 3.3 PC Point-BERT F IF/ICL/R Full LLaMA 7B/13B 8 A100 Linear layer 12/23 GPT4Point [268] 3.3 PC Point-BERT F IF/ICL/R LF OPT/Flan-T5 6.7B/3B 8 A100 QFormer 12/23 Chat-3D v2 [172] 3.3 PC Uni-3D F IF/ICL/R AF Vicuna 7B 4 A40 MLP 12/23 LiDAR-LLM [271] 3.3 PC/VM VoxelNet F IF/ICL/R AF LLaMA 7B 4 A100 Transformer 12/23 3DMIT [269] 3.3 PC EPCL/Uni3D F IF/ICL/R LoRA Vicuna 7B 8 A100 Linear layer 01/24 Scene-LLM [266] 3.3 PC/VM ConceptFusion F IF/ICL/R LF LLaMA-2 7B 32 A100 Linear layer 03/24 Point-Bind [272] 3.4 PC I2P-MAE F IF/ICL/R LF LLaMA 7B 8 A100 Linear layer 09/23 JM3D-LLM [279] 3.4 PC PointNet++/PointMLP/Point-BERT F IF/ICL/R LF Vicuna 7B 3 A100 MLP 10/23 LLM-Planner [12] 3.5 VM Seg&Depth F IF/ICL/R None GPT3 175B - - 03/23 SayPlan [275] 3.5 SG SG generator F IF/ICL/R None GPT4 - - - 07/23 VoxPoser [13] 3.5 RGB-D OWL-ViT/SAM F IF/ICL/R None GPT4 - - - 07/23 UniHSI [277] 3.5 RGB-D - - IF/ICL/R None GPT3.5/4 - 1 A100 - 09/23 LAN-grasp [14] 3.5 MVI OWL-ViT F IF/ICL/R None GPT4 - - - 10/23 Agent3D-Zero [17] 3.3/3.5 MVI - - IF/ICL/R PT GPT4V - - - 03/24 NaviLLM [11] 3.5 MVI EVA-CLIP-Large F IF/ICL/R LF Vicuna 7B 8 A100 Transformer 12/23 MultiPLY [24] 3.3/3.4/3.5 PC ConceptGraph T IF/ICL/R AF Vicuna 13B 128 V100 Linear layer 01/24 ManipLLM [15] 3.5 RGB-D CLIP F IF/ICL/R AF LLaMA 7B 1 A100 Linear layer 03/24 3D-VLA [276] 3.3/3.5 PC/MVI Mask2Former/SAM F IF/ICL/R LoRA Flan-T5 3B 384 V100 QFormer 03/24 PolyGen [280] 3.6 n-gon Mesh Custom Transformer F - Full Custom - 4 V100 - 02/20 LLMR [281] 3.6 Unity, Mesh Dall-E-2, CLIP F IF/ICL/R None GPT4 - 1 3080 - 09/23 3D-GPT [26] 3.6 Blender, Mesh - - IF/R None GPT3.5/4 - - - 10/23 MeshGPT [282] 3.6 Mesh GPT-2-style Transformer F - Full GPT-2-style $\sim$345M 4 A100 - 11/23 ShapeGPT [283] 3.6 SDF 3D VQ-VAE, CLIP F IF/R Full T5 - 4 A100 T5 11/23 Holodeck [284] 3.6 Mesh - - IF/R None GPT4 - 8 RTX 8k CLIP/SBERT 12/23 GALA-3D [25] 3.6 3DGS MVDream, Control-Net F IF/R None Any (GPT3.5) - - - 02/24
-TABLE I: Summarization of 3D-LLMs methods.
-The 3D Geometry column collects the 3D geometric information being used for each method, such as Point Cloud (PC), Multi-view Images (MVI), RGB+Depth (RGB-D), 3D Bounding Box (3D-BB), Scene Graph (SG), Voxel Map (VM) and NeRF.
-The tuning column outlines whether the vision model is finetuned during training (True/False).
-IF, ICL, R, and WK denote the LLM Abilities of Instruction-following, In-Context Learning, Reasoning, and World Knowledge.
-The Fine-tuning column summarizes how LLM components are fine-tuned, such as Prompt Tuning (PT), Low-Rank Adaptation (LoRA), Adaptive Fine-tuning (AF), Layer Freezing (LF), or Full Fine-tuning (Full).
-The Hardware column shows methods that are trained either with numbers of Nvidia GPU and the specific GPU type or with no training involved.
 4.2.1 Knowledge-enhanced approaches
@@
 Here, the knowledge embedded in LLMs is used to determine the room category based on the object category information contained in the room.
-Firstly, this approach creates scene graphs from Matterport3D [285] data with nodes for regions and objects, and object nodes linked to room nodes.
+Firstly, this approach creates scene graphs from Matterport3D [190] data with nodes for regions and objects, with object nodes linked to room nodes.
 Next, key objects are selected to form queries per room type.
 LLMs score descriptions extracted from the selected objects, with the highest score predicting the room label.
-Spatial information like sizes or locations can also be provided.
 ViewRefer [20] uses LLMs to expand the grounding texts with view-related descriptions.
@@
 By rephrasing input texts and their opposite-view synonyms several times, the model improves cross-view grounding.
-It also adopts a fusion transformer with inter-view attention and includes learnable multi-view prototypes, which capture inter-view knowledge and further boost 3D grounding performance.
 Abdelreheem et al.
-[22] address the problem of semantic correspondence in 3D shapes.
-They classify 3D shapes by feeding rendered views into a BLIP2 model to generate lists of class proposals.
-ChatGPT [286] unifies these into a single class per shape.
-ChatGPT also generates semantic part names and pairwise mappings (e.g.
-arm $\rightarrow$ wing).
-A 3D segmenter then segments shapes based on semantic regions, utilizing the part mappings to produce sparse correspondence maps.
-The knowledge-enhanced strategies described above enable strong performance especially in zero-shot scenarios where no labeled 3D data is available for a particular object or scene type.
+[22] address the problem of semantic correspondence in 3D shapes by using a BLIP2 model to generate lists of text class proposals from different rendered views of a 3D shape.
+ChatGPT [191] then unifies these proposals into a single class proposal for the entire object.
+The knowledge-enhanced strategies described above enable strong performance, especially in zero-shot scenarios where no labeled 3D data is available for a particular object or scene type.
 This allows open-ended reasoning about object parts, relationships, and semantics beyond fixed ontologies, as demonstrated by (i) Chen et al.
@@
 [22] generating and matching object part semantics across shapes.
-Method
+4.2.2 Reasoning-enhanced approaches
+Along with world knowledge, the reasoning abilities of LLMs also help tackle other 3D tasks, particularly visual grounding in complicated 3D scenes with detailed geometry and multiple objects.
+In such cases, text descriptions of the objects should include their appearance and spatial relationship with surrounding items.
+Ordinary grounding methods [192] often struggle in this setting due to their inability to understand detailed text descriptions.
+LLM-Grounder [21], Transcribe3D [23], and zero-shot 3DVG [16] approach this problem by leveraging the reasoning capability of LLMs to analyze the text descriptions and generate a sequence of instructions for locating the objects using existing grounding toolboxes.
+Specifically, an LLM first identifies anchor and target objects from the text descriptions.
+It then analyzes the spatial relationship (or described attributes) between multiple candidate objects based on their coordinates returned by the grounding tool to select the candidate that best matches the text description.
+A common drawback of these approaches is the “blindness” of the LLM, since it is provided only with abstract textual descriptions of the 3D scene instead of the original point cloud.
+This may lead to the loss of crucial scene details.
+Consequently, when a 3D scene contains multiple objects of the same class, the absence of necessary scene details means that ambiguities of the text-based references cannot be resolved, which limits overall performance.
+Besides visual grounding, the reasoning capability of LLMs can facilitate other tasks.
+3DAP [149] leverages GPT-4V to infer 3D information of an object from its 2D images using a visual prompting technique, where it annotates the input image with a 3D axis to enhance the LLM’s awareness of the 3D scale.
+ConceptFusion [18] uses GPT3 to generate instructions using pre-defined elementary spatial comparison modules to enable more complex spatial reasoning using their proposed 3D feature map.
+SpatialPIN [150] utilizes GPT-4V to infer interactive objects, potential tasks and plan of actions from a single RGB image.
+4.3 LLMs for 3D Multi-Task Learning
+Method 3D Captioning 3D Grounding 3D Conversation 3D Embodied Agents 3D Generation
 Object
-Captioning
 Scene
-Captioning
 Dense
-Captioning
-Grounding QA
-Situated
-QA
-Dialogue Planning Navigation Manipulation
+Single
+Multi
+QA Situated-QA Dialogue Plan. Nav. Manip. object
 3D-LLM [115]
@@
 Agent3D-Zero [17]
+ShapeLLM [158]
 Scene-LLM [147]
+Spartun3D-LLM [193]
+Chat-Scene [178]
+LLaVA-3D [162]
+Robin-3D [163]
+Video-3D LLM [165]
+GPT4Scene [166]
+LSceneLLM [167]
+Ross3D [170]
+3D-LLaVA [171]
+VG-LLM [173]
+3DRS [176]
 TABLE II: Comparison of tasks included in various 3D multi-task learning methods. QA stands for Question Answering. The methods are sorted chronologically from top (oldest) to bottom (most recent).
-4.2.2 Reasoning-enhanced approaches
-Along with world knowledge, the reasoning abilities of LLMs also help tackle other 3D tasks, particularly visual grounding in complicated 3D scenes with detailed geometry and multiple objects.
-In such cases, text descriptions of the objects should include their appearance and spatial relationship with surrounding items.
-Ordinary grounding methods [287] often struggle in this setting due to their inability to understand detailed text descriptions.
-LLM-Grounder [21], Transcribe3D [23], and zero-shot 3DVG [16] approach this problem by leveraging the reasoning capability of LLMs to analyze the text descriptions and generate a sequence of instructions for locating the objects using existing grounding toolboxes.
-Specifically, an LLM first identifies anchor and target objects from the text descriptions.
-It then analyzes the spatial relationship (or described attributes) between multiple candidate objects based on their coordinates returned by the grounding tool to select the candidate that best matches the text description.
-In addition, (i) Transcribe3D [23] and LLM-Grounder [21] employ a multi-round interactive question-and-answer process to help users clarify their intentions, prompting them to provide instructions with more information to arrive at more accurate results, whereas (ii) LLM-Grounder includes multiple grounding tool choices, such as OpenScene [29] or LERF [30] to accommodate different 3D representations like point clouds or NeRF.
-A common drawback of these approaches is the “blindness” of the LLM since it is provided only with abstract textual descriptions of the 3D scene instead of the original point cloud of the scene.
-This may lead to the loss of crucial scene details.
-Consequently, when a 3D scene contains multiple objects of the same class, the absence of necessary scene details means that ambiguities of the text-based references cannot be resolved, which limits overall performance.
-Besides visual grounding, the reasoning capability of LLMs facilitates other tasks as well.
-3DAP [273] leverages GPT-4V to infer 3D information of an object from its 2D images using a visual prompting technique, where it annotates the input image with a 3D axis to enhance the LLM’s awareness of the 3D scale.
-ConceptFusion [18] uses GPT3 to generate instructions using pre-defined elementary spatial comparison modules to enable more complex spatial reasoning using their proposed 3D feature map.
-4.3 LLMs for 3D Multi-Task Learning
 Many works focus on using the instruction-following and in-context learning capabilities of LLMs to unify multiple 3D tasks into a single language space.
-By employing different text prompts to denote different tasks, these studies aim for LLMs to serve as a unifying dialogue interface.
+By using different text prompts to denote different tasks, these works allow for LLMs to serve as a unifying dialogue interface.
 Achieving multi-task learning using an LLM typically involves several key steps, beginning with the construction of 3D-text data pairs [19, 115, 152].
@@
 Alignment modules [80, 115, 81, 147] are subsequently used to (i) align 3D features with text embeddings from the LLMs across multiple levels (object-level, relationship-level, and scene-level) and (ii) translate the 3D features into tokens interpretable by the LLMs.
-Finally, an appropriate training strategy [272, 19, 153, 270, 269] needs to be selected, such as single or multi-stage 3D-language alignment training and multi-task instruction fine-tuning.
-In the remainder of this section, we explore these aspects in detail. We additionally summarize the scope and capabilities of each method reviewed in this section in Table II.
+Finally, an appropriate training strategy [19, 115, 177, 152, 157, 196] needs to be selected, such as single or multi-stage 3D-language alignment training and multi-task instruction fine-tuning.
+In the remainder of this section, we explore these aspects in detail.
+We additionally summarize the scope and capabilities of each method reviewed in this section in Table II.
 4.3.1 Data for Multi-Task Learning
-As demonstrated in Table II, we divide tasks into four categories: captioning, grounding, question answering (QA), and embodied agent tasks (i.e., planning, navigation, and manipulation).
+As shown in Fig.
+2, we divide tasks into five categories: captioning, grounding, question answering (QA), embodied agent tasks (i.e., planning, navigation, and manipulation), and generation.
 Accordingly, each task’s textual output follows a predefined format.
@@
 Existing methods follow these guidelines to construct their multi-task instruction fine-tuning datasets.
-Once the text format has been decided upon, different methods use different strategies to obtain text annotations for their datasets.
-Several approaches utilize human labelers to generate ‘ground truth’ annotations for each sample [218, 228, 234, 233], however this can be an expensive and time-consuming process.
-Another approach is to use ChatGPT [286] to generate text annotations for each sample, which is a strategy utilized by 3DMIT [269], LiDAR-LLM [271], Chat-3D [171] and Chat-3D v2 [172].
+Once the text format has been decided upon, different works use different strategies to obtain text annotations for their datasets, constructing 3D-text pairs for training.
+Several approaches ask human labelers to generate ‘ground truth’ annotations for each sample [101, 111, 58, 59].
+However this can be an expensive and time-consuming process.
+Another approach [157, 155, 80, 158, 81] is to use an LLM (e.g.
+ChatGPT [191]) to generate text annotations for each sample.
 Here, 3D scene data is converted into text (usually by describing object bounding boxes and spatial relations textually), and a task description is created to describe the desired output.
 To guide ChatGPT towards the expected output format for the task, demonstration examples are provided, which allows ChatGPT to perform in-context learning to generate plausible text annotations for other 3D scenes.
-Alternatively, other multi-task datasets [19, 266] are constructed simply be merging existing 3D Vision-Language (VL) datasets [218, 228, 234, 153].
-Some multi-task datasets are constructed using a combination of these three approaches, as seen in LEO [270], LiDAR-LLM [271] and 3D-LLM [153], seeking to combine the accuracy of human annotations with the scalability provided by using LLM-generated annotations.
-4.3.2 Training an LLM for multiple 3D tasks
-The first step in training an LLM for multiple 3D tasks involves obtaining meaningful 3D features, where the extraction method varies depending on the type of 3D scene.
-For individual object point clouds, Point-LLM [267], Chat-3D [171], and GPT4Point [268] utilize Point-BERT [288] to extract 3D object features.
-For indoor scenes, LEO [270] uses PointNet++ [34] for feature extraction, whereas Chat-3D v2 [172] and 3DMIT [269] segment the scene and use Uni-3D [289] to extract features for each segmented part.
-Meanwhile, MultiPLY [24] integrates the extracted object features into a scene graph [290] to represent the entire scene.
-3D-LLM [153] and Scene-LLM [266] lift features from 2D multi-view images into a 3D representation.
-3D-LLM [153] lifts 2D semantic features from Mask2Former [291] or SAM [211].
-Scene-LLM [266] follows ConceptFusion [18] to fuse both global information and local details, mapping pixel-wise CLIP features into point-wise 3D features.
-For outdoor 3D scenes, LiDAR-LLM [271] uses VoxelNet [292] to extract 3D voxel features.
-For alignment modules, as discussed in Sec.
-4.1, various network architectures are used.
-Notably, MultiPLY [24] employs distinct linear layers to align features from each modality.
-Chat-3D [171] and Chat-3D v2 [172] use a one-layer vanilla transformer to allow 3D object features to attend to each other during alignment.
-LEO [270] and LiDAR-LLM [271] utilize modified transformers as their alignment modules to better suit different types of 3D data (dense point clouds vs.
-sparse LiDAR).
-LEO [270] modifies the self-attention mechanism to explicitly encode spatial relations between object pairs in point clouds.
-In contrast, LiDAR-LLM [271] employs both self-attention and cross-attention mechanisms to align Bird’s Eye View (BEV) features with text features.
-3D-LLM [153] and GPT4Point [268] adopt Q-Former, while LL3DA[19] adds an additional branch on top of Q-Former that allows query tokens to interact with visual prompts provided by users.
+Alternatively, other multi-task datasets [19, 147] are constructed by merging existing 3D Vision-Language (VL) datasets [101, 111, 58, 115].
+Several recent multi-task datasets adopt a hybrid strategy that combines these three approaches [152, 155, 115], aiming to leverage both the accuracy of human annotations and the scalability of LLM-generated ones.
+4.3.2 3D Feature Tokenization for LLM training
+The first step in training an LLM for multi-task 3D understanding lies in extracting semantically and spatially meaningful 3D features.
+We group existing methods into two main categories depending on their input modality and spatial encoding strategies.
+Point-based methods operate directly on point clouds, offering explicit spatial grounding and high geometric precision.
+They can be further divided into: (1) Scene-centric models [115, 162, 19, 154, 157, 158, 153], which process raw scene point clouds using pretrained or custom point encoders such as Point-BERT [194] to obtain the 3D scene features, then alignment module (such as MLP) is used to project 3D features to 3D tokens; (2) Object-centric models [152, 80, 81, 156, 168, 163, 158, 153, 159, 161], instead of processing the whole 3D scenes, each 3D object is segmented via 3D segmentation, such as Mask3D [197] and then encoded to 3D object feature.
+Additional 3D positional embeddings are incorporated into each object feature to generate the final 3D tokens.
+While effective at modeling discrete entities, they often lack comprehensive scene-level spatial understanding; and (3) Voxel-centric methods [147, 155], which discretize the 3D space into voxel grids and apply an encoder such as VoxelNet [198] to get 3d voxel feature.
+Despite the advantage of maintaining spatial structures and metric accuracy, these methods generally suffer from limited semantic richness due to annotation scarcity and the sparsity of 3D data.
+Moreover, due to the relatively small scale of existing point cloud datasets, 3D-LLMs trained solely on point-based representations typically underperform their 2D counterparts in terms of general-purpose semantic understanding [199].
+Multi-view image-based methods directly encode multi-view images of a 3D scene using a 2D image encoder, and then lift the 2D image features into 3D space to obtain 3D tokens.
+The advantage of this paradigm is that it combines the semantic richness and well-aligned text-image representation of large-scale 2D vision-language models (VLMs) with 3D spatial reasoning.
+These can be categorized as: (1) Projection-based methods such as LEO-VL [175], which project 2D pixel features from 2D images into 3D coordinates using calibrated camera intrinsics, thus obtaining a feature for each 3D point.
+By further embedding the XYZ coordinates through 3D positional encoding, the final 3D tokens are generated; (2) Reconstruction-based methods [169, 160], which first reconstruct a 3D scene (e.g., via 3D gaussian splatting or NeRF) and then align it with language; and (3) Geometry-augmented methods [173, 172, 174], which use VGGT [200]-like encoder to generate both geometry-aware and semantic representations without explicit projection or reconstruction.
+These approaches benefit from pretrained VLMs and rich 2D image-text datasets, yet the additional lifting step can introduce projection errors and lead to imprecise spatial relationships in the constructed 3D space.
+(4) Temporal-augmented methods [162, 165, 166, 170, 176] repurpose video streams as implicit 3D observations, leveraging temporal continuity as a proxy for spatial structure.
+these methods feed video frames to existing 2D-VLMs, without architectural modifications, directly inherit their strong language grounding and reasoning capabilities.
+Additionally, videos are widely available, and adjacent frames are naturally correlated in both time and space, implicitly encoding 3D geometry through motion.
+This category currently yields the best empirical performance among 3D-LLMs due to the scalability and semantic richness of video-text data, as shown in Table IV.
+4.3.3 Training Strategy for Multi-Task Learning
 LLMs can be fine-tuned to incorporate multiple 3D tasks using different strategies discussed in Sec.
 2.2.3.
-LEO [270], and 3DMIT [269] use low-rank adaptation (LoRA) for fine-tuning.
+LEO [152], MiniGPT-3D [159], and 3DMIT [157] use low-rank adaptation (LoRA) for fine-tuning.
 As a result, the total trainable parameters, including those of alignment modules and 3D encoders, amount to less than 10% of the original LLMs’ parameters, significantly enhancing training efficiency.
@@
 […]
-In contrast to these single-stage alignment methods, LiDAR-LLM [271], Chat-3D [171], and Chat-3D v2 [172] each employ a 2-stage 3D-language alignment process.
+By contrast, other works [80, 81, 155, 159] employ a multi-stage alignment process, with each stage designed to enhance different aspects of the LLM’s 3D capabilities.
 LiDAR-LLM [155] focuses on enhancing both local and global scene perception through 3D captioning tasks in two phases: first, concentrating on single-view captions and then expanding to panoramic scene descriptions.
 They develop instance-level perception capabilities through a blend of captioning and grounding tasks.
-Chat-3D [171] first aligns 3D objects with text using datasets for 3D object classification [296, 293, 297] aiming to maximize the cosine similarity between mapped 3D object features and object category word embeddings by updating only the alignment module.
-In the second stage of scene-level alignment, it utilizes ScanRefer [218] to enable captioning capabilities, updating an additional transformer layer specifically to model spatial relationships of objects.
-Similarly, Chat-3D v2 [172] incorporates object-level and scene-level alignments, with the second stage additionally training a position embedding layer.
+Chat-3D [80] first aligns 3D objects with text using datasets for 3D object classification [204, 201, 205], aiming to maximize the cosine similarity between mapped 3D object features and object category word embeddings.
+In the second stage, it utilizes ScanRefer [101] to enable captioning capabilities, updating an additional transformer layer specifically to model spatial relationships of objects.
+Similarly, Chat-3D v2 [81] incorporates object-level and scene-level alignments, with the second stage training an additional position embedding layer.
+MiniGPT-3D [159] additionally fine-tunes a ‘Mixture of Query’ module to enhance their model’s 3D perception ability from multiple perspectives.
+GreenPLM [161] propose to only align text caption of 3D data to LLM at the first two stages and then achieve object-level alignment by tuning the MLP layer.
 For training efficiency, both LL3DA [19] and 3DMIT [157] skip the alignment stages and focus solely on the stage of instruction tuning described below.
-Nearly all multitask learning methods ultimately require the ability to complete various 3D tasks based on instructions.
+Nearly all multi-task learning methods ultimately require the ability to complete various 3D tasks based on instructions.
 Thus, as the final stage of training, each method typically uses their own constructed multitask instruction-following dataset for instruction fine-tuning [^1] Instruction fine-tuning refers to the process of further fine-tuning LLMs on a dataset consisting of (INSTRUCTION, OUTPUT) pairs [206].
-Since all task outputs are unified into textual form, the training loss used is the standard autoregressive loss used in LLMs.
+Since all task outputs are unified into textual form, the training loss used is the standard auto-regressive loss used in LLMs, except that Ross3D [170] uses an additional reconstructive loss.
 This stage usually involves jointly training the alignment module and the LLM.
-An exception is Agent3D-Zero [17], which completes various 3D tasks by feeding GPT4V with 2D images from different viewpoints, thus it does not require any training.
+An exception is Agent3D-Zero [17], which completes various 3D tasks by feeding GPT-4V with 2D images from different viewpoints, thus it does not require any training.
 4.4 LLMs as 3D Multi-Modal Interfaces
@@
 Compared with methods on object-level scenes, approaches that can process large scenes reduce costs by relying on pre-trained encoders to bridge the modality gap, instead of learning new ones from scratch.
-4.5 LLMs for Embodied Agents
+4.5 LLMs for 3D Embodied Agents
 3D embodied agents can be created using the planning, tool-using, and decision-making abilities of LLMs.
@@
 LEO [152] and LLM-Planner [12] utilize LLMs to generate step-by-step plans which they adjust dynamically based on environmental perceptions.
-LEO [270] emphasizes scene-aware planning grounded in the current scene’s configurations, whereas LLM-Planner [12] adopts GPT3 [127] to divide planning into high-level subgoals and low-level actions, and replan when the agent gets stuck during task execution.
+LEO [152] emphasizes scene-aware planning grounded in the current scene’s configurations, whereas LLM-Planner [12] adopts GPT-3 [70] to divide planning into high-level subgoals and low-level actions, and replan when the agent gets stuck during task execution.
 3D-VLA [182] integrates 3D perception, reasoning, and action through a generative world model.
@@
 While the aforementioned approaches focus on planning within a single scene, SayPlan [180] can handle multiple rooms and floors by (i) leveraging 3D scene graphs for semantic search and (ii) integrating classical path planning with an iterative replanning pipeline for plan refinement.
+NLMap [214] proposes an open-vocabulary, queryable scene map for LLM planners, enabling them to see and query available objects before generating plans.
 4.5.2 3D Navigation
@@
 It incorporates historical data from previous viewpoints to refine its navigation path towards specific goals, such as finding a printer in an office setting.
-LLM-Planner [12] employs a hierarchical approach, that first generates high-level plans as sequences of subgoals, which are then translated into a series of primitive actions by a low-level planner.
+LLM-Planner [12] employs a hierarchical approach that first generates high-level plans as sequences of subgoals, which are then translated into a series of primitive actions by a low-level planner.
 This makes the overall process adaptable to the immediate environment.
@@
 GPT4Point [154] uses a two-stream approach - aligning point cloud geometry with text via Point-QFormer, which then feeds into coupled LLM and diffusion paths for text comprehension and high-fidelity 3D object generation conforming to text input.
-In contrast, MeshGPT [282] and PolyGen [280] do not condition the generation on text, but they still adopt an autoregressive approach akin to sequence modelling in LLMs.
+In contrast, MeshGPT [185], LLaMA-Mesh [188], and PolyGen [183] do not condition the generation on text, but they still adopt an autoregressive approach akin to sequence modelling in LLMs.
 MeshGPT uses graph convolutions to encode mesh geometry/topology into rich embeddings compressed via Residual Vector Quantization and fed into a GPT-style transformer to autoregressively predict tokens/embeddings for generating meshes with desired properties.
@@
 It consists of a vertex model that unconditionally models mesh vertices, and a face model for mesh faces conditioned on input vertices using autoregressive networks to output face indices and vertex coordinates for generating diverse high-quality meshes.
+LLaMA-Mesh [188] represents 3D meshes in a purely textual form by expressing vertex coordinates and face indices as plain text sequences, with coordinate values discretized into integers for compactness, allowing the direct use of pretrained LLM tokenizers and architectures without introducing a mesh-specific vocabulary.
 4.6.2 Scene-scale Generation
@@
 While Section 4 discussed methods integrating LLMs in 3D tasks, a vast body of research has explored various aspects of 3D understanding through the lens of 2D Vision-Language Models (VLMs).
-VLMs include much richer visual information that can be linked directly to 3D.
-This section reviews contributions from a series of recent papers covering language-driven open-world understanding, instance-level understanding, unified end-to-end architectures, spatial reasoning, generation, and beyond.
+VLMs include richer visual priors that can be linked to 3D representations.
+This section reviews contributions from recent papers spanning open-world understanding, instance-level recognition, unified end-to-end architectures, spatial reasoning, and generation.
 5.1 Open-Vocabulary 3D Scene Understanding
-Open-vocabulary 3D scene understanding aims to recognize and describe scene elements using natural language descriptions instead of predefined category labels.
-OpenScene [29] takes a zero-shot approach by predicting dense features for 3D scene points that are co-embedded with CLIP’s text and image pixel embeddings in a shared feature space, enabling task-agnostic training and open-vocabulary querying to identify objects, materials, affordances, activities, and room types.
-CLIP-FO3D [304] follows a similar approach, modifying CLIP to extract dense pixel features from 3D scenes which are projected to point clouds, then training a 3D model via distillation to transfer CLIP’s knowledge.
-Semantic Abstraction [305] extracts relevancy maps from CLIP as abstracted object representations to generalize to novel semantics, vocabularies, and domains.
-Open-Fusion [306] combines the SEEM [307] vision-language model with TSDF 3D mapping for real-time open-vocabulary scene creation and querying, utilizing region-based embeddings and confidence maps.
-Approaches such as PLA [308] and RegionPLC [309] leverage contrastive learning to combine captions with 2D and 3D data modalities to associate visual and semantic information.
-PLA [308] uses 3D-caption pairs and contrastive learning to associate multi-view images with captions for learning visual-semantic representations, while RegionPLC [309] proposes region-aware contrastive learning by combining region-level captions from 2D models mapped to 3D points.
-OVIR-3D [310] fuses 2D region proposals and text-aligned features from an off-the-shelf 2D detector into 3D instances for efficient open-vocab retrieval.
-CoDA [311] uses 3D geometry priors from annotated base categories and CLIP’s 2D semantic priors in its 3D Novel Object Discovery (3D-NOD) strategy.
-Its Discovery-driven Cross-Modal Alignment (DCMA) aligns 3D and image/text features for novel object localization and classification.
-Instance-level scene understanding works such as OpenMask3D [312] and Open3DIS [313] utilize predicted class-agnostic 3D instance masks and 2D segment-level CLIP embeddings to enable open-vocabulary 3D instance segmentation.
-OpenIns3D [314] achieves open-vocabulary understanding without aligned images, using a “Mask-Snap-Lookup” pipeline that predicts 3D mask proposals, generates synthetic scene images, and assigns categories to masks via a language module.
-Rozenberszki et al.[315] propose to leverage CLIP features to ground 3D feature learning for 3D semantic and instance segmentation.
-Language grounding with NeRFs has shown promising results in open-vocabulary scene understanding.
-Several methods, such as DFF [316], LERF [30], VL-Fields [317], and 3D-OVS [318], distill the knowledge of 2D feature extractors like DINO or CLIP into 3D feature fields by minimizing the error of volume-rendered features with respect to the 2D features, enabling query-based local editing and grounding language into neural implicit representations.
-LERF [30] optimizes a dense, scale-conditioned 3D language field by volume rendering CLIP embeddings.
-LangSplat [319] and N2F2 [320] demonstrate efficient open-vocabulary querying and interaction within a 3D Gaussian Splatting representation by leveraging hierarchical supervision and multiscale feature fields.
+Open-vocabulary 3D scene understanding recognizes and describes scene elements using natural language instead of predefined category labels.
+OpenScene [217] adopts a zero-shot strategy by predicting dense 3D features co-embedded with CLIP’s text and image embeddings in a shared space, enabling task-agnostic querying of objects, materials, affordances, and room types.
+CLIP-FO3D [218] similarly distills dense pixel features from CLIP to 3D point clouds via feature projection and training a 3D model to absorb CLIP knowledge.
+Semantic Abstraction [219] extracts CLIP relevance maps as abstract object representations to generalize across vocabularies and domains.
+Open-Fusion [220] integrates SEEM [221] with TSDF 3D mapping for real-time open-vocabulary scene creation using region-based embeddings and confidence maps.
+PLA [222] and RegionPLC [223] use contrastive learning to associate captions with 2D and 3D modalities.
+PLA aligns multi-view images with captions via 3D-caption contrastive pairs, while RegionPLC aligns region-level 2D captions to 3D points.
+OVIR-3D [224] fuses 2D region proposals with CLIP-aligned features to support efficient open-vocabulary 3D instance retrieval.
+CoDA [225] incorporates annotated 3D geometry priors with 2D semantics from CLIP to discover novel objects using Discovery-driven Cross-Modal Alignment (DCMA).
+For instance-level understanding, OpenMask3D [197] and Open3DIS [226] use class-agnostic 3D instance masks and 2D segment-level CLIP embeddings for open-vocabulary 3D instance segmentation.
+OpenIns3D [227] bypasses image alignment by predicting 3D mask proposals, rendering synthetic images, and assigning mask labels via a language model.
+Rozenberszki et al.
+[228] align 3D features with CLIP for semantic and instance-level segmentation.
+Diff2Scene [229] distills semantic knowledge from frozen text-to-image diffusion models, using geometry- and saliency-aware masks to enable label-free open-vocabulary 3D segmentation and grounding.
+Language grounding with NeRFs has shown promise in open-vocabulary 3D understanding.
+Methods like DFF [230], LERF [231], VL-Fields [232], and 3D-OVS [233] distill 2D VLM knowledge into implicit 3D feature fields by minimizing rendering error w.r.t.
+2D features.
+LERF further learns a scale-conditioned language field over 3D space.
+LangSplat [234] and N2F2 [235] extend this to 3D Gaussian Splatting, enabling efficient querying via hierarchical supervision and multiscale feature fields.
 5.2 Text-Driven 3D Generation
-Section 4.6 covered methods that used LLMs for 3D generation.
-Here, we survey text-to-3D generation methods that leverage guidance from 2D VLMs [174] and text-to-image diffusion models [321, 322] using differentiable rendering.
-Early works such as DreamFields [323], CLIP-Mesh [32], CLIP-Forge [324] and Text2Mesh [325] explore zero-shot 3D generation guided by CLIP.
-DreamFusion [31] introduced Score Distillation Sampling (SDS) where the parameters of a 3D representation are optimized by making its renderings from arbitrary perspectives appear highly realistic, as evaluated by a pre-trained 2D diffusion model.
-It uses the text-to-image Imagen model [322] to optimize a NeRF representation via SDS.
-Magic3D [326] proposes a two-stage framework: generating a coarse model with a low-resolution diffusion prior and sparse 3D hash grid, then optimizing a textured 3D mesh model with an efficient differentiable renderer and high-resolution latent diffusion model [321].
-Fantasia3D [327] disentangles geometry and appearance, using a hybrid DMTet [328] representation and spatially varying BRDFs.
-ProlificDreamer [329] introduced Variational Score Distillation (VSD), a particle-based framework treating 3D parameters as random variables to improve fidelity and diversity.
-Dream3D [330] utilizes explicit 3D shape priors and text-to-image diffusion models for enhanced text-guided 3D synthesis.
-MVDream [331] employs a multi-view consistent diffusion model trainable on few-shot data for personalized generation.
-Text2NeRF [332] combines NeRF representations with pre-trained text-to-image diffusion models to generate diverse indoor/outdoor 3D scenes from language.
-Apart from simultaneously generating geometries and appearance, several studies explore the possibility of solely synthesizing the textures based on given geometries [333, 334, 335].
-For human avatars, AvatarCraft [336] uses diffusion models to guide neural implicit field geometry/texture learning from text prompts.
-Additionally, it enables animating these human avatars by deforming the neural implicit field with an explicit warping field that maps the target human mesh to a template human mesh.
-AvatarCLIP [337] proposes a zero-shot CLIP-supervised framework for 3D avatar generation, geometry sculpting, texture mapping and motion synthesis from text.
-CG-HOI [338] uses diffusion models to characterize dynamic human-object interactions from text.
-GenZI [339] distills information about human interactions via a pre-trained vision-language model, producing zero-shot 3D human-scene interaction synthesis from text prompts.
-Exploring compositional generation, CG3D [340] generates scalable 3D scenes using explicit 3D Gaussian radiance fields by compositing individual objects without bounding boxes.
+Section 4.6 reviewed LLM-guided generation; here, we survey methods leveraging 2D VLMs [82] and text-to-image diffusion models [236, 237] with differentiable rendering.
+Early methods like DreamFields [238], CLIP-Mesh [239], CLIP-Forge [240], and Text2Mesh [241] explore CLIP-guided zero-shot 3D optimization.
+DreamFusion [92] introduced Score Distillation Sampling (SDS), optimizing NeRF parameters so rendered views are realistic under a text-to-image diffusion model (Imagen [237]).
+Magic3D [242] adds a two-stage pipeline - coarse NeRF generation followed by mesh optimization using high-res latent diffusion.
+Fantasia3D [243] disentangles geometry and appearance via a DMTet [244] hybrid with spatially varying BRDFs.
+ProlificDreamer [245] introduces Variational SDS (VSD), treating 3D parameters as particles for more diverse generation.
+Dream3D [246] enhances synthesis using explicit 3D shape priors.
+MVDream [247] trains multi-view consistent diffusion on few-shot data.
+Text2NeRF [248] uses pre-trained diffusion models to guide NeRF generation from language.
+CrossOver [249] learns a modality-agnostic space across images, point clouds, CAD models, floorplans, and text, enabling cross-modal scene retrieval and generation.
+Some works target text-guided texture synthesis only [250, 251, 252].
+For avatars, AvatarCraft [253] uses diffusion to learn geometry/texture from text and animate avatars via warping fields.
+AvatarCLIP [254] provides a zero-shot CLIP-supervised framework for geometry, texture, and motion synthesis.
+GenZI [255] and CG-HOI [256] model human-object interactions from text.
+Compositional approaches like CG3D [257] assemble Gaussian radiance fields from object parts.
 Po et al.
-[341] introduce locally conditioned diffusion for granular scene control via text prompts and bounding boxes.
-GraphDreamer [342] generates compositional scenes from scene graphs by decomposing them into global-local descriptions to optimize object SDFs.
-Overall, these approaches combine diffusion models, vision-language models, neural representations, and 3D priors for text-to-3D generation of objects, avatars, and scenes.
+[258] enable local scene control via text and bounding boxes.
+GraphDreamer [259] composes scenes from graph-derived global-local descriptions using object SDF optimization.
 5.3 End-to-End Architectures for 3D Vision & Language
-Transformer models pre-trained on large 3D-text datasets learn powerful joint representations that bridge the visual and language modalities.
-3D-VisTA [343] is a transformer model that employs self-attention to jointly model 3D visual and text data, enabling effective pre-training on objectives like masked language/object modeling and scene-text matching.
-UniT3D [227] takes a unified transformer approach, combining a PointGroup 3D detection backbone, BERT text encoder, and multi-modal fusion module, with joint pre-training on synthetically generated 3D-language data.
-SpatialVLM [344] takes a different tack, and co-trains VLMs on a large synthetic 3D spatial reasoning dataset, boosting performance on 3D spatial visual question answering tasks and enabling applications like chain-of-thought reasoning for robotics.
-Multi-CLIP [345] pre-trains a 3D scene encoder to align scene features with CLIP’s text and image embeddings, aiming to transfer CLIP’s knowledge for improved 3D understanding on tasks like visual question answering.
-In addition to pre-training approaches, researchers have explored architectures that unify 3D perception with language capabilities in an end-to-end framework.
-D3Net [27] combines dense captioning and visual grounding with a 3D object detector, a speaker for generating captions from detections, and a listener for discriminating objects using the captions.
-Uni3DL [28] operates on point clouds, with modules for text encoding, point encoding, semantic/mask prediction, and diverse task outputs like segmentation, detection, grounding, and captioning.
-InstanceRefer [346] uses panoptic segmentation and linguistic cues to filter instance candidates based on language descriptions for visual grounding tasks in 3D point clouds, while LanguageRefer [347] combines language embeddings with spatial embeddings from 3D bounding boxes.
-3DVG-Transformer [348] also tackles 3D grounding in point clouds, with a coordinate-guided contextual aggregation module and multiplex attention for effective feature fusion.
-Dataset     Object Captioning     Scene Captioning     Dense Captioning     Single Object Grounding     Multiple Object Grounding     Question Answering     Situated Question Answering     Dialogue     Task Planning     Real     Synthetic     Object     Indoor     Outdoor     Human-Annotated     Model-Annotated     Template-Based
-Cap3D [349]
-Text2Shape [350]
-SceneVerse [351]
-nu-Caption [271]
-nu-Grounding [271]
-ScanRefer [218]
-ReferIt3D [228]
-Multi3DRefer [230]
-Chat-3D v2 [172]
-EmbodiedScan [352]
-ScanEnts3D [353]
-WildRefer [354]
-RIORefer [355]
-ARKitSceneRefer [356]
-ScanERU [229]
-DenseGrounding [231]
-ScanQA (Azuma et al.[234]
-ScanQA (Ye et al.[357]
-3DMV-VQA [236]
-NuScenes-QA [235]
-CLEVR3D [358]
-SQA-3D [233]
-3D-LLM [153]
-ScanScribe [294]
-M3DBench [359]
-GPT4Point [268]
-LAMM [360]
+Transformer-based 3D-text models learn joint visual-language representations.
+3D-VisTA [260] models both modalities with self-attention and is pre-trained on masked modeling and scene-text matching.
+UniT3D [110] combines PointGroup, BERT, and multi-modal fusion, trained on synthetic 3D-text data.
+SpatialVLM [261] trains VLMs on synthetic 3D spatial reasoning tasks for improved 3D spatial VQA and chain-of-thought reasoning.
+Multi-CLIP [262] aligns 3D scene encodings with CLIP’s embedding space for better cross-modal transfer.
+Lexicon3D [263] benchmarks a wide range of foundation models across 3D vision-language tasks, identifying key strengths (e.g., DINOv2 for vision, diffusion for grounding) and challenges in multimodal alignment.
+End-to-end 3D-language models unify perception, grounding, and captioning.
+D3Net [264] includes a detector, a speaker module for captioning, and a listener for grounding.
+Uni3DL [265] operates directly on point clouds with shared encoders for text and points, predicting outputs for segmentation, detection, grounding, and captioning.
+InstanceRefer [266] and LanguageRefer [267] combine linguistic and spatial cues for grounding in 3D point clouds.
+3DVG-Transformer [268] uses coordinate-guided attention and contextual feature fusion for grounding descriptions to 3D objects.
+SAGA [269] enables prompt-based segmentation using 3D Gaussian splats with SAM-derived affinity fields, supporting real-time open-vocabulary interaction across granularities.
+6 Datasets and Benchmarking
+Dataset Object Captioning Scene Captioning Dense Captioning Single Object Grounding Multiple Object Grounding Question Answering Situated Question Answering Dialogue Task Planning Real Synthetic Object Indoor Outdoor Human-Annotated Model-Annotated Template-Based # Language Pairs # Scenes
+Cap3D [270] 660k 660k
+Text2Shape [271] 75k 15k
+SceneVerse [272] 2.5M 68k
+nu-Caption [155] 420k -
+nu-Grounding [155] 280k -
+ScanRefer [101] 51k 800
+ReferIt3D [111] 125k 707
+Multi3DRefer [113] 60k 800
+Chat-3D v2 [81] 705 705
+EmbodiedScan [273] 970k 3.4k
+ScanEnts3D [274] 84k 700
+WildRefer [275] 30.8k 3.8k
+RIORefer [276] 1.4k 63k
+ARKitSceneRefer [277] 15k 1.6k
+ScanERU [112] 46k 706
+PhraseRefer [278] 170k 700
+DenseGrounding [114] - -
+3DVQA [279] 484k 707
+ScanQA (Azuma et al.) [58] 41k 800
+ScanQA (Ye et al.) [280] 10k 806
+3DMV-VQA [118] 50k 5k
+NuScenes-QA [117] 460k 34k
+CLEVR3D [281] 171k 8.7k
+SQA-3D [59] 33.4k 650
+MSQA/MSNN [282] 251k 1.7k
+3D-LLM [115] 200k 900
+ScanScribe [202] 278k 1.2k
+M3DBench [283] 327k -
+GPT4Point [154] 1M 1M
+LAMM [284] - -
+3D-GRAND/3D-POPE [285] 6.2M 40k
+LEO [152] - -
+3DMIT [157] 75k 1.5k
+Spartun3D [193] 133k -
+3D-CoT [286] 1.51M -
+Beacon3D [287] 5k 30
+SPAR-Bench [288] 7M 4k
+SpatialVQA [172] 1.3k -
+Space3D-Bench [289] 1k -
+City-3DQA [290] 450k -
+MMScan [291] 6.9M 5.2k
+Reason3D 2.5k -
+LEO-VL 700k -
+SpaCE-10 6k 811
 TABLE III: An overview of datasets used for 3D-Related tasks using Large Language Models.
@@
 We focus primarily on newer datasets which are used to evaluate methods in recent research papers.
-6 Datasets
 We now provide a high-level overview of the datasets that are used to train and evaluate 3D vision-language models.
-In Table III, we provide a list of datasets alongside the tasks they are used for, as well as information about the 3D scans and annotations.
+In Table III, we provide a list of datasets alongside the tasks they are used for, as well as information about their 3D data and annotations.
 In Fig.
-4, we present these datasets on a timeline, showing where each dataset sourced 3D information from.
+5, we present these datasets on a timeline, showing where each dataset sourced 3D information from.
 Current 3D vision-language datasets are almost exclusively generated by taking existing 3D vision datasets, and applying human, model or templated annotations to samples.
-As seen in Table III, a majority of existing datasets focus on real, indoor scenes, and this is partially explained by observing that the majority of existing datasets use 3D scans from ScanNet [361] and 3RScan [295].
+As seen in Table III, a majority of existing datasets focus on real, indoor scenes due to using 3D scans from either ScanNet [57] or 3RScan [203].
 Many of the datasets presented here share the same 3D data, and instead differ primarily through their choice of annotation strategy, and the 3D vision-language task they are designed to be used for.
+Among outdoor datasets, the majority use nuScenes [292], which focuses on LiDAR data for autonomous driving.
+There are three current strategies for collecting language data for these datasets: human-based, template-based and model-based annotation.
+Prior to ChatGPT [191], language data was collected from human annotators labelling data with natural language, or by having humans construct natural language ‘templates’, where appropriate classes and spatial references are taken from the dataset (such as from scene graphs) and placed into the template phrases.
+However, human-based annotation is expensive and time-consuming, and template-based annotation can have limited linguistic diversity.
+Recently, datasets have been using LLMs to analyze and annotate 3D data, either by asking the LLM to analyze a generated text description of each 3D scene, or by using 2D vision or vision-language models to analyze multiple 2D images of the scene and collating the responses.
+This has allowed 3D vision language datasets to contain orders of magnitude more language annotations, but can fail when the LLM does not generate accurate and naturalistic annotations for the samples.
+See Table III for a breakdown of how each dataset is annotated.
+Due to LLMs’ ability to act as multi-task learners, several recent datasets [154, 157, 152, 202, 283, 284, 285] have focused on collecting data for many different tasks, attempting to unify the input and output formats for each task (see section 4.3).
+These datasets typically draw 3D data from a greater range of datasets to expand data size and diversity (see Fig.
+5).
+There are a few strategies to improve the data diversity by leveraging multiple datasets, such as randomly replacing 10% of the objects in scenes with a same-category object from Objaverse [202], or randomly replacing objects in the scene graphs with objects from a generated object pool [281].
+Many datasets can also be used beyond their intended task.
+Dense captioning methods such as Scan2Cap [100] can use the referring expressions from grounding datasets like ScanRefer [101] to create the captions needed to train a model.
+Datasets that were designed for text-to-3D generation tasks, such as [271], can also be used for 3D-to-text captioning tasks.
+Other datasets are introduced as a way of testing the cross-dataset generalization abilities of models [276].
 Datasets for 3D navigation and manipulation using language are often designed around specific requirements and have a large overlap with existing bodies of research.
@@
 We omit further discussion here due to prior extensive coverage, and because many methods use 2D vision-language data rather than 3D-specific datasets.
+In Fig.
+IV, we provide a quantitative comparison between a number of existing 3D-VLM methods on the common testing datasets ScanRefer [101], Multi3DRefer [113], Scan2Cap [100], ScanQA [58] and SQA3D [59].
+We include these results to highlight the variety of metrics and datasets used even among similar methods.
 Figure 5: Timeline of datasets.
@@
 Datasets in orange are foundational 3D datasets without language annotations and datasets in blue are the annotated datasets used in 3D vision language tasks.
-Many existing datasets use 3D data from the same sources, such as ScanNet and 3RScan, and instead differ primarily through their choice of annotation strategy and targeted 3D vision task.
 Note that WildRefer also introduces new 3D data and annotations for vision-language tasks.
-Cap3D [349] is a 3D object captioning dataset developed upon 660k objects from the Objaverse [293] dataset.
-It is constructed by generating 2D image captions from multiple views of the 3D object, and consolidating those captions using image-text alignment and LLMs.
-Text2Shape [350] is a human-captioned form of the 8,447 tables and 6,591 chairs found in ShapeNet [296], combined with a procedurally generated dataset of primitive shapes labeled with template-based captions.
-It was originally used for generative text-to-3D shape tasks.
-SceneVerse [351] is a large-scale, multi-purpose dataset of annotated scenes made by compiling 68k scenes from existing 3D datasets.
-SceneVerse contains a total of 2.5 million vision-language pairs for object captioning, scene captioning and generating relative descriptions, primarily generated through the use of 3D scene graphs and LLMs.
-nu-Caption [271] is a captioned version of 420k LiDAR scans from the nuScenes [364] dataset, annotated using GPT-4 and 2D MLLMs.
-Captions include general scene descriptions, detailed descriptions of objects and their relationships, and the identification of potential risks on the road.
-nu-Grounding [271] builds on nu-Caption by focusing on the task of grounding, using the annotations from nuScenes to to create 280k pairs of questions and answers for visual grounding and grounded captioning.
-ScanRefer [218] introduces the task of 3D RGB-D grounding using a natural language expression by creating 51,583 human-annotated ‘referring expressions’ of 11,046 objects found in 800 scenes of ScanNet [361].
-The input consists of a point cloud of a scanned 3D scene alongside a free-form description of a specified target object, and the output is the corresponding bounding box of the object.
-ScanRefer provides an evaluation server and online benchmark  [^2] https://kaldir.vc.in.tum.de/scanrefer_benchmark/ to enable comparison between different methods.
-ReferIt3D [228] introduces a number of datasets (Nr3D, Sr3D and Sr3D+) consisting of objects from 707 ScanNet scenes.
-Similarly to ScanRefer, these object have been annotated with referring expressions, with a focus on queries where the scene contains multiple instances of the target class, and the referring expression is required to disambiguate between them.
-Nr3D contains 41,503 human-annotated, free-form utterances to refer to objects in 3D scenes, Sr3D contains 83,572 template-based utterances, and SR3D+ is a version of Sr3D with augmented utterances.
-ReferIt3D also provides an evaluation server and online benchmark  [^3] https://referit3d.github.io/benchmarks.html to enable comparison between different methods.
-Multi3DRefer [230] is a modified version of the ScanRefer dataset.
-Instead of referring expressions that always refer to one object in the scene, Multi3DRefer contains 6688 zero-target, 42,060 single target and 13178 multiple target referring descriptions collected for 11,609 objects in 800 ScanNet scenes.
-ChatGPT [286] is also used to rephrase the referring expressions.
-Chat-3D v2 [172] is another modified form of ScanRefer, where the referring expressions from 705 scenes in ScanNet are used to construct scene captions that describe the relationships between objects in the scene.
-The scene captions are generated using GPT-4 [365] by providing the model with ground-truth information about objects.
-The generated captions contain references to explicit ‘object identifiers’ that directly represent each object in the scene.
-EmbodiedScan [352] is an annotated combination of Matterport3D [285], 3RScan [295] and ScanNet [361], designed as a multi-modal, ego-centric dataset for 3D scene understanding.
-Segment Anything [366] and other annotation tools are used to provide 3D bounding boxes, semantic occupancy, and 970k template-based language descriptions across a total of 5185 scenes.
-ScanEnts3D [353] extends both ScanRefer [218] and ReferIt3D [228] by using professional annotators to link every object mentioned in referential sentences to their respective instances within a 3D scene.
-In the original paper this dataset is only used for training purposes, where it is found to improve the model’s performance on other visual grounding and captioning datasets.
-WildRefer [354] presents the STRefer and LifeRefer datasets, emphasizing human-centric in the wild settings with comprehensive 3D and linguistic human-annotations for 3D grounding.
-STRefer contains 5,458 referring expressions for objects in 662 scenes from the STCrowd [367] dataset, whereas LifeRefer contains 25,380 referring expressions for objects in 3,172 scenes from a new set of 3D scans acquired for this dataset.
-RIORefer [355] is a human-annotated version of 3RScan [295], used for 3D grounding.
-It consists of 63k descriptions of objects across 1,380 scenes.
-This dataset is introduced as a way of testing the cross-dataset generalization abilities of models, such as in the proposed “ScanRefer to RIORefer generalization” and “RIORefer to ScanRefer generalization” tasks.
-ARKitSceneRefer [356] is an annotated version of ARKitScenes [368] emphasizing the 3D grounding of small, everyday objects found in real-world indoor environments.
-It contains 15k descriptions of objects found in 1,605 scenes.
-ScanERU [229] is a modified and human-annotated version of ScanNet that combines 46k referring expressions from ScanRefer with 706 ScanNet scenes that have been modified to include a 3D human model gesturing to the referred object, using locations designated by human annotators.
-DenseGrounding [231], like Multi3DRefer, seeks to extend the task of 3D grounding to include multiple objects, however instead of a single referring expression referring to multiple objects, each input is a combined paragraph of referring queries which each correspond to a single object.
-These paragraphs are constructed using the nearest neighbors of random objects from ScanRefer and ReferIt3D, and combining their referring expressions together to form a paragraph.
-ScanQA (Azuma et al.) [234] is an annotated version of ScanNet with 41k question-answer pairs generated across 800 scenes.
-Questions are automatically generated using the referring expressions in ScanRefer and then refined by human annotators, whereas answers are derived entirely from human-annotators.
-This is the dataset that is typically referred to by the name “ScanQA”.
-ScanQA (Ye et al.) [357] was released contemporaneously with ScanQA (Azuma et al.) and is also a human-annotated form of ScanNet to be used as a 3D question answering dataset.
-Ye et al.contains 10k question-answer pairs for 806 ScanNet scenes.
-While Azuma et al.initially generates questions using referring expressions from ScanRefer, Ye et al.use human annotators for the entirety of question creation.
-3DMV-VQA [236] is an annotated version of 5k scenes from the Habitat-Matterport 3D Dataset (HM3D) [369], using the semantic information from HM3DSem [370] to generate 50k questions of four types: “concept”, “counting”, “relation” and “comparison”.
-These questions are generated as templates, and then transformed into natural language questions.
-NuScenes-QA [235] consists of 34k scenes from nuScenes [364] annotated with 460k template-style question-answer pairs generated using a constructed scene graph.
-Questions are divided into 5 types: “existence”, “counting”, “query-object”, “query-status” and “comparison”, and can include spatial reasoning.
-CLEVR3D [358] is an annotated form of 3RScan [295] designed for indoor 3D QA tasks.
-Template-based questions and answers are generated using the scene graph annotations from 3DSSG [371].
-Initially, 44k questions are generated for 1,333 scenes, however they use a “compositional scene manipulation” technique to randomly replace objects in the scene graphs with objects from a generated object pool, artificially leading to 171k questions across 8,771 scenes.
-SQA-3D [233] is another human-annotated version of ScanNet, proposing the problem of “situated question answering”, where each query consists of a description of the agent’s current location and situation, and a query relating to the scene.
-SQA-3D contains 33.4k questions across 650 scenes.
-3D-LLM [153] is an annotated version of Objaverse, ScanNet and HM3D/HM3DSem designed to handle a broad variety of tasks involving 3D vision and language.
-BLIP-2 [195] and ChatGPT [286] are used to generate 300k instances of ‘3D-language data’ that are used across a range of 3D vision-language tasks.
-ScanScribe [294] is an annotated version 1,185 scenes from ScanNet and 3RScan designed to act as a large-scale 3D scene-text pair dataset for pre-training 3D vision-language models.
-It utilizes ScanQA, ScanRefer and ReferIt3D for ScanNet data, and uses GPT-3 prompting to generate annotations for 3RScan to create a total of 278k scene descriptions.
-To increase the diversity of the dataset, scenes are synthetically generated by random replacing 10% of the objects in scenes with a same-category object from Objaverse.
-M3DBench [359] is a multi-modal instruction-following dataset consisting of 327k instruction-response pairs across a broad range of tasks.
-Data is collected from a large set of existing datasets (see Fig.
-4), alongside using GPT-prompting to generate additional annotations.
-GPT4Point [268] is an annotated form of over 1 million objects from Objaverse-XL, designed for 3D Captioning, 3D QA and other 3D tasks.
-Annotations are generated automatically through a hierarchical pipeline that fuses and improves captions from multiple views.
-LAMM [360] is a large scale multi-modal instruction-tuning dataset that covers tasks for both 2D and 3D vision.
-The 3D Data is sourced from 3RScan [295], CLEVR3D [358], 3DSSG [371] and ShapeNet [296], and is annotated using GPT [365] and template-based responses.
+Model ScanRefer Multi3DRefer Scan2Cap ScanQA (val) SQA3D (test)
+Acc@0.25 Acc@0.5 F1@0.25 F1@0.5 B-4@0.5 C@0.5 M@0.5 R@0.5 M R C EM B-1 B-4 EM EM-R
+Specialists
+ScanRefer [101] 37.3 24.3 - - - - - - - - - - - - - -
+Scan2Cap [100] - - - - 22.4 35.2 22.0 44.8 - - - - - - - -
+Vote2Cap-DETR++ [109] - - - - 34.5 61.8 26.2 54.4 - - - - - - - -
+ScanQA [58] - - - - - - - - 13.1 33.3 64.9 21.1 30.2 10.1 47.2 -
+SQA3D [59] - - - - - - - - 13.5 34.5 - - 30.5 11.2 46.6 -
+D3Net [264] - 37.9 - 32.2 35.7 62.6 25.7 53.9 - - - - - - - -
+3D-ViSTA [202] 50.6 45.8 - - 34.0 66.9 - - 13.9 35.7 69.6 22.4 - - 48.5 -
+3DVG-Transformer [268] 45.9 34.5 - 25.5 - - - - - - - - - - - -
+SIG3D [151] - - - - - - - - - - - - - - 52.6 -
+Generalists (Point-Based Methods)
+Chat-3D [80] - - - - - - - - - - 53.2 - - - - -
+Chat-3D v2 [81] 42.5 38.4 45.1 41.6 31.8 63.9 - - 16.1 40.1 87.6 - 38.4 7.3 54.7 -
+LL3DA [19] - - - - 36.0 62.9 26.0 55.0 15.9 37.3 76.8 - - 13.5 - -
+Scene-LLM [147] - - - - - - - - 16.6 40.0 80.0 27.2 - - 53.6 -
+LEO [152] - - - - 38.2 72.4 27.9 58.1 16.2 39.3 101.4 21.5 - 11.5 50.0 52.4
+Chat-Scene [156] 55.5 50.2 57.1 52.4 36.3 77.1 - - 18.0 41.6 87.7 21.6 43.2 14.3 54.6 57.5
+3D-LLaVA [171] 51.2 40.6 - - 36.9 78.8 27.1 57.7 18.4 43.1 92.6 27.0 - 17.1 54.5 56.6
+Robin3D [163] 60.8 55.1 64.9 59.7 38.4 87.2 - - 19.2 44.0 - - - - 56.9 59.8
+PerLA [164] - - - - - - - - 17.4 39.6 78.1 - - 14.5 - -
+Grounded-3DLLM [295] 47.9 44.1 45.2 40.6 35.5 70.6 - - 15.2 37.1 72.7 - - - -
+Inst3D-LLM [168] 57.8 51.6 58.3 53.5 38.3 79.7 - - - - 88.6 24.6 - - - -
+Spartun3D-LLM [193] - - - - - - - - - - - - - - 55.0 -
+LSceneLLM [167] - - - - - - - - 18.0 40.8 88.2 - - - - -
+Generalists (Multi-View Image-Based Methods)
+3D-LLM [115] 30.3 - - - - - - - 14.5 35.7 69.4 20.5 39.3 12.0 - -
+LLaVA-3D [162] 54.1 42.4 - - 41.1 79.2 30.2 63.4 - - 91.7 27.0 - - 55.6 -
+Video-3D LLM [165] 58.1 51.7 58.0 52.7 41.3 83.8 28.5 61.7 19.8 49.0 102.1 30.1 47.1 16.2 58.6 -
+3DRS [176] 62.9 56.1 60.4 54.9 41.6 86.1 - - - - 104.8 30.3 - - 60.6 -
+GPT4Scene-HDM [166] 62.6 57.0 64.5 59.8 - - - - 18.9 46.5 96.3 - 44.4 15.5 60.6 63.3
+VeBrain [296] 66.4 60.2 67.8 62.7 - - - - 20.1 48.2 101.5 - 47.7 17.3 - -
+VG LLM [173] 51.0 45.1 - - 38.7 74.1 28.0 61.0 - - - - - - - -
+Ross3D [170] 61.1 54.4 59.6 54.3 43.4 81.3 30.3 66.9 20.9 50.7 107.0 30.8 - 17.9 63.0 65.7
+LEO-VL [175] - - - - - - - - 17.4 39.6 78.1 - - 14.5 - -
+Spatial-MLLM [172] - - - - - - - - 18.4 45.0 91.8 - 44.8 14.8 55.9 58.7
+Human (from [297]) - - - - - - - - - - - 51.6 - - 90.1 -
+Best Model 66.4 60.2 67.8 62.7 43.4 87.2 30.3 66.9 20.9 50.7 107.0 30.8 47.7 17.9 63.0 65.7
+Model vs. Human Gap - - - - - - - - - - - 40.6% - - 30.1% -
+TABLE IV: Quantitative comparison of different models on common 3D vision-language benchmarks.
+Results are from original papers where applicable.
+Best results in each category (Specialists and Generalists) are highlighted in bold.
+The bottom row shows the performance gap between the best AI models and human performance.
 7 Challenges and Opportunities
-Despite progress in integrating LLMs with 3D data, challenges remain in data representation, computational efficiency, and benchmarks, necessitating innovative solutions.
-Representation Choice has a large impact on the performance of 3D vision language models.
-Currently, point clouds are predominantly used to represent both indoor (e.g.
-vertices of meshes) and outdoor (e.g.
-LiDAR point clouds) environments due to their simplicity and neural network compatibility.
-However, they struggle to capture the fine details crucial for accurate, rich spatial models.
-Developing novel 3D scene representations that more effectively bridge the gap between spatial information and language could unlock new levels of understanding and interaction.
-By finding innovative ways to encode language and semantic information in 3D representations, such as using distilled language and semantic embeddings [30, 316], could help bridge the gap between these two modalities.
-Computational Demands of both 3D data processing and LLMs pose significant challenges.
-Scalability remains a concern as the complexity of 3D environments and the size of language models increase.
-Advances in LLM architectures designed for adaptivity and computational efficiency could significantly broaden their application scope.
-Improving Benchmarks is essential for comprehensively evaluating and advancing the capabilities of multi-modal LLMs in 3D tasks.
-Current benchmarks have limited scope, especially regarding 3D reasoning, hampering evaluation of spatial reasoning skills as well as development of 3D decision-making/interaction systems.
-Moreover, the metrics in use today fall short of capturing the full spectrum of LLMs’ capabilities in 3D environments.
-Crafting task-specific metrics that more precisely measure performance across diverse 3D tasks is essential.
-Finally, the granularity of current scene understanding benchmarks is overly simple, limiting insight into complex 3D environment comprehension.
-A more diverse array of tasks is required.
-Safety and Ethical Implications are imperative to consider when employing LLMs for 3D understanding.
-LLMs can hallucinate and output inaccurate, unsafe information leading to erroneous decisions in critical 3D applications.
-Moreover, LLMs often fail in ways that are unpredictable and challenging to interpret.
-They may also inherit societal biases present in training data, disproportionately disadvantaging certain groups when making predictions in real-world 3D scenes.
-It is crucial to approach the use of LLMs in 3D contexts with caution, employing strategies to create more inclusive datasets, robust evaluation frameworks for bias detection and correction, and mechanisms to minimize hallucinations, ensuring responsible and equitable outcomes.
+Despite the progress made in integrating LLMs with 3D data, challenges remain in data representation, model training, evaluation and safety, necessitating further research.
+7.1 Better 3D Representations for Language Alignment
+A central challenge for 3D-LLMs is the choice of 3D data representation, as it directly affects both geometric fidelity and language alignment.
+Point clouds dominate due to their simplicity but lack fine-grained detail, while voxel grids and meshes provide richer geometry at the cost of high memory use or poor neural compatibility.
+Recent works have thus explored representations [231, 230, 298] that combine semantics with geometry, including differentiable formats such as NeRFs [49] and 3D Gaussian Splats [53], as well as VGGT-style geometry-aware encoders [200, 299].
+Another promising direction is structured, part-level representations [300], which support hierarchical alignment between language and 3D across multiple granularities—from scenes, to objects, down to fine components.
+7.2 Need for 3D-centric Pre-training and Architectures
+Existing 3D-LLM methods typically rely on a frozen 3D encoder combined with an LLM that has been pretrained purely on text sequences.
+During fine-tuning, the majority of existing methods use alignment modules that transform native 3D features to sequences of 1D tokens that are directly aligned with the text embedding space of the pretrained LLMs.
+This reduction from inherently three-dimensional structures to one-dimensional text space leads to a substantial loss of spatial information, and these methods have little direct knowledge of 3D spatial structures beyond what can be inferred from linguistic priors.
+This highlights an important opportunity: to move toward 3D-centric pre-training and architectures that treat 3D structure as a first-class representation.
+Promising directions include introducing pretraining objectives that explicitly teach LLMs 3D relationships, such as masked autoencoding in a 3D token space, geometric consistency, and auxiliary reconstruction losses [170], and jointly pretraining the 3D encoder and the LLM so that both components co-adapt.
+In parallel, alignment modules should be redesigned from simple projection into language space toward bidirectional, structure-aware interfaces that better integrate 3D and text.
+7.3 From Static to Dynamic Geometric Understanding
+Most existing 3D-LLMs focus on static and coarse-grained comprehension of object categories, attributes (e.g.
+size, color), positions, and spatial relationships.
+However, many real-world tasks also require reasoning about dynamic properties from static 3D scans, such as predicting object affordances [301], functionality, and physical characteristics [302, 303] (e.g.
+material, fragility, deformability), as well as motion dynamics (e.g.
+free fall) and inter-object interactions (e.g.
+collision, friction [304]).
+Beyond this, fine-grained geometric reasoning, including symmetry, perspective, occlusion [137], and containment remain essential for precise spatial understanding.
+A next step is to move from inferring dynamics from static geometry to directly modeling dynamic 3D scenes (4D, i.e.
+3D + time) [179, 305, 306], enabling LLMs to capture both spatial and temporal aspects of the physical world.
+7.4 Acquisition of Higher Spatial Reasoning Ability
+Recent progress such as OpenAI-o3 and DeepSeek-R1 [307] have shown that reinforcement learning (RL) can significantly enhance the reasoning ability of LLMs by providing reward signals beyond next-token prediction.
+However, current 3D-LLMs typically generate a single textual answer without engaging in step-by-step multi-hop reasoning or grounding their outputs to corresponding 3D structures.
+This lack of explicit reasoning chains and spatial grounding limits interpretability, which is essential for higher-order spatial reasoning.
+An important opportunity is to leverage RL for training 3D-LLMs with explicit reasoning traces and verifiable grounding.
+In particular, constructing 3D-specific chain-of-thought (3D-CoT) datasets could teach models to decompose problems into multi-step spatial reasoning processes.
+Together with well-designed reward functions and multi-stage training schemes (e.g.
+pre-training followed by RL-based post-training), 3D-LLMs may gradually acquire more robust, interpretable, and generalizable spatial reasoning skills, which is a largely unexplored direction.
+7.5 Lack of 3D-text Paired Data as a Bottleneck
+The success of LLMs is largely driven by massive text corpora, and 2D-VLMs rely on billions of image-text pairs such as LAION-5B [308].
+In contrast, existing 3D vision-language datasets remain orders of magnitude smaller, with the largest containing only millions of 3D-text pairs (Tab.
+III).
+While VLMs can generate textual descriptions for 3D datasets, the scale of available 3D data is still far behind that of 2D images.
+Moreover, most existing annotations provide only coarse object-level descriptions, whereas fine-grained alignment across spatial, geometric, and semantic levels is critical for 3D-LLMs.
+This scarcity in both quantity and quality of 3D-text data is a major bottleneck for scaling and capability development in 3D-LLMs.
+Promising directions to alleviate this include: (1) using simulators [309] to synthesize diverse 3D data of objects, scenes, and interactions, combined with automated annotation tools for multi-granularity language alignment; and (2) leveraging 3D generative models [310, 311] to produce 3D content directly from textual descriptions, thereby expanding data scale while ensuring consistent 3D-text alignment.
+7.6 Need for Accurate and Robust Evaluation Metrics
+Current evaluations of 3D-LLMs largely follow a VQA-style protocol [58], where performance is measured only at the text level via similarity scores.
+However, complex 3D geometric relationships are often difficult to capture through language alone, and text-based answers do not necessarily reflect genuine spatial understanding.
+In particular, hallucinations pose a major issue: models may output the correct answer by exploiting linguistic priors rather than grounding their reasoning in the underlying 3D scene [312].
+Existing metrics fail to detect this behavior, raising concerns about reliability.
+Going forward, the community needs more robust and interpretable evaluation methods that explicitly test 3D-grounded understanding and reasoning.
+7.7 Safety Issues of 3D Embodied Agents
+The ultimate goal of understanding 3D scenes is to enable actions, interactions, and transformations in the 3D world, where LLMs can play a crucial role.
+In Sec.
+4.5, we discuss how current research leverages LLMs to empower 3D embodied agents in tasks such as planning, navigation, and manipulation, illustrating the feasibility of LLMs directly generating action outputs.
+However, a critical yet often overlooked challenge is the potential safety risks when LLMs generate physical actions.
+For instance, when controlling a robotic arm to assist an elderly person during feeding, the agent must carefully regulate its strength and speed to prevent harm.
+LLMs, however, may lack precise physical awareness or produce hallucinations, leading to excessive force or other hazardous actions.
+Thus, safety and reliability of 3D embodied agents must be ensured through more detailed mechanisms, including the introduction of more physical constraints [313], real-time feedback systems, and human-value alignment [314].
+These measures are essential to achieving truly safe, reliable and trustworthy actions in complex 3D environments.
 8 Conclusion
@@
 Systematically reviewing methodologies, applications, and emergent abilities of LLMs in processing, understanding, and generating 3D data, the survey underlines the transformative potential of LLMs across a spectrum of 3D tasks.
-From enhancing spatial comprehension and interaction within 3D environments to driving forward the capabilities of embodied AI systems, LLMs emerge as pivotal in advancing the field.
 Key findings include the identification of LLMs’ unique advantages such as zero-shot learning, advanced reasoning, and extensive world knowledge, which are instrumental in bridging the gap between textual information and spatial interpretation.
@@
 The exploration of other 3D vision-language methods alongside LLMs reveals a rich landscape of research aiming to deepen our understanding of the 3D world.
-Furthermore, the survey highlights significant challenges such as data representation, model scalability, and computational efficiency, suggesting that overcoming these hurdles is crucial for the full realization of LLMs’ potential in 3D applications.
+Furthermore, the survey highlights significant challenges such as data representation, model scalability, acquisition of spatial reasoning, data bottleneck and evaluation metrics, suggesting that overcoming these hurdles is crucial for the full realization of LLMs’ potential in 3D applications.
 In conclusion, this survey not only offers a comprehensive overview of the current state of 3D tasks using LLMs but also sets the stage for future research directions.